### v1

# On computing and the complexity of computing higher-order $U$-statistics, exactly

Abstract
Higher-order $U$-statistics abound in fields such as statistics, machine learning, and computer science, but are known to be highly time-consuming to compute in practice. Despite their widespread appearance, a comprehensive study of their computational complexity is surprisingly lacking. This paper aims to fill that gap by presenting several results related to the computational aspect of $U$-statistics. First, we derive a useful decomposition from an $m$-th order $U$-statistic to a linear combination of $V$-statistics with orders not exceeding $m$, which are generally more feasible to compute. Second, we explore the connection between exactly computing $V$-statistics and Einstein summation, a tool often used in computational mathematics, quantum computing, and quantum information sciences for accelerating tensor computations. Third, we provide an optimistic estimate of the time complexity for exactly computing $U$-statistics, based on the treewidth of a particular graph associated with the $U$-statistic kernel. The above ingredients lead to a new, much more runtime-efficient algorithm of exactly computing general higher-order $U$-statistics. We also wrap our new algorithm into an open-source Python package called u-stats. We demonstrate via three statistical applications that u-stats achieves impressive runtime performance compared to existing benchmarks. This paper aspires to achieve two goals: (1) to capture the interest of researchers in both statistics and other related areas further to advance the algorithmic development of $U$-statistics, and (2) to offer the package u-stats as a valuable tool for practitioners, making the implementation of methods based on higher-order $U$-statistics a more delightful experience.
Keywords: $U$-statistics, Statistical/Scientific Computing, Tensors, Einstein summation, Treewidth
1 Introduction
Higher-order $U$-statistics are prevalent in various disciplines, including statistics (robins2008higher; bonvini2022fast; bonvini2024doubly; neykov2021minimax; yao2018testing; breunig2024adaptive; shao2025u), machine learning (clemenccon2008ranking; clemenccon2013maximal; shen2025engression), random graphs (chatterjee2024higher), Bayesian networks and statistical physics (pearl1988probabilistic; koller2009probabilistic; shpitser2011efficient; li2020random), just to name a few. It is well-known in the statistical community that $U$-statistics are highly time-consuming to compute at large. In the worst-case, an $m$-th order $U$-statistic over a sample of size $n$ takes roughly $O(n^{m})$ steps to compute, because by definition (see Definition 1) one needs to sum over all $m$-out-of-$n$ permutations (resp. over $n^{m}$ terms). If $n=10^{4}$ and $m=7$, $n^{m}\approx 10^{28}$!
More specifically, in causal inference, robins2008higher; robins2016technical; robins2017minimax developed a so-called “higher-order influence function (HOIF)” framework for constructing nearly rate-optimal estimators of average treatment effects under ignorability (hernan2023causal) under various structural assumptions, using higher-order $U$-statistics with order roughly of $\sqrt{\log n}$. Despite their appealing statistical properties, implementing such complex estimators remains challenging in practice and computational complexity is one important bottleneck. Here we quote a sentence from a recent graduate-level textbook on causal inference by wager2024causal (page 44) to highlight such a quandary:
One challenge with the HOIF estimator of robins2017minimax, however, is that to date it has been challenging to implement in practical applications …
This paper was initially motivated by (at least partially) addressing the computational aspect of this challenge, but we will present general-purpose theory and algorithm for computing general higher-order $U$-statistics beyond those for HOIF estimators. We will revisit HOIF estimators in Section 4.1.
To tackle the difficulty of computing higher-order $U$-statistics, statisticians and computer scientists have long considered the so-called “incomplete $U$-statistics”, where one sums over only a random subset of all $m$-out-of-$n$ permutations (blom1976some; janson1984asymptotic; nowicki1988subgraph; shao2025u). Another natural approach is to choose a subsample of size $n^{\prime}<n$ from the whole sample and then compute a complete $U$-statistic from that subsample. Both strategies are akin to randomized or sketching algorithms in the theoretical computer science literature (mahoney2011randomized). The randomization approach certainly injects extra randomness in the final statistical results, which may result in inconsistent statistical conclusions across different randomly selected permutations/subsamples (guo2025rank). For a restricted class of higher-order $U$-statistics with decomposable and symmetric (or more precisely permutation-invariant) kernels, kong2018estimating cleverly recognize that exact computation can be accomplished by repeated matrix multiplications. A somewhat similar strategy has also been considered in zhang2025adaptive, in which an incomplete $U$-statistic with a special kernel structure is computed more efficiently via dynamic programming. For asymmetric kernels, their strategy essentially computes an incomplete $U$-statistic instead. All the above approaches may suffer from efficiency loss by not using all the available information.
The literature on (the complexity of) exactly computing higher-order $U$-statistics is rather scarce. This is hardly surprising because the exact time complexity of computing higher-order $U$-statistics seems to simply equal the total number of summations. However, as we will demonstrate in the paper, exactly computing $U$-statistics can be improved compared to a brute-force summation over all summands (using for-loops) when the kernel enjoys a particular decomposable structure specified in Definition 9 later. Interestingly, as we will exhibit in Section 4, such a structure holds in many statistical applications. It will also be clear later that it is helpful to bring in tools such as the Einstein summation from the computational mathematics community to obtain improved algorithms of computing $U$-statistics by leveraging such a structure. It is worth noting that Section 3 of liu2024hypothesis considered to use di-graphs to represent the structure of a class of $U$-statistics, a strategy conceptually similar to ours. Yet our new algorithm is designed to cover general higher-order $U$-statistics.
Main contributions
The main contributions of our paper can be summarized as below.
  • We design a new algorithmic framework for exactly computing higher-order $U$-statistics and another closely related objects called $V$-statistics by leveraging the Einstein summation operation and related concepts from combinatorics and graph theory. Thanks to the connection with graph theory, we can provide an optimistic estimate of the time complexity of exactly computing $U$-statistics (see Section 3.3 for details). Although the idea is conceptually simple, to the best of our knowledge, it has not been thoroughly explored in the existing statistical literature.
  • Despite the ubiquity of $U$-statistics (see later this section), a dedicated, general-purpose software package for computing them seems to be lacking. To fill in this void, we developed a Python package called u-stats, which implemented the aforementioned algorithm. We demonstrate that the runtime of u-stats achieves the state-of-the-art performance in various examples, including computing HOIF estimators, motif counts of networks, and the distance covariance in the independence testing problem. In fact, as will be clear in Section 4, u-stats can dramatically improve the runtime compared to popular benchmarks. u-stats currently supports both CPU and GPU infrastructure. Having access to GPU can further improve the runtime because our algorithm heavily utilizes tensor operations such as the Einstein summation. We envision that the package u-stats can be a valuable tool for practitioners.
Remark 1.
We focus on scalar-valued $U$-statistics in this paper, but we plan to upgrade our package u-stats to handle vector- or matrix-valued cases in its next version. In this paper, one floating point operation is considered as the basic computation unit.
Related Literature
In this part, we provide an incomplete survey on the related literature of $U$-statistics, and the relevant techniques/tools for developing our new algorithm.
Statistics and machine learning
$U$-statistics are frequently found in statistics and machine learning. For example, in certain machine learning tasks (clemenccon2008ranking; clemenccon2013maximal; shen2025engression), the empirical training risk can be written as a $U$-statistic instead of a sample mean. Various (conditional) dependence measures can also be represented as $U$-statistics (yao2018testing; neykov2021minimax; chatterjee2021new; azadkia2021simple; shao2025u; zhang2025adaptive); see Section 4.3 for an example. As mentioned before, in causal inference, HOIF has been shown to be a theoretically appealing framework for constructing nearly rate-optimal estimators for various causal effect parameters. Estimators based on HOIF are higher-order $U$-statistics; see Section 4.1 for concrete forms. As mentioned, due to the computational difficulty, HOIF estimators have not been comprehensively applied in real-life applications.
Motif counts in random graphs
Higher-order $U$-statistics also appear very often in statistical inference over networks, often modeled as random graphs (chatterjee2024higher; jin2025counting). For instance, chatterjee2024higher estimate motif counts (such as triangles, 4-cliques, and etc.) by higher-order $U$-statistics and use these estimated motif counts to test certain properties of the underlying network. Motif counts also have widespread applications in bioinformatics (wernicke2006fanmod). As will be shown in Section 4.2, using the state-of-the-art software package igraph, for a moderately dense network, computing motif counts can be quite time consuming, necessitating the development of new algorithmic tools for $U$-statistics.
Tensor operations
Since computing UU/$V$-statistic is closely related to numerical integration using Monte Carlo sampling for which tensor operations are quite useful, our work is inspired by related techniques developed in the computational mathematics and physics communities. In particular, our algorithm utilizes the Einstein summation operation, an important tool for tensor computations. For example, in quantum computing/information, tensor networks are often used to describe many-body quantum systems and a very useful operation called tensor contraction is implemented by the Einstein summation in practice (markov2008simulating; stoudenmire2018learning). As will be seen later, our proposed algorithm is similar to tensor contraction in spirit.
Probabilistic graphical models and statistical physics
Our work is also related to the literature on probabilistic graphical models (pearl1988probabilistic; shpitser2011efficient), in which a common problem is to compute certain marginal distributions from the joint distribution, (nested) factorized according to Bayesian network. It can be reduced to computing $U$-statistics with kernels satisfying the decomposable structure to be defined later. Similarly, in both classical (edwards1975theory; li2020random) and quantum many-body systems (bernaschi2024quantum), the corresponding log-Gibbs measure can often be written as a multi-linear sum over all particles. The summation is taken over all possible configurations of each particle. Thus the log-Gibbs measure can be viewed as (a linear combination of) higher-order $U$-statistics.
Notation and organization
In this section, we first collect some frequently used notation. We adopt the following convention: we use the demarcation $\{\}$ for a set, representing an unordered collection of distinct elements; whereas we use the demarcation $()$ for a tuple, meaning an ordered collection of elements which may contain duplicates. Given a positive integer $m$, we let $m]\coloneqq\{1,2,\ldots,m\}$. To reduce clutter, we write $\bar{i}_{m}\equiv(i_{1},\cdots,i_{m})$ for short. $\mathbb{N}$ denotes the set of natural numbers including zero, while $\mathbb{N}^{+}$ denotes the set of all positive integers. Given a tuple $\bm{s}$, we let $\mathsf{set}(\bm{s})\coloneqq\{s_{k}\mid k\in[m]\}$ denote the set containing all elements appear in $\bm{s}$. Given a non-empty and finite set $S$ and a positive integer $m$, all $m$-permutations of elements in $S$ is denoted by $\mathsf{Perm}(S,m)$. All partitions of $S$ is denoted by $\Pi_{S}$. If $S$ happens to be integer set $[m]$, $\Pi_{S}$ is simplified to $\Pi_{m}$. Given sets $S_{1},\cdots,S_{m}$, $\times_{i=1}^{m}S_{i}$ denotes their Cartesian product. When $S_{1}=\cdots=S_{m}=S$, we simply write $S^{m}\equiv\times_{i=1}^{m}S_{i}$. $S^{*}\coloneqq\cup_{m\in\mathbb{N}^{+}}S^{m}$ denotes all tuples possibly generated by $S$. Given a tuple $\bm{s}\in S^{m}$ and a tuple $\alpha\in[m]^{*}$, we let $\bm{s}[\alpha]\coloneqq(s_{\alpha_{k}})_{k=1}^{|\alpha|}$.
Terminologies from graph theory will be frequently encountered. Throughout this paper, all graphs are assumed to be finite and undirected without self-loops, but are allowed to have isolated vertices, unless stated otherwise. A graph $G$ is denoted by a pair $(V,E)$, where $V\equiv V(G)$ is the finite vertex set, and $E\equiv E(G)$ is the edge set of $G$. Two vertices $u$ and $v$ are said to be adjacent if $\{u,v\}\in E$. For any vertex $v\in V$, its neighbor set is denoted as $N_{G}(v)\coloneqq\{u\in V\mid\{u,v\}\in E\}$, and the degree of $v$ is defined as $\deg(v,G)\coloneqq|N_{G}(v)|$. A graph $G^{\prime}$ is a subgraph of graph $G$, if $V(G^{\prime})\subseteq V(G)$ and $E(G^{\prime})\subseteq E(G)$. For any subset $V^{\prime}\subseteq V$, the induced subgraph $G_{V^{\prime}}$ has vertex set $V^{\prime}$ and edge set $E_{V^{\prime}}\coloneqq E(G_{V^{\prime}})\equiv\{\{u,v\}\in E\mid\forall\,u,v\in V^{\prime}\}$. A subset of vertices $C\subseteq V$ is said to be a clique if $G_{C}$ is complete, where a complete graph is simply a graph with every pair of vertices connected by an edge.
The remainder of our paper is structured as follows. Section 2 introduces and reviews several concepts central to the paper, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation (Section 2.2), and notions (including vertex elimination, treewidth, and quotient graphs) from graph theory (Section 2.3). Readers who are familiar with related materials are recommended to skip the relevant subsections. In Section 3, the new algorithm for exactly computing $U$-statistics is presented. We also illustrate how the algorithm proceeds using concrete examples to facilitate understanding. We then demonstrate the utility of our algorithm using three examples in Section 4, including computing HOIF estimators, motif counts in networks (Section 4.2), and computing distance covariances (Section 4.3). In our own opinion, Section 4 is the most important part of this paper. Section 5 concludes the paper with a discussion of future directions.
2 Preliminaries
In this section, we will review concepts important for later development, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation operation (Section 2.2), and useful notions from graph theory (Section 2.3). The way that we define $U$- or $V$-statistics may look strange to statisticians, but we find such definitions more convenient to work with when combined with the Einstein summation formulation.
2.1 UU/$V$-Statistics and Tensors
We define $U$-statistics and $V$-statistics as follows. As we just focus on computation, the sample space is regarded as a non-empty set.
Definition 1 ($U$-statistics).
Let $\mathbb{X}$} be a non-empty set, $m,n$ be positive integers satisfying $m\leq n$, and $h:{\mathbb{X}}^{m}\rightarrow{\mathbb{R}}$ be a function. Given a tuple $\bm{X}\in{\mathbb{X}}^{n}$, any $U$-statistic of order $m$ with kernel $h$ takes the following form
$\mathbb{U}(h,\bm{X})\coloneqq\sum_{\alpha\in\mathsf{Perm}([n],m)}h(\bm{X}[\alpha]).$
Definition 2 ($V$-statistics).
Let $\mathbb{X}$} be a non-empty set, $m,n$ be positive integers and $h:{\mathbb{X}}^{m}\rightarrow{\mathbb{R}}$ be a function. Given a tuple $\bm{X}\in{\mathbb{X}}^{n}$, any $V$-statistic of order $m$ with kernel $h$ takes the following form
$\mathbb{V}(h,\bm{X})\coloneqq\sum_{\alpha\in[n]^{m}}h(\bm{X}[\alpha]).$
Definition 3 (Tensors).
Given positive integers $m,n>0$, an $m$-th order tensor (i.e. an $m$-dimensional array) $T$ of size $n$ is defined as a map from $n]^{m}$ to $\mathbb{R}$}: for any $\alpha\in[n]^{m}$, $T(\alpha)\in{\mathbb{R}}$; and there is no time complexity to access an entry in $T$, i.e. to evaluate $T$ on $\alpha$, for any $\alpha\in[n]^{m}$. Here we can safely assume that each dimension of a tensor has the same size (i.e. number of entries in each dimension) unless explicitly mentioned otherwise.
In statistical applications, $\mathbb{X}$} and the tuple $\bm{X}$, respectively, correspond to the sample space and the observed data sample of size $n$. Unlike $U$-statistics, computing $V$-statistics amounts to summing over all $n]^{m}$ without being restricted to permutations. As will be clear later in this paper, $V$-statistics can be represented as summations over higher-order tensors. This unique property renders computing $V$-statistics, but not $U$-statistics, amenable for tensor operations such as the Einstein summation (einstein1916grundlage). In particular, if the kernel function $h$ satisfies certain decomposable structure (see Definition 9 later), the computational complexity of $V$-statistics can be notably reduced. Our algorithmic framework is thus to first turn a $U$-statistic into a linear combination of $V$-statistics and then compute each individual $V$-statistic via the Einstein summation.
2.2 The Einsum Notation and Einsum Operation
As mentioned, our algorithm to be revealed will rely on the Einstein summation operation (abbreviated as the Einsum operation) over tensors. In this section, we give a brief but (hopefully) pedagogical overview for readers unfamiliar with these related concepts. We also steer readers to klaus2023compiling or blacher2024einsum for more details.
The Einsum operation to be defined serves as a fundamental computational primitive in our proposed algorithm for efficiently and exactly computing UU/$V$-statistics. The Einsum operation generalizes the classical Einstein summation convention introduced in (einstein1916grundlage). The modern Einsum operation, now implemented in popular Python packages such as numpy and pytorch, can be used for efficient tensor computation. To explain the Einsum operation, we first introduce a concept called the Einsum notation.
Definition 4 (The Einsum Notation).
Let $m$ and $K$ be two positive integers. An Einsum notation with respect to [w.r.t.] $[m]$ is a pair
${\mathcal{N}}=\big{(}{\mathcal{A}};B\big{)}\quad\text{with}\quad{\mathcal{A}}\coloneqq(A_{1},\cdots,A_{K}),A_{k},\,B\in[m]^{*},\;k\in[K],$
where we recall that $m]^{\ast}$ denotes all tuples possibly generated by elements in $[m]$, and $B$ needs to satisfy the following constraint:
  • $B[i]\neq B[j]$, for any $i,j\in[|B|]suchthati\neq j$,
  • $\mathsf{set}(B)\subseteq\bigcup_{k\in[K]}\mathsf{set}(A_{k})$.
${\mathcal{A}}\coloneqq(A_{k})_{k=1}^{K}$ is referred to as the input tuple list of $\mathcal{N}$} and $B$ is referred as the output tuple. The index set of $\mathcal{N}$} is the set of all indices appearing in the input tuple list, and is denoted by $\mathsf{Indices}({\mathcal{N}})$. Without loss of generality, we always take $m$ to be the smallest $m$ such that $m]=\mathsf{Indices}({\mathcal{N}})$. If $B=\emptyset$, we then simply write $\mathcal{A}$} as ${\mathcal{N}}=({\mathcal{A}};\emptyset)$ and call $\mathcal{A}$} as the Einsum notation.
It is noteworthy that in the above definition, the index set $[m]$ is simply chosen for notational convenience and in principle the indices do not have to be integers or even numbers. We now define the Einsum operation based on the Einsum notation.
Definition 5 (The Einsum Operation).
Let ${\mathcal{N}}=({\mathcal{A}};B)$ be an Einsum notation w.r.t. the index set $I=[m]$. Suppose that a set of tensors ${\mathcal{T}}\coloneqq(T_{k})_{k=1}^{K}$, each of size $n$ and of order $(m_{k})_{k=1}^{K}$, can be represented by $\mathcal{N}$}, i.e. $m_{k}=|A_{k}|$ for any $k\in[K]$. An Einsum operation, denoted as $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$, takes $\mathcal{T}$} and $\mathcal{N}$} as inputs and outputs another tensor $T^{\rm out}:[n]^{|B|}\rightarrow{\mathbb{R}}$ as follows: for any $\alpha\in[n]^{|B|}$,
$T^{\rm out}(\alpha)=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{m}\\ \alpha^{\prime}[B]=\alpha\end{subarray}}\prod_{k=1}^{K}T_{k}(\alpha^{\prime}[A_{k}]).$
As a special case, if $B=\emptyset$, then $T^{\rm out}(\alpha)$ reduces to the following scalar:
$T^{\rm out}=\sum_{\alpha^{\prime}\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha^{\prime}[A_{k}]).$
We now use several concrete examples to illustrate the above definitions.
Example 1.
We take matrix multiplication as an example to illustrate the Einsum operation just defined. Let $T_{1},T_{2}$ be two input matrices, which are also 2nd-order tensors. Therefore ${\mathcal{T}}=(T_{1},T_{2})$. Suppose we are interested in multiplying $T_{1}$ and $T_{2}$, denoted as $T^{\rm out}$: for any $i,j\in[n]$,
$T^{\rm out}(i,j)=\sum_{k=1}^{n}T_{1}(i,k)T_{2}(k,j).$
The matrix multiplication corresponds to the Einsum operation with Einsum notation ${\mathcal{N}}=({\mathcal{A}};B)$, with input tuple list ${\mathcal{A}}=((1,2),(2,3))$ and output tuple $B=(1,3)$ and $|B|=2$. The index set is simply $I=[3]$. This means that the operation sums over the dimensions of $T_{1}$ and $T_{2}$ corresponding to index $2$ and reserves indices 11 and $3$. $T^{\rm out}(\alpha)$ for any $\alpha=(i,j)\in[n]^{2}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ because by definition:
$\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ $\equiv\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}\prod_{k=1}^{2}T_{k}(\alpha^{\prime}[A_{k}])$
$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])$
$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[1,2])T_{2}(\alpha^{\prime}[2,3])$
$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[3])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha[3])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(i,\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],j),$
so the claim holds by identifying $\alpha^{\prime}[2]=k$.
Example 2.
Here we consider an example involving higher-order tensors. Given three tensors ${\mathcal{T}}=(T_{1},T_{2},T_{3})$, each of size $n$, where $T_{1}$ and $T_{2}$ are 3rd-order tensors while $T_{3}$ is a matrix, suppose that we are interested in computing the following 3rd-order tensor: for any $(i,j,k)\in[n]^{3}$
$T^{\rm out}(i,j,k)=\sum_{a=1}^{n}\sum_{b=1}^{n}\sum_{c=1}^{n}T_{1}(i,a,b)T_{2}(a,c,j)T_{3}(b,k).$
$T^{\rm out}(i,j,k)$ is in fact a partial trace, a common operation taken in quantum mechanics. The corresponding Einsum notation is ${\mathcal{N}}=({\mathcal{A}};B)$ with ${\mathcal{A}}=(A_{1}=(1,2,3),A_{2}=(2,4,5),A_{3}=(3,6))$ with $B=(1,5,6)$ and the index set $I=[6]$. $T^{\rm out}(\alpha)$ for any $\alpha=(i,j,k)\in[n]^{3}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ because by definition:
$\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ $\equiv\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])T_{3}(\alpha^{\prime}[A_{3}])$
$=\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
$=\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}\sum_{\alpha^{\prime}[3]=1}^{n}\sum_{\alpha^{\prime}[4]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2],\alpha^{\prime}[3])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[4],\alpha[5])T_{3}(\alpha^{\prime}[3],\alpha[6]),$
so the claim holds by identifying $\alpha^{\prime}[2]=a$, $\alpha^{\prime}[3]=b$, and $\alpha^{\prime}[4]=c$.
Example 3.
The last example that we consider will also be used in Section 3 at various stages. Given three second-order tensors (matrices) ${\mathcal{T}}=(T_{1},T_{2},T_{3})$, suppose we are interested in computing a scalar $T^{\rm out}$:
$T^{\rm out}=\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}\sum_{l=1}^{n}T_{1}(i,j)T_{2}(j,k)T_{3}(k,l).$
The corresponding Einsum notation is ${\mathcal{A}}=(A_{1}=(1,2),A_{2}=(2,3),A_{3}=(3,4))$ with $B=\emptyset$ and the index set $I=[4]$. $T^{\rm out}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ because by definition:
$\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ $\equiv\sum_{\alpha\in[n]^{4}}\prod_{k=1}^{3}T_{k}(\alpha[A_{k}])=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[A_{1}])T_{2}(\alpha[A_{2}])T_{3}(\alpha[A_{3}])$
$=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[1,2])T_{2}(\alpha[2,3])T_{3}(\alpha[3,4])$
$=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[1],\alpha[2])T_{2}(\alpha[2],\alpha[3])T_{3}(\alpha[3],\alpha[4]),$
so the claim holds by identifying $\alpha[1]=i,\alpha[2]=j,\alpha[3]=k,\alpha[4]=l$.
Remark 2.
The Einsum operation has been implemented in highly optimized Python libraries such as numpy.einsum and pytorch.einsum. These efficient algorithmic subroutines, enabling efficient Einsum operation in practice, serve as the building blocks for our algorithm. Notably, developed for training large overparameterized neural networks in tensor forms, pytorch.einsum is also designed to leverage GPU computing to further enhance the scalability and speed of the Einsum operation, as we will demonstrate using an example from statistics in Section 4.1.
Remark 3.
The Einsum operation is frequently employed in scientific computing problems. For example, the Einsum operation serves as a basic algorithmic subroutine for computation over tensor (hyper)networks, a useful mathematical language adopted by the quantum computing and quantum information community (markov2008simulating; stoudenmire2018learning). Although we do not directly adopt the tensor network formulation, our strategy is quite similar in that we also establish a correspondence between the Einsum operation and operations on graphs – see Section 3.2 for more details.
2.3 Useful Concepts from Graph Theory
In this section, we recall three key concepts from graph theory that are central to our complexity-theoretic characterization: vertex elimination and treewidth, which captures the dominant computational complexity; and quotient graph, which is related to the decomposition of a $U$-statistic into a linear combination of $V$-statistics. Here we adopt the terminologies from bodlaender2010treewidth.
Definition 6 (Vertex Elimination).
Let $G=(V,E)$ be a graph and $v\in V$. We define the operation of eliminating $v$ from $V$ as first connecting $N_{G}(v)$ (the neighbor of $v$ w.r.t. $G$) by edges to form a clique and then removing the vertex $v$ and all edges incident to $v$. The resulting graph is denoted by $\mathsf{elim}_{v}(G)$.
Definition 7 (Treewidth).
Let $G=(V,E)$ be a graph with $|V|=n$. For $\sigma\in\mathsf{Perm}(V)$, $(G_{i}^{\sigma})_{i=1}^{n}$ is a sequence of graphs recursively defined as $G_{i}^{\sigma}=\mathsf{elim}_{\sigma[i]}(G_{i-1}^{\sigma})$ with $G_{0}^{\sigma}=G$. Then the treewidth of the graph $G$ is defined as
$\mathsf{tw}(G)\coloneqq\min_{\sigma\in\mathsf{Perm}(V)}\max_{i\in[n]}\deg(\sigma[i],G_{i-1}^{\sigma}).$
Definition 8 (Quotient Graph).
Let $G=(V,E)$ be a graph. A quotient graph associated with $G$ w.r.t. a partition $\pi$ of $V$ is defined as $G/\pi\coloneqq(\pi,E_{\pi})$, where
$E_{\pi}=\left\{\{B_{i},B_{j}\}\subseteq\pi\mid\exists\,v_{1}\in B_{i},v_{2}\in B_{j},\text{ such that }\{v_{1},v_{2}\}\in E\right\}.$
3 A New Algorithm of Exactly Computing $U$-Statistics
We are ready to present our new algorithm for exactly computing $U$-statistics, along with an analysis of the time complexity. Our algorithm has been implemented in an open source Python package u-stats.
We begin by introducing the multiplicative-decomposable (MD) structure of the UU/$V$-statistic kernel in Section 3.1. We then describe our algorithm and the complexity analysis for $V$-statistics in Section 3.2, followed by those for $U$-statistics in Section 3.3.
3.1 The Multiplicative-Decomposable Kernels
[…]
Definition 9 (Multiplicative Decomposition of a Function).
Let $\mathbb{X}$} be a non-empty set, $m$ a positive integer, and $h:{\mathbb{X}}^{m}\to{\mathbb{R}}$. Suppose ${\mathcal{A}}\coloneqq(A_{k})_{k=1}^{K}\in[m]^{**}$ satisfies $\bigcup_{k=1}^{K}\mathsf{set}(A_{k})=[m]$, and let ${\mathcal{H}}\coloneqq(h_{k})_{k=1}^{K}$ be a tuple of functions where each $h_{k}$ is defined on ${\mathbb{X}}^{|A_{k}|}$. ${\mathcal{H}}\times{\mathcal{A}}$, comprised of ordered pairs of the form $(h_{k},A_{k})$ for $k\in[K]$, is called a multiplicative decomposition of the function $h$, if for any $\bm{x}\in{\mathbb{X}}^{m}$,
$h(\bm{x})=\prod_{k=1}^{K}h_{k}\big{(}\bm{x}[A_{k}]\big{)}.$
In the sequel, we refer to the function tuple $\mathcal{H}$} as the kernel components and the index tuple $\mathcal{A}$} as the decomposition signature. For convenience, we say that a kernel $h$ permits a (nontrivial) MD structure ${\mathcal{H}}\times{\mathcal{A}}$ when $K>1$.
$U$-statistic kernels with MD structures are pervasive in statistical applications. We will provide three examples in Section 4.
To utilize Einsum operation to compute UU/$V$-statistics, we need to store pieces of the underlying statistics in tensor format. We term this step as tensorization. Let $h$ be the kernel of some UU/$V$-statistic permitting a MD structure ${\mathcal{H}}\times{\mathcal{A}}$, where ${\mathcal{H}}\coloneqq(h_{k})_{k=1}^{K}$ and ${\mathcal{A}}\coloneqq(A_{k})^{K}_{k=1}$. Given a sample $\bm{X}\in{\mathbb{X}}^{n}$, the tensorization step constructs a tuple of tensors ${\mathcal{T}}\coloneqq(T_{k})_{k=1}^{K}$ as follows:
$T_{k}(\alpha_{k})=h_{k}(\bm{X}[\alpha_{k}]),\quad\forall\,\alpha_{k}\in[n]^{|A_{k}|},\,\forall\,k\in[K].$
Then following Definition 2 and Definition 9, we have
$\mathbb{V}(h,\bm{X})=\sum_{\alpha\in[n]^{m}}h(\bm{X}[\alpha])\equiv\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}h_{k}(\bm{X}[\alpha[A_{k}]])\equiv\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha[A_{k}]),$ (1)
which implies that the $V$-statistic $\mathbb{V}(h,\bm{X})$ can be equivalently expressed by using the tuple of tensors $\mathcal{T}$} and the decomposition signature $\mathcal{A}$}. As will be seen in Section 3.3, we can also use $\mathcal{T}$} and $\mathcal{A}$} to represent $U$-statistics. Hence $\mathbb{V}(h,\bm{X})$ and $\mathbb{U}(h,\bm{X})$ can be equivalently denoted as $\mathbb{U}({\mathcal{T}},{\mathcal{A}})$ and $\mathbb{V}({\mathcal{T}},{\mathcal{A}})$, respectively. We summarize the tensorization step in Algorithm 1 below, which constitutes the first step of our new algorithm for exactly computing a $V$-statistic.
Algorithm 1 Tensorization of Kernel Function
1:
2:A kernel $h$ with decomposition components ${\mathcal{H}}\coloneqq(h_{k}:{\mathbb{X}}^{m_{k}}\to{\mathbb{R}})_{k=1}^{K}$
3:A sample $\bm{X}=(X_{1},\dots,X_{n})\in{\mathbb{X}}^{n}$ from a non-empty set $\mathbb{X}$}
4:A tuple of tensors ${\mathcal{T}}\coloneqq(T_{k}:[n]^{m_{k}}\to{\mathbb{R}})_{k=1}^{K}$, where $T_{k}$ is the tensorized version of $h_{k}$
5:function Tensorization(${\mathcal{H}},\bm{X}$)
6:  for $k\in[K]$ do $\triangleright$ Iterate over each factor of kernel $h$
7:   for $\alpha\in[n]^{m_{k}}$ do $\triangleright$ Iterate over all $m_{k}$-dimensional indices
8:     $T_{k}(\alpha)\leftarrow h_{k}(\bm{X}[\alpha])$ $\triangleright$ Evaluate $h_{k}$ on $\bm{X}[\alpha]$
9:   end for
10:  end for
11:  return ${\mathcal{T}}\coloneqq(T_{k})_{k=1}^{K}$
12:end function
The MD structure plays an essential role in reducing the complexity of exactly computing $V$-statistics. Without a (nontrivial) MD structure, the time complexity of computing an $m$-th order $V$-statistic is generally $O(n^{m})$ as we need to evaluate a function of $m$ arguments on all tuples of $n]^{m}$. But if a particular MD structure holds, the time complexity may scale with $O(n^{m_{\mathrm{max}}})$ where $m_{\mathrm{max}}\coloneqq\max_{k\in[K]}|A_{k}|$ since we just need to evaluate each $h_{k}$ on all tuples of $n]^{|A_{k}|}$ for $k\in[K]$. As $m_{\mathrm{max}}$ can be smaller than $m$, it is possible to improve the time complexity dramatically in practice. More details on the complexity analysis and intuition can be found in Section 3.2 below.
Example 3(a).
Recalling Example 3. The Einsum operation can be seen as a $V$-statistic whose kernel function $h$ has a MD structure ${\mathcal{H}}\times{\mathcal{A}}$ with ${\mathcal{H}}=(h_{1},h_{2},h_{3})$, signature ${\mathcal{A}}=((1,2),(2,3),(3,4))$ and $n$ sample points $\bm{X}\in{\mathbb{X}}^{n}$. Tensorization of $\mathcal{H}$} outputs tensors ${\mathcal{T}}=(T_{1},T_{2},T_{3})$ such that:
$T_{i}(j,k)=h_{i}(\bm{X}_{j},\bm{X}_{k}),\,(j,k)\in[n]^{2},\,i=1,2,3.$
The $V$-statistic and $U$-statistic with kernel $h$ are respectively:
$\mathbb{V}(h,\bm{X})=\mathbb{V}({\mathcal{T}},{\mathcal{A}})=\textsf{Einsum}{}({\mathcal{T}},{\mathcal{A}}),\quad\mathbb{U}(h,\bm{X})=\mathbb{U}({\mathcal{T}},{\mathcal{A}}).$
3.2 The Algorithm and Time Complexity of Computing $V$-Statistics
In this section, we show that $V$-statistics can be computed via the Einsum operation. Let $\mathbb{V}(h,\bm{X})$ be an $m$-th order $V$-statistic with kernel $h$ and sample $\bm{X}$ of size $n$. Suppose that the kernel $h$ satisfies MD structure $({\mathcal{H}}\coloneqq(h_{k})_{k=1}^{K})\times({\mathcal{A}}\coloneqq(A_{k})_{k=1}^{K})$. As we have seen, $\mathbb{V}(h,\bm{X})$ has an equivalent tensorized form $\mathbb{V}({\mathcal{T}},{\mathcal{A}})$, where $\mathcal{T}$} is the output of Algorithm 1.
By definition, $\mathcal{A}$} constitutes an Einsum notation and $\mathcal{T}$} can be represented by $\mathcal{A}$} (see Definition 5), then we have
$\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})=\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha[A_{k}]).$
Comparing the above equation with (1), we immediately have
$\mathbb{V}(h,\bm{X})=\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}),$
i.e. computing a $V$-statistic is the same as performing an Einsum operation. Algorithm 2 describes how $V$-statistics can be computed by the Einsum operation.
Algorithm 2 Tensorization Algorithm of $V$-statistics
1:
2:A kernel function $h:{\mathbb{X}}^{m}\to{\mathbb{R}}$ permitting a MD structure ${\mathcal{H}}\times{\mathcal{A}}$
3:A sample $\bm{X}\in{\mathbb{X}}^{n}$ from a non-empty set $\mathbb{X}$}
4:$V$-statistic $\mathbb{V}(h,\bm{X})$
5:function Vstats(${\mathcal{H}},{\mathcal{A}};\bm{X}$)
6:  $\mathcal{T}$$\leftarrow$ Tensorization(${\mathcal{H}},\bm{X}$) $\triangleright$ see Algorithm 1
7:  $v\leftarrow\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ [^1] In our implementation of u-stats, we directly call numpy.einsum or pytorch.einsum, with the summation ordering optimized by opt-einsum.
8:  return $v$
9:end function
The time complexity of computing a $V$-statistic can obviously be characterized by that of the Einsum operation. The essential idea is to establish a correspondence between the Einsum operation and the vertex elimination operation (Definition 6) over a graph associated with the Einsum notation. The time complexity can in turn be characterized by the treewidth (Definition 7) of that graph. Similar idea has also been exploited in markov2008simulating, in which the time complexity of as the so-called tensor contraction operation (a special case of the Einsum operation) is also characterized via the treewidth of the underlying tensor network.
To facilitate presentation, we first introduce the notion of “decomposition graph” associated with a given Einsum notation, which captures all the information needed to establish the time complexity of computing a particular $V$-statistic.
Definition 10 (The Decomposition Graph Associated with an Einsum Notation).
Let $\mathcal{A}$} be an Einsum notation with no output. The decomposition graph associated with $\mathcal{A}$} (or equivalently the decomposition signature $\mathcal{A}$}) is a graph $G_{{\mathcal{A}}}=(V_{{\mathcal{A}}},E_{{\mathcal{A}}})$ where
  • $V_{{\mathcal{A}}}=\mathsf{Indices}({\mathcal{A}})$,
  • $E_{{\mathcal{A}}}=\big{\{}\{i,j\}\subseteq V_{{\mathcal{A}}}\mid\exists\,A\in{\mathcal{A}};\text{ such that }i,j\in\mathsf{set}(A)\big{\}}$.
We use several concrete examples to illustrate the above definition.
Example 4 (Correspondence between the Einsum notation and the decomposition graph).
Figure 1 below illustrates several examples of decomposition graphs associated with different decomposition signatures. The underlying idea is to treat each index in $\mathcal{A}$} as a vertex and to add an edge between two vertices whenever the corresponding indices co-occur in the same tuple. Repeated tuples do not create multiple edges (see Figures 1(b) and 1(e)). Moreover, different decomposition signatures can yield the same decomposition graph (see Figures 1(f) and 1(h)).
1234
(a) ${\mathcal{A}}=((1,2),(2,3),(3,4))$
123
(b) ${\mathcal{A}}=((1,2),(2,1),(1,3))$
123
(c) ${\mathcal{A}}=((1,2),(2,3),(3,1))$
123
(d) ${\mathcal{A}}=((1,2),(2,3),(3,2))$
12
(e) ${\mathcal{A}}=((1,2),(2,1),(1,2))$
1234
(f) ${\mathcal{A}}=((1,2),(1,3),(1,4),$
$(2,3),(2,4),(3,4))$
1234
(g) ${\mathcal{A}}=((1,2),(3,4))$
1234
(h) ${\mathcal{A}}=((1,2,3),(1,3,4),$
$(1,2,4),(2,3,4))$
Figure 1: Examples of decomposition graphs associated with different decomposition signatures.
We are ready to present the following “optimistic” estimate of the time complexity of the Einsum operation. The proof of the proposition below is deferred to Appendix A.
Proposition 1.
Let $\mathcal{A}$} be an Einsum notation with no output, and $G_{\mathcal{A}}$ be the corresponding decomposition graph of $\mathcal{A}$}. Then there exists an algorithm with a particular Einsum ordering such that for any set of tensors $\mathcal{T}$} that can be represented by the Einsum notation $\mathcal{A}$}, the time complexity of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(|{\mathcal{A}}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$.
Remark 4 (The Einsum ordering in practice).
We say that the time complexity given in the above proposition is only an optimistic estimate, because although it can be achieved with the optimal Einsum ordering corresponding to the treewidth (Definition 7), this particular ordering is usually difficult to find in practice. arnborg1987complexity have shown that finding the ordering that gives rise to the treewidth is an NP-complete problem. While it is ideal to evaluate $\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ in Algorithm 2 using this optimal ordering, we decide to directly plug in the Python library opt-einsum (daniel2018opt) to find the ordering in practice, because such a preexisting library has been optimized and stabilized over the years. opt-einsum offers both heuristic strategies such as greedy algorithms or exhaustive search such as dynamic programming, but the latter can be time-consuming when dealing with large-scale problems. As mentioned in the Introduction, dynamic programming is also used in zhang2025adaptive for a special class of $U$-statistics. Since the ordering only depends on the MD structure of the kernel $h$, for commonly encountered kernels, one could in principle find the optimal ordering via exhaustive search without accessing the data sample and re-use this ordering whenever the same type of kernel appears. Finally, we briefly remark that opt-einsum uses a different scheme to describe the summation ordering from the one used in this paper. We will not describe such a difference in any detail as doing so will necessarily introduce extra notations and definitions that are only tangential to the main theme of the paper.
With Proposition 1, we immediately obtain the following optimistic estimate of the time complexity of exactly computing a $V$-statistic.
Corollary 1.
Let $\mathbb{X}$} be a non-empty set, $m,n$ be positive integers and $\bm{X}\in{\mathbb{X}}^{n}$. Suppose that $h:{\mathbb{X}}^{m}\rightarrow{\mathbb{R}}$ be a kernel that permits a MD structure ${\mathcal{H}}\times{\mathcal{A}}$ and for each $h_{k}\in{\mathcal{H}}=(h_{k})_{k=1}^{K}$, the time complexity of evaluating $h_{k}$ is of $O(1)$ and does not depend on $n$. Let $\mathcal{T}$} be the tuple of tensors after applying the tensorization algorithm (Algorithm 1) to $\mathcal{H}$} and $G_{\mathcal{A}}$ be the decomposition graph of $\mathcal{A}$}. Then there exists an algorithm such that its time complexity of computing $\mathbb{V}(h,\bm{X})$ exactly is $O(|{\mathcal{A}}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$.
3.3 The Algorithm and Time Complexity of Computing $U$-Statistics
As discussed in Section 2.1, unlike $V$-statistics, $U$-statistics cannot be directly computed via the Einsum operation. Therefore, we first decompose a given $U$-statistic into a linear combination of $V$-statistics. Each of the $V$-statistic terms can then be efficiently computed through Einsum, and finally recombined to recover the original $U$-statistic.
To present how to decompose a $U$-statistic into $V$-statistics, several notions need to be introduced. Recall that $\Pi_{m}$ denotes all partitions of $[m]$.
Definition 11 (V-set).
Let $m,n$ be positive integers satisfying $m\leq n$ and $\pi$ be a partition of the set $[m]$. A V-set of size $n$ w.r.t. partition $\pi$ is defined as
${\mathcal{V}}(n,\pi)\coloneqq\left\{\bar{s}_{m}\in[n]^{m}\mid\text{where for any $i,j\in[m]$},i,j\in Q\text{ for some $Q\in\pi$}\Longrightarrow s_{i}=s_{j}\right\}.$
If $\pi$ happens to be the finest partition of $[m]$, i.e. $\pi=\{\{1\},\{2\},\cdots,\{m\}\}$, the corresponding V-set is denoted by ${\mathcal{V}}(n,m)$ and in fact, ${\mathcal{V}}(n,m)=[n]^{m}$. The cardinality of the given partition $\pi$ is referred to as the order of ${\mathcal{V}}(n,\pi)$.
Definition 12 (Restricted $V$-statistics).
Let $\mathbb{X}$} be a non-empty set, $m,n$ be positive integers, $\pi$ be a partition of set $[m]$ and $h:{\mathbb{X}}^{m}\rightarrow{\mathbb{R}}$ be a function defined on ${\mathbb{X}}^{m}$. For any tuple $\bm{X}\in{\mathbb{X}}^{*}$, the $V$-statistic with kernel $h$ restricted by partition $\pi$ takes the following form
$\mathbb{V}[\pi](h,\bm{X})\coloneqq\sum_{\alpha\in{\mathcal{V}}(n,\pi)}h(\bm{X}[\alpha]).$
Having introduced the above concepts, we state the following lemma that describes how to decompose a $U$-statistic into a linear combination of $V$-statistics.
Lemma 1 (Decomposition of $U$-statistics into $V$-statistics).
Let $\mathbb{X}$} be a non-empty set and $h:{\mathbb{X}}^{m}\rightarrow{\mathbb{R}}$ be a kernel function. Then for any tuple $\bm{X}\in{\mathbb{X}}^{n}$ with $n\geq m$,
$\mathbb{U}(h,\bm{X})=\sum_{\pi\in\Pi_{m}}\mu_{\pi}\mathbb{V}[\pi](h,\bm{X}),$
where
$\mu_{\pi}=(-1)^{(m-|\pi|)}\prod_{C\in\pi}(|C|-1)!.$
The proof of the above lemma can be found in Appendix B and leverages Möbius inversion from combinatorics. We now discuss how to extend the tensorization algorithm from $V$-statistics (Algorithm 1) to $U$-statistics. Let $\mathbb{U}(h,\bm{X})$ be an $m$-th order $U$-statistic with kernel function $h$ and sample $\bm{X}\in{\mathbb{X}}^{n}$, where $h$ permits a MD structure ${\mathcal{H}}\times{\mathcal{A}}$. On the surface, it seems that we have to apply Algorithm 2 to each resulting $V$-statistics in the decomposition given in Lemma 1. But in fact, we only need to apply Algorithm 2 once since all the resulting $V$-statistics can be tensorized by the same tensor tuple ${\mathcal{T}}=\textsc{Tensorizaion}({\mathcal{H}},\bm{X})$.
Given any $\pi\in\Pi_{m}$, we can also compute the restricted $V$-statistic $\mathbb{V}[\pi](h,\bm{X})$ by turning it into a new Einsum notation ${\mathcal{A}}_{\pi}$ constructed from $\mathcal{A}$} based on $\pi$. ${\mathcal{A}}_{\pi}$ simply identifies all the indices that happen to belong to the same subset of the partition $\pi$. The procedure is presented in Algorithm 3. We then have the following identity that relates the restricted $V$-statistic by the partition $\pi$ and the Einsum operation under the induced decomposition signature ${\mathcal{A}}_{\pi}$:
$\mathbb{V}[\pi](h,\bm{X})=\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}_{\pi}).$ (2)
One can immediately realize that (2) holds because, by Definitions 11 and 12, $\mathbb{V}[\pi](h,\bm{X})$ is a summation over all tuples of $n]^{m}$, under the restriction that indices belonging to the same subset of the partition $\pi$ need to be identified.
Algorithm 3 Induced Einsum Notation via Partition
1:
2:An Einsum notation ${\mathcal{A}}=(A_{k})_{k=1}^{K}$ with index set $[m]$
3:A partition $\pi\in\Pi_{m}$ of size $J$
4:The Einsum notation ${\mathcal{A}}_{\pi}$ induced by $\pi$ on $\mathcal{A}$}
5:function InducedNotation(${\mathcal{A}};\pi$)
6:  Let $(Q_{j})_{j=1}^{J}$ be a permutation of $\pi$ $\triangleright$ Label subsets by $j\in[J]$
7:  Define $\psi\colon[m]\to[J]$ as the subset assignment map:
8:  function $\psi$(ii)
9:   return $j$ if $i\in Q_{j}$ $\triangleright$ Map all indices in $Q_{j}$ to $j$
10:  end function
11:  Initialize ${\mathcal{A}}_{\pi}\leftarrow{\mathcal{A}}$
12:  for $k\in[K]$ do
13:   for $i\in[|A_{k}|]$ do
14:     ${\mathcal{A}}_{\pi}[k][i]\leftarrow\psi(A_{k}[i])$ $\triangleright$ Replace index with its subset label
15:   end for
16:  end for
17:  return ${\mathcal{A}}_{\pi}$
18:end function
According to Lemma 1, when computing an $m$-th order $U$-statistic, the number of $V$-statistics that we need to compute is the same as the total number of partitions of $[m]$, or equivalently the Bell number $\mathsf{B}_{m}$. Bell number $\mathsf{B}_{m}$ grows super-exponentially with $m$. As a result, when the order $m$ increases, the number of $V$-statistics that needs to be computed grows rapidly. Fortunately, we discover a sparsification trick to reduce the number of $V$-statistics in the decomposition when possible. To this end, we first introduce the following sparsified tensor list $\widetilde{{\mathcal{T}}}$, which gives the same $U$-statistic as the original tensor list $\mathcal{T}$}.
Lemma 2.
Let $h$ be a kernel function that permits a MD structure $({\mathcal{H}}=(h_{k})_{k=1}^{K})\times({\mathcal{A}}=(A_{k})_{k=1}^{K})$. A tensor tuple ${\mathcal{T}}=(T_{k})_{k=1}^{K}$ is constructed from $\mathcal{H}$} as in Algorithm 1. Let $\widetilde{{\mathcal{T}}}\coloneqq(\widetilde{T}_{k})_{k=1}^{K}$ be constructed as follows:
$\begin{split}\widetilde{T}_{k}(\alpha)=\begin{cases}h_{k}(\bm{X}[\alpha])&\text{if }\forall\,i,j\in[|A_{k}|]:i\neq j,\,\alpha[i]\neq\alpha[j]\\ 0&\text{otherwise},\end{cases}\,\forall\,\alpha\in[n]^{|A_{k}|},\,\forall\,k\in[K].\end{split}$ (3)
Then it holds that
$\mathbb{U}({\mathcal{T}},{\mathcal{A}})=\mathbb{U}(\widetilde{{\mathcal{T}}},{\mathcal{A}}).$
The proof of Lemma 2 is straightforward as none of the indices $\alpha$ such that $\widetilde{T}_{k}(\alpha)\neq T_{k}(\alpha)$ (when the second case of (3) holds) contributes to the value of the underlying $U$-statistic, because by definition $U$-statistic sums over distinct indices. Lemma 2 suggests that some Einsum operations can be skipped when computing $U$-statistics.
Remark 5.
After applying the sparsification trick, the only remaining terms required to compute the $U$-statistic $\mathbb{U}(\widetilde{{\mathcal{T}}},{\mathcal{A}})$ are the restricted $V$-statistics $\mathbb{V}[\pi](\widetilde{{\mathcal{T}}},{\mathcal{A}})$ associated with partitions in the set
$\Pi_{m}^{\mathcal{A}}\coloneqq\left\{\pi\in\Pi_{m}\,\middle|\,\forall\,Q\in\pi,\ \forall A\in{\mathcal{A}},\ |Q\cap\mathsf{set}(A)|<2\right\}.$ (4)
That is, only the Einsum operations with notation ${\mathcal{A}}_{\pi}$, for which any $Q\in\pi$ contains $\leq 1$ index from $\mathsf{set}(A)$ for any $A\in{\mathcal{A}}$, need to be evaluated when computing $\mathbb{U}(h,{\mathcal{A}})$. To be more concrete, under a given partition $\pi$, if some indices are in the same component of the decomposition signature $\mathcal{A}$}, the corresponding Einsum operation can be skipped because the result must be zero. Example 3(b) will illustrate this sparsification trick using a toy example. In Table 1 of Section 4.1, we use higher-order influence functions (HOIFs) mentioned in the beginning of this paper as a realistic example to also exhibit the effectiveness of this sparsification trick in reducing the number of $V$-statistics needed to compute.
With the sparsification trick, the final algorithm of computing $U$-statistics is displayed in Algorithm 4 below.
Algorithm 4 Computation of $U$-statistics via Tensorization
1:
2:A kernel function $h:{\mathbb{X}}^{m}\to{\mathbb{R}}$ satisfying MD ${\mathcal{T}}\times{\mathcal{A}}$
3:A sample $\bm{X}=(X_{1},\dots,X_{n})\in{\mathbb{X}}^{n}$
4:The $U$-statistic $\mathbb{U}(h,\bm{X})$
5:function Ustats(${\mathcal{T}},{\mathcal{A}};\bm{X}$)
6:  Let $\widetilde{{\mathcal{T}}}$ be constructed as in Lemma 2
7:  Initialize $u\leftarrow 0$
8:  for $\pi\in\Pi_{m}$ do
9:   if $\forall Q\in\pi,\forall A\in{\mathcal{A}};|Q\cap\mathsf{set}(A)|<2$ then
10:     $\mu_{\pi}\leftarrow(-1)^{m-|\pi|}\prod_{C\in\pi}(|C|-1)!$ $\triangleright$ See Lemma 1
11:     ${\mathcal{A}}_{\pi}\leftarrow$ InducedNotation(${\mathcal{A}};\pi$) $\triangleright$ See Algorithm 3
12:     $v_{\pi}\leftarrow\mathsf{Einsum}(\widetilde{{\mathcal{T}}},{\mathcal{A}}_{\pi})$ (see footnote 1)
13:     $u\leftarrow u+\mu_{\pi}v_{\pi}$
14:   end if
15:  end for
16:  return $u$
17:end function
Example 3(b).
To illustrate how the above sparsification trick manifests in concrete examples, we consider a $4$-th order $U$-statistics, with a kernel function $h$ has a decomposition signature ${\mathcal{A}}=((1,2),(2,3),(3,4))$ as in Example 3(a). This $U$-statistic can be decomposed into a linear combination of $V$-statistics as follows:
$\ \mathbb{U}(\widetilde{{\mathcal{T}}},{\mathcal{A}})=\mathbb{V}[\{1\}\{2\}\{3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})-\cancel{\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}-\mathbb{V}[\{1,3\}\{2\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})$
$-\mathbb{V}[\{1,4\}\{2\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})-\cancel{\mathbb{V}[\{2,3\}\{1\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}-\mathbb{V}[\{2,4\}\{1\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})-\cancel{\mathbb{V}[\{3,4\}\{1\}\{2\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
$+\cancel{\mathbb{V}[\{1,2\}\{3,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\mathbb{V}[\{1,3\}\{2,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})+\cancel{\mathbb{V}[\{1,4\}\{2,3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
$+\cancel{2\mathbb{V}[\{1,2,3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\cancel{2\mathbb{V}[\{1,2,4\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\cancel{2\mathbb{V}[\{1,3,4\}\{2\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
$+\cancel{2\mathbb{V}[\{2,3,4\}\{1\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}-\cancel{\mathbb{V}[\{1,2,3,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}.$
As shown in the above display, there were originally 15 $V$-statistics in total, whereas only 5 $V$-statistics are left after applying the sparsification trick! Taking the second term $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})$ on the RHS of the above display as an example: $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})\equiv 0$ because the intersection between $Q=\{1,2\}$ and $A_{1}=(1,2)$ is 2, violating the criterion described in (4).
Armed with the above preparations, we can bound the time complexity of exactly computing $U$-statistics using Proposition 1. Given a partition $\pi$ and the decomposition signature $\mathcal{A}$}, the decomposition graph corresponding to induced Einsum notation ${\mathcal{A}}_{\pi}$ can be characterized by the quotient graph $G_{\mathcal{A}}/\pi$ (see Definition 8) of the decomposition graph $G_{\mathcal{A}}$ induced by $\pi$.
Corollary 2.
Let $G_{\mathcal{A}}$ be the decomposition graph of $\mathcal{A}$}. Under the same conditions of Corollary 1, there exists an algorithm such that the time complexity of computing the $U$-statistic $\mathbb{U}(h,\bm{X})$ is $O(|{\mathcal{A}}|\sum_{l=1}^{M}N_{l}n^{l+1})$, where
$M$ $\coloneqq\max\{\mathsf{tw}(G_{\mathcal{A}}/\pi)\mid\pi\in\Pi_{m}^{\mathcal{A}}\},$
$N_{l}$ $\coloneqq|\{\pi\in\Pi_{m}^{{\mathcal{A}}}\mid\mathsf{tw}(G_{{\mathcal{A}}}/\pi)=l\}|,\;\forall\,l\in[M].$
and $\Pi_{m}^{\mathcal{A}}$ is defined in (4).
Remark 6.
In the above result, quantities $N_{l}$ and $M$ are introduced. $N_{l}$ is the number of partitions in $\Pi_{m}^{{\mathcal{A}}}$ such that the treewidth of the decomposition graph $G_{{\mathcal{A}}}$ quotient by $\pi$ is $l$, is defined. Although abstract, $N_{l}$ and $M$ can in principle be computed for concrete examples. For instance, in Section 4.1, we report the values of $M$ for HOIF estimators between orders $2$ to $12$; see the last column of Table 1. At this moment, we are not yet aware of a good strategy to count $N_{l}$, except a loose upper bound by the total number of $V$-statistics in the decomposition; see the third column of Table 1.
Remark 7.
In this paper, we have used the treewidth of the decomposition graph (or its quotient as in Corollary 2) to characterize the optimistic time-complexity of exactly computing UU/$V$-statistics, and the Einsum operation. Although treewidth has been a widely used concept for establishing the time-complexity of algorithms on graphs (dong2021nearly), it is in general nontrivial to compute or derive a tight bound. To our knowledge, Theorem 4.7 in (kneis2009bound) proved a bound for treewdith linear in the edge size. We do not know if this bound can be further improved. We will discuss this as an open problem at the end of this paper (Section 5).
Before moving on to applications of our algorithm in the next section, we summarize the steps of exactly computing a $U$-statistic using our proposed algorithm. Given a $U$-statistic kernel $h$ and a sample $\bm{X}$:
  1. (1)
    tensorize $h$ into $\mathcal{T}$} (Algorithm 1) and construct $\widetilde{{\mathcal{T}}}$ as in Lemma 2;
  2. (2)
    for each possible partition $\pi$ in $\Pi_{m}$, obtain the corresponding Einsum notation ${\mathcal{A}}_{\pi}$ (Algorithm 3) and evaluate the corresponding $V$-statistic using the Einsum operation;
  3. (3)
    add all evaluated $V$-statistics together sequentially as in Algorithm 4.
4 Examples and Numerical Results
In this section, we evaluate the performance of our new algorithm for computing UU/$V$-statistics exactly in three concrete statistical applications, including computing higher-order influence functions (Section 4.1), computing motif counts in random graphs (Section 4.2) and computing distance covariance measures (Section 4.3).
As mentioned, we have made our algorithm available for public use as a Python package u-stats, which is available to download from either its GitHub repository (clicking u-stats) or the Python package index PyPi by typing the following on the command line:
pip install u-stats
We also refer interested readers to the GitHub link for how to implement u-stats in practice.
4.1 Application 1: Higher-Order Influence Functions
As mentioned in the Introduction, Higher-Order Influence Functions (HOIFs) (robins2008higher; robins2016technical) can help construct rate-optimal estimators for important parameters such as the treatment-specific mean under ignorability under certain complexity-reducing assumptions on the data generating mechanism. However, estimators constructed via HOIFs are high-order $U$-statistics, which are time-consuming to compute in practice. Fortunately, the HOIF estimator of ATE is a sum of $U$-statistics, each of which has kernel satisfying the MD structure as in Definition 9. In the setting of an unconfounded observational study, we often have access to a sample tuple $\bm{X}\equiv(X_{i}\equiv(Z_{i},A_{i},Y_{i}))_{i=1}^{n}$ of size $n$, where $Z\in{\mathbb{R}}^{d}$ denotes the baseline covariates, $A\in\{0,1\}$ denotes a binary treatment variable and $Y$ denotes the outcome of interest. According to robins2008higher, with a $k$-dimensional dictionary $\phi:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{k}$, an $m$-th order influence function estimator of the treatment-specific mean reads as
$\hat{\tau}_{m}=\sum_{j=2}^{m}(-1)^{j}\mathbb{U}(h^{(j)},\bm{X}),$
where for every $j=2,\cdots,m$, given $\bar{i}_{j}\in[n]^{j}$,
$h^{(j)}(\bm{X}[\bar{i}_{j}])=A_{i_{1}}\phi(Z_{i_{1}})^{\top}\left\{\prod_{s=3}^{j}\left(A_{i_{s}}\phi(Z_{i_{s}})A_{i_{s}}\phi(Z_{i_{s}})^{\top}-\mathbb{I}\right)\right\}\phi(Z_{i_{2}})Y_{i_{2}}$
and we assume for simplicity that the population mean of $A\phi(Z)\phi(Z)^{\top}$ is $\mathbb{I}$.
$\hat{\tau}_{m}$ can be rewritten as a sum over $j$-th order $U$-statistic with kernel $h^{(j)}$ satisfying the MD structure, for $j=2,\cdots,m$. To see this, without loss of generality we consider $j=m$ and rewrite $\mathbb{U}(h^{(m)},\bm{X})$ as
$\mathbb{U}(h^{(m)},\bm{X})=\sum_{j=0}^{m-2}\binom{m-2}{j}\frac{1}{\binom{n}{j}j!}\mathbb{U}(h_{j}^{(m)},\bm{X}).$
where for any $\bar{i}_{j}\in[n]^{j}$,
$h_{j}^{(m)}(\bm{X}[\bar{i}_{j}])\coloneqq A_{i_{1}}\phi(Z_{i_{1}})^{\top}\left\{\prod_{s=3}^{j}A_{i_{s}}\phi(Z_{i_{s}})A_{i_{s}}\phi(Z_{i_{s}})^{\top}\right\}\phi(Z_{i_{2}})Y_{i_{2}}.$ (5)
Then obviously $h_{j}^{(m)}$ permits a MD structure by letting $f_{l}(X_{1},X_{2})\coloneqq A_{1}\phi(Z_{1})^{\top}A_{2}\phi(Z_{2})$ for $l<j$ and $f_{j}(X_{1},X_{2})\coloneqq A_{1}\phi(Z_{1})^{\top}\phi(Z_{2})Y_{2}$. The corresponding decomposition signature of $h_{j}^{(m)}$ is $((i,i+1))_{i=1}^{j-1}$, which is denoted by ${\mathcal{A}}_{\mathsf{HOIF},j}$, so the decomposition graph of ${\mathcal{A}}_{\mathsf{HOIF},j}$ is a chain graph with e=j1e=j-1 edges.
We also demonstrate the effectiveness of the sparsification trick using the HOIF example (see Lemma 2). The following table enlists the total number of $V$-statistics and the number of floating-point operations (FLOPs) that need to be computed before and after applying the sparsification trick when the sample size is $10^{4}$. The rightmost column of Table 1 also records the value of $M$ defined in Corollary 2. The derivation of this column can be found in Appendix C.2 (and unfortunately we do not know a more systematic way of deriving these).
Table 1: Exact number of floating-point operations (FLOPs), $V$-statistic decompositions, and intermediate tensor sizes across HOIF orders from $2$ to $12$, with and without the sparsification trick, where the Einsum ordering is found by applying opt-einsum with optimization scheme greedy. Sample size is fixed at $n=10000$.
Order ($m$)
V-stat terms
(Bell number)
V-stat terms
FLOPs
FLOPs
$M$
2 2 1 $2.00020\times 10^{8}$ $2.00000\times 10^{8}$ 11
3 5 2 $060\times 10^{12} 2.00020$$020\times 10^{12}$ 11
4 15 5 $8.00310\times 10^{12}$ $4.00150\times 10^{12}$ 22
5 52 15 $5.20133\times 10^{13}$ $60045\times 10^{13} 22$
6 203 52 $2.94069\times 10^{14}$ $1.20020\times 10^{14}$ 22
7 877 203 $1.78236\times 10^{15}$ $6.32091\times 10^{14}$ 22
8 4140 877 $31054\times 10^{17} 2.2$$23430\times 10^{17}$ 33
9 21147 4140 $5.29126\times 10^{18}$ $47959\times 10^{18} 33$
10 115975 21147 $2.07990\times 10^{21}$ $2.04048\times 10^{21}$ 33
11 678570 115975 $5.50026\times 10^{22}$ $3.72144\times 10^{22}$ 44
12 4213597 678570 $1.13401\times 10^{24}$ $6.62181\times 10^{23}$ 44
In Tables 2 and 3, we report the average runtime over 3 runs when computing the $U$-statistic $\mathbb{U}(h_{j}^{(m)},\bm{X})$ where kernel function $h_{j}^{(m)}$ is as in (5) with $m=7$ and $j\in\{2,3,\cdots,7\}$ and sample size $n\in\{1000,2000,4000,8000,10000\}$, respectively using CPU parallel computation and using a single GPU. We would like to highlight that even when $m=7$, $n=10000$, it only takes about 5 minutes using our new package u-stats, with CPU parallel computation. If one instead has access to a single GPU, the runtime even reduces to about 12 seconds, a substantial speedup! We remark in passing that all the runtime reported in Table 2 and Table 3 excludes the time of kernel evaluation and tensorization.
Table 2: Average runtime (in seconds) using CPU parallel computation. Experiments were conducted on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB, evaluated across varying sample sizes and orders of HOIF-type $U$-statistics.
Order 1000 2000 4000 8000 10000
2 0.64 0.00141 0.01298 0.03174 0.04746
3 0.00396 0.01518 0.07392 0.26849 0.45976
4 0.02321 0.07313 0.32765 2.09545 2.36766
5 0.09853 0.36239 1.71419 9.11349 14.21350
6 0.38878 1.47719 7.44444 40.22143 58.85063
7 1.91677 6.75947 34.08805 195.31414 290.54295
Table 3: Average runtime (in seconds) using a single GPU (NVIDIA RTX 4090, 24GB) with parallel computation, across varying sample sizes and orders of HOIF-type $U$-statistics.
Order 1000 2000 4000 8000 10000
2 0.00184 0.00151 0.00155 0.00238 0.00339
3 0.00172 0.00147 0.00193 0.00577 0.00745
4 0.00305 0.00353 0.00707 0.03612 0.06245
5 0.00835 0.01089 0.03456 0.19807 0.36160
6 0.02608 0.03906 0.16981 1.12072 2.13328
7 0.12031 0.19132 0.91884 6.45225 12.21350
4.2 Application 2: Motif Counts in Random Graph Models and Networks
Motif (or subgraph) counts refer to the number of occurrences of a particular subgraph type $\mathsf{r}$ within a larger graph $G$, denoted by $\mathsf{C}(\mathsf{r},G)$. These counts, also known as network moments, are crucial for understanding the structural properties of random graph models often used to model real-world networks. Interestingly, motif counts can be equivalently represented as $U$-statistics. Leveraging this observation, we show that various motif counts can be computed using a single Einsum operation. While this approach may not capture fine-grained structural information such as sparsity of the adjacency matrix, it is highly efficient in the context of dense graphs, significantly outperforming a popular software called igraph designed for computing motif counts.
Motif counts in a simple graph $G$ with $n$ vertices and the adjacency matrix $A\in\{0,1\}^{n\times n}$ can be represented as $U$-statistics of the following form:
$\mathsf{C}(\mathsf{r},G)=\frac{1}{|\mathrm{Aut}(\mathsf{r})|}\sum_{\bar{i}_{m}\in\mathsf{Perm}(n,m)}h_{\mathsf{r}}(\bar{i}_{m};A),$
where the kernel function $h_{\mathsf{r}}$ encodes the motif structure, and $|\mathrm{Aut}(\mathsf{r})|$ denotes the number of automorphisms of motif $\mathsf{r}$, which counts the number of symmetric repetitions.
To be concrete, letting $B\equiv 1-A$, the corresponding $U$-statistics for motif counts for specific $3$-vertex motifs and $4$-vertex motifs (see Figure 2 for which motifs $\mathsf{r}_{1}$ to $\mathsf{r}_{8}$ encode) read as follows:
$\mathsf{C}(\mathsf{r}_{1},G)$ $=\frac{1}{2}\sum_{\bar{i}_{3}\in\mathsf{Perm}(n,3)}A_{i_{1},i_{2}}A_{i_{2},i_{3}}B_{i_{3},i_{1}},$
$\mathsf{C}(\mathsf{r}_{2},G)$ $=\frac{1}{6}\sum_{\bar{i}_{3}\in\mathsf{Perm}(n,3)}A_{i_{1},i_{2}}A_{i_{2},i_{3}}A_{i_{3},i_{1}},$
$\mathsf{C}(\mathsf{r}_{3},G)$ $=\frac{1}{6}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}B_{i_{2},i_{3}}B_{i_{2},i_{4}}B_{i_{3},i_{4}},$
$\mathsf{C}(\mathsf{r}_{4},G)$ $=\frac{1}{2}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}B_{i_{1},i_{3}}B_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
$\mathsf{C}(\mathsf{r}_{5},G)$ $=\frac{1}{2}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}B_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
$\mathsf{C}(\mathsf{r}_{6},G)$ $=\frac{1}{8}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}B_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
$\mathsf{C}(\mathsf{r}_{7},G)$ $=\frac{1}{4}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}A_{i_{2},i_{4}}B_{i_{3},i_{4}},$
$\mathsf{C}(\mathsf{r}_{8},G)$ $=\frac{1}{24}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}A_{i_{2},i_{4}}A_{i_{3},i_{4}}.$
We compare our package u-stats with igraph (csardi2006igraph) on exactly computing all 3-vertex and all 4-vertex motif counts. igraph computes motif counts by using a highly optimized algorithm based on wernicke2006fanmod.
We generate random graphs using the classical Erdös-Rényi model $G(n,p)$, where each of the $\binom{n}{2}$ possible edges between $n$ vertices is present independently with probability $p$. We report the runtime comparisons in Table 4 and Table 5. All experiments were conducted on a server equipped with Intel Xeon Platinum 8358 CPUs (2.6GHz, 64 cores) and memory of 512 GB. Our implementation makes full use of all 64 CPU cores via multiprocessing, while igraph runs in a single-threaded setting, due to its current implementation limitation.
To quantify the performance, we consider two metrics:
  • Speedup: the ratio of igraph runtime to the runtime of u-stats with parallel computation;
  • Single-core speedup: a very rough estimate of per-core speedup, computed as the speedup divided by the number of cores (i.e. $64$).
As shown in Table 4 and Table 5, igraph has better runtime performance than u-stats when $p$ is small, corresponding to a sparse graph. Because u-stats is designed as a general purpose package for computing $U$-statistics, we do not further take into account the potential sparsity of the adjacency matrix. When $p$ increases (i.e., as the graph becomes denser), the runtime of igraph grows considerably even when $p$ is as low as $0.1$. Yet u-stats still maintains roughly the same runtime throughout, which is much faster than that of igraph when $p$ grows.
(a) $\mathsf{r}_{1}$: V-shape
(b) $\mathsf{r}_{2}$: Triangle
(c) $\mathsf{r}_{3}$ : 3-star
(d) $\mathsf{r}_{4}$ : Fork
(e) $\mathsf{r}_{5}$ : Tailed triangle
(f) $\mathsf{r}_{6}$: Square
(g) $\mathsf{r}_{7}$: House
(h) $\mathsf{r}_{8}$ : 4-clique
Figure 2: All non-degenerate isomorphism classes of 3-vertex and 4-vertex simple undirected graphs.
Table 4: Runtime comparison of exact all 33-vertex motif counts using u-stats and igraph on Erdös-Rényi graphs $G(n,p)$ with $n=5623$. Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB. “Speedup ratio” compares igraph to our parallel method; “single core speedup” assumes single-core execution.
Edge
Prob. $p$
u-stats
Time (s)
igraph
Time (s)
Speedup
Ratio
Single-Core
Speedup
0.001 0.617 0.025 0.040 0.001
0.005 0.649 0.353 0.544 0.008
0.010 0.692 1.795 2.595 0.041
0.020 0.783 10.594 13.529 0.211
0.050 1.038 136.272 131.288 2.051
0.080 1.298 514.702 396.415 6.194
0.100 1.453 960.299 660.690 10.323
0.150 1.864 3003.490 1611.389 25.178
0.200 2.275 6725.515 2955.656 46.182
Table 5: Runtime comparison of exact all 44-vertex motif counts using u-stats and igraph on Erdös-Rényi graphs $G(n,p)$ with $n=2000$. Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB. “Speedup ratio” compares igraph to our parallel method; “single core speedup” assumes single-core execution.
Edge
Prob. $p$
u-stats
Time (s)
igraph
Time (s)
Speedup
Ratio
Single-Core
Speedup
0.001 70.959 0.004 0. 0.
0.005 71.947 0.168 0.002 0.
0.010 68.463 1.453 0.021 0.
0.020 65.634 15.573 0.237 0.004
0.050 65.786 394.532 5.997 0.094
0.080 66.866 2371.373 35.465 0.554
0.100 68.219 5396.057 79.099 1.236
0.150 65.969 23003.041 348.692 5.448
0.200 65.259 61796.760 946.940 14.796
4.3 Application 3: Distance Covariance Measures for Independence Testing
In the last example, we revisit the problem considered in Section 5.1 of shao2025u, which concerns the problem of computing the squared distance covariance $\mathsf{dCov}^{2}$, a recently proposed dependence measure between two random vectors (szekely2007measuring; yao2018testing). In particular, let $O_{i}\coloneqq(X_{i},Y_{i}),i\in[n]$ be the observed sample. $\mathsf{dCov}^{2}(X,Y)$ can then be written as a $4$-th order $U$-statistic
$\mathsf{dCov}^{2}(X,Y)=\frac{(n-4)!}{n!}\sum_{\bar{i}_{4}\in\mathsf{Perm}([n],4)}$ $\Bigg{(}h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{3}},O_{i_{4}})+h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{1}},O_{i_{2}})$
$-h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{1}},O_{i_{3}})-h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{2}},O_{i_{4}})\Bigg{)},$
where $h_{1}(O_{i},O_{j})\coloneqq\|X_{i}-X_{j}\|_{2}$ and $h_{2}(O_{i},O_{j})\coloneqq\|Y_{i}-Y_{j}\|_{2}$ for $i,j\in[n]$. Here $\|\cdot\|_{2}$ denotes the vector $\ell_{2}$-norm.
Quoting from Table 5 of shao2025u, it takes “8099.73 seconds” to compute $\mathsf{dCov}^{2}$ with $n=138$ among 11 random vectors. Each vector corresponds to a different stock sector—such as Health Care or Industrials—and its components represent the monthly log-returns of individual stocks within that sector, using historical S&P 500 data. To bypass such a high computational burden, shao2025u resort to computing a randomized incomplete $U$-statistic instead. In the same Table 5, shao2025u demonstrated that there is indeed a significant speedup when using randomized incomplete $U$-statistics. It is noteworthy that another motivation of shao2025u to use randomized incomplete $U$-statistic is due to its asymptotic normality (chen2019randomized) even when the original complete $U$-statistic has a Gaussian chaos limiting distribution. But we will not further discuss this issue as it is not directly relevant to the main theme of our paper.
In Table 6, we display the average runtime over 10 runs of exactly computing $\mathsf{dCov}^{2}$ using our package u-stats, or computing $\mathsf{dCov}^{2}$ exactly or approximately with the randomized incomplete $U$-statistic using the MATLAB code attached in the supplementary materials of shao2025u. We rerun the code of shao2025u together with u-stats on the same computing platform to ensure a fair comparison. In particular, the algorithm proposed in shao2025u will randomly choose $O(n^{\alpha})$ permutations, with $\alpha>0$ a tuning parameter determined by the available computation budget. By choosing $\alpha<4$, the running time will be reduced compared to $O(n^{4})$. We vary the tuning parameter $\alpha$ within the set $\{1.5,2,2.5\}$. Our package u-stats again exhibits quite impressive runtime performance: u-stats without parallel computation takes slightly shorter time than the randomized incomplete $U$-statistic when $\alpha=2$, whereas u-stats with parallel computation takes less than half of the time as the case when $\alpha=1.5$!
Table 6: Runtime (in seconds) comparison of various methods for computing $\mathsf{dCov}^{2}$. Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB.
u-stats shao2025u’s MATLAB code
No Parallel
Parallel
Randomized
$\alpha=1.5$
Randomized
$\alpha=2.0$
Randomized
$\alpha=2.5$
Complete
4.0928 0.1847 0.4211 4.5744 53.0265 3395.4001
5 Concluding Remarks
In this paper, we have conducted a delicate and systematic study on computing and the complexity of computing higher-order $U$-statistics (and also $V$-statistics), exactly. Explicit connections between the complexity of computing higher-order UU/$V$-statistics and concepts from tensor computation and graph theory are established. An accompanying Python package u-stats is provided to help practitioners compute UU/$V$-statistics more efficiently in applications. Although Python becomes more and more popular among statisticians (dai2023rehline), an R interface of u-stats will also be available in due course. It will also be interesting to test u-stats on other problems involving $U$-statistics (qiu2024neural; shen2025engression).
One limitation of our work is that we prioritize time complexity over space complexity. As a result, when we need to store a very high-order tensor, the sample size $n$ cannot be too large – this is why we have only evaluated the performance of u-stats on computing HOIF estimators with order $m\leq 7$ when $n=10000$. An important future direction is thus to explore how to balance the time-complexity and space-complexity when the sample size $n$ is large.
Finally, we conclude our paper with an open problem that we ourselves find intriguing. As we indicated in Remark 7 at the end of Section 3, the treewidth of the decomposition graph associated with the Einsum notation of the underlying $V$-statistic relates to its time complexity. But the treewidth of a graph is not easy to compute either analytically (harvey2018treewidth) or numerically (arnborg1987complexity). It will be interesting to get a sharp bound on treewidth using the number of vertices or edges of the graph. As mentioned in Remark 7, kneis2009bound has proved that the treewidth is bounded at most linearly by the edge size, or more precisely bounded by $e/5.769+O(\log n)$ where $n\coloneqq|V|$ denotes the number of vertices in $G$. Yet we do not know if this bound can be further improved. Below we provide the maximum tree-width possible for edge size no greater than $15$, to hopefully serve as a starting point of fully addressing this problem.
Observation 1 (Treewidth of Graphs with a Given Number of Edges).
Let
$\mathrm{t}(e)\coloneqq\max\left\{\mathsf{tw}(G)\;\middle|\;G\text{ is a graph where }|E(G)|=e\right\}.$
We have computed the exact values of $\mathrm{t}(e)$ for all $1\leq e\leq 15$ as follows:
[…]
The proof can be found in Appendix C.1. However, extending the above result to $e>15$ is currently beyond our reach.
Appendix A Proof of Proposition 1
Our proposed algorithm for computing a tensor $\mathcal{T}$} is based on iterating the Einsum operation w.r.t. the corresponding Einsum notation $\mathcal{A}$}, summing over one index at a time. We refer readers to Definition 4 in the main text for the meaning of the index set of an Einsum notation.
On a high level, the proof proceeds by establishing a correspondence between the Einsum notation and the associated decomposition graph (Definition 10). With the latter, we can leverage concepts from graph theory and combinatorics to obtain a time complexity upper bound of computing $\mathcal{T}$}. In particular, we will demonstrate that the summation over each index in the Einsum notation corresponds to eliminating one vertex of the decomposition graph following Definition 6. This characterization can help demonstrate that the treewidth of (Definition 7) the decomposition graph gives an optimistic estimate of the time complexity of the Einsum operation.
To ease exposition, we decide to present the proof in a slow pace. The roadmap of the proof is as follows. We first construct a sequence of Einsum notations, described in Definition 13 and Lemma 3, to mimic the computation procedure presented in Section 3.2. Next, in Lemma 4, we obtain the computational complexity of summing over an index, which corresponds to the degree of the associated vertex in the decomposition graph. Finally, we complete the proof by showing that the optimal ordering of summation over indices is equivalent to identifying the best vertex elimination order of the decomposition graph. This is why the treewidth of the decomposition graph can be used to give an optimistic estimate of the time complexity.
Definition 13 (Einsum Notation Sequence under a Particular Summation Ordering).
Let ${\mathcal{A}}=(A_{1},\cdots,A_{K})$ be an Einsum notation (recall from Definition 4) with no output of. Let $I=[m]$ be the index set of $\mathcal{A}$} and $\sigma$ be a permutation over $I$. The Einsum notation sequence of $\mathcal{N}$} under $\sigma$ is the sequence ${\mathcal{A}}^{\sigma}_{0},{\mathcal{A}}^{\sigma}_{1},\ldots,{\mathcal{A}}^{\sigma}_{m}$ defined recursively as follows:
  • ${\mathcal{A}}^{\sigma}_{0}={\mathcal{A}}$.
  • For each $i\in I$, let $s_{\sigma[i]}$ be a permutation of the following set
    $\bigcup_{\begin{subarray}{c}A\in{\mathcal{A}}^{\sigma}_{i-1}\\ \sigma[i]\in A\end{subarray}}(\mathsf{set}(A)\setminus\{\sigma[i]\}).$
    . Then ${\mathcal{A}}^{\sigma}_{i}$ is obtained from ${\mathcal{A}}^{\sigma}_{i-1}$ by:
    • Removing tuples in ${\mathcal{A}}_{i-1}^{\sigma}$ containing $\sigma[i]$.
    • Inserting $s_{\sigma[i]}$ as the last tuple into the tuple list if $s_{j}\neq\emptyset$.
    • Denoting the new tuple list as ${\mathcal{A}}^{\sigma}_{i}$.
We use the following concrete example to further illustrate the above definition.
Example 4.
Let ${\mathcal{A}}=((1,2),(1,3),(1,4),(2,3),(2,4),(3,4))$ and $\sigma=(1,2,3,4)$. Then the Einsum notation sequence of $({\mathcal{A}};\emptyset)$ under $\sigma$ is comprised of:
${\mathcal{A}}^{\sigma}_{0}$ $=((1,2),(1,3),(1,4),(2,3),(2,4),(3,4)),$
${\mathcal{A}}^{\sigma}_{1}$ $=((2,3),(2,4),(3,4),(2,3,4)),$
${\mathcal{A}}^{\sigma}_{2}$ $=((3,4),(3,4)),$
${\mathcal{A}}^{\sigma}_{3}$ $=((4)),$
${\mathcal{A}}^{\sigma}_{4}$ $=(()).$
The following lemma establishes a one-to-one correspondence between the Einsum notation sequence under a particular summation ordering $\sigma$ defined in Definition 13 and vertex elimination given in Definition 6.
Lemma 3 (The Einsum Notation Sequence, Vertex Elimination Correspondence).
Let $\mathcal{A}$} be an Einsum notation with index set $I=[m]$ and $\sigma$ be a permutation of $I$. Let the sequence ${\mathcal{A}}^{\sigma}_{0},{\mathcal{A}}^{\sigma}_{1},\ldots,{\mathcal{A}}^{\sigma}_{m}$ be the Einsum notation sequence w.r.t. $\mathcal{A}$} under permutation $\sigma$ (see Definition 13) and $G_{{\mathcal{A}}^{\sigma}_{0}},G_{{\mathcal{A}}^{\sigma}_{1}},\ldots,G_{{\mathcal{A}}^{\sigma}_{m}}$ be the corresponding decomposition graphs. Then for each $i\in[m]$, it holds that
$G_{{\mathcal{A}}^{\sigma}_{i}}=\mathsf{elim}_{\sigma[i]}(G_{{\mathcal{A}}^{\sigma}_{i-1}}).$
Proof.
Recalling from Definition 6 about the vertex elimination operation over a graph and Definition 10 about the decomposition graph associated with a tuple list $\mathcal{A}$} and tet $s_{\sigma[i]}$ be a permutation of the following set
$\bigcup_{\begin{subarray}{c}A\in{\mathcal{A}}^{\sigma}_{i-1}\\ \sigma[i]\in A\end{subarray}}(\mathsf{set}(A)\setminus\{\sigma[i]\}),$
we then recognize that removing all tuples containing $\sigma[i]$ from the tuple list ${\mathcal{A}}^{\sigma}_{i-1}$ corresponds to removing the vertex $\sigma[i]$ and all edges incident to $\sigma[i]$ in the decomposition graph $G_{{\mathcal{A}}^{\sigma}_{i-1}}$ associated with ${\mathcal{A}}^{\sigma}_{i-1}$ (some edges connected neighbors of $\sigma[i]$ may also be removed but will be re-added latter), while inserting the tuple $s_{\sigma[i]}$ into ${\mathcal{A}}^{\sigma}_{i-1}$ corresponds connecting the neighbors of $\sigma[i]$ into a clique, which completes the proof. ∎
Lemma 4.
[…]
Proof.
We immediately have |B|=d1|B|=d-1, which implies that the result of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ has $n^{d-1}$ entries. For each entry, as one index is summed over, it takes n1n-1 summation operations and $(|{\mathcal{A}}|-1)n$ multiplication operations. Therefore, the total time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ is $\big{(}(n-1)n^{d-1}+(|{\mathcal{A}}|-1)n^{d}\big{)}$, which is $\Theta(|{\mathcal{A}}|n^{d})$. ∎
[…]
Proof of Proposition 1.
Let $m\coloneqq|V(G_{\mathcal{A}})|$. Let $\sigma$ be a permutation of $[m]$, and $G_{{\mathcal{A}}^{\sigma}_{0}},G_{{\mathcal{A}}^{\sigma}_{1}},\ldots,G_{{\mathcal{A}}^{\sigma}_{m}}$ be the corresponding decomposition graphs. We define the ${\mathcal{A}}^{\sigma}_{0},{\mathcal{A}}^{\sigma}_{1},\ldots,{\mathcal{A}}^{\sigma}_{m}$ be the Einsum notation sequence under $\sigma$ as in Definition 13.
We first show that there is an algorithm such that the time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(n^{C_{\sigma}+1})$ where
$C_{\sigma}\coloneqq\max_{i\in[m]}\deg(\sigma[i],G_{{\mathcal{A}}^{\sigma}_{i-1}}).$
To see this, we use Einsum notation to characterize this procedure. First, for $i\in[m]\cup{0}$, we define the tensor tuple ${\mathcal{T}}^{\sigma}_{i}$ corresponding to the Einsum notation sequence ${\mathcal{A}}^{\sigma}_{i}$ as follows: ${\mathcal{T}}^{\sigma}_{0}\coloneqq{\mathcal{T}}$ and for each $i\in[m]$:
  • $\sigma[i]$ is the index to be eliminated at the $i$-th step;
  • Let ${\mathcal{A}}^{\sigma,\mathrm{in}}_{i}\sqsubseteq{\mathcal{A}}^{\sigma}_{i-1}$ be the sublist of ${\mathcal{A}}^{\sigma}_{i-1}$ that contains $\sigma[i]$, i.e.
    $A\in{\mathcal{A}}^{\sigma,\mathrm{in}}_{i}\iff\sigma[i]\in A,$
    and similarly let ${\mathcal{T}}^{\sigma,\mathrm{in}}_{i}\sqsubseteq{\mathcal{T}}^{\sigma}_{i-1}$ be the tensor tuple corresponding to ${\mathcal{A}}^{\sigma,\mathrm{in}}_{i}$;
  • Let $B^{\sigma,\mathrm{out}}_{i}$ be a permutation of the set $\bigcup_{\begin{subarray}{c}A\in{\mathcal{A}}^{\sigma}_{i-1}\\ \sigma[i]\in A\end{subarray}}(\mathsf{set}(A)\setminus\{\sigma[i]\})$ as the output tuple and $T^{\sigma,\mathrm{out}}_{i}\coloneqq\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$;
  • Update the tensor tuple ${\mathcal{T}}^{\sigma}_{i}$ from ${\mathcal{T}}^{\sigma}_{i-1}$ by
    • removing all tensors in ${\mathcal{T}}^{\sigma,\mathrm{in}}_{i}$ from ${\mathcal{T}}^{\sigma}_{i-1}$, and then
    • inserting $T^{\sigma,\mathrm{out}}_{i}$ to the end of ${\mathcal{T}}^{\sigma}_{i-1}$ to form ${\mathcal{T}}^{\sigma}_{i}$.
According to the distributive property of multiplication over summation, it holds that for any $i,j\in[m]$,
$\mathsf{Einsum}({\mathcal{T}}^{\sigma}_{i},{\mathcal{A}}^{\sigma}_{i})\equiv\mathsf{Einsum}({\mathcal{T}}^{\sigma}_{j},{\mathcal{A}}^{\sigma}_{j})\equiv\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}).$
Thus the time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ just comes from transferring from one step to the next step, which is $\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$. Following Lemma 4, we obtain that the time complexity of computing $\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$ is $\Theta(|{\mathcal{A}}^{\sigma,\mathrm{in}}_{i}|n^{\deg(x,G_{{\mathcal{A}}^{\sigma}_{i-1}})})$, which is also $\Theta(|{\mathcal{A}}|n^{\deg(\sigma[i],G_{{\mathcal{A}}^{\sigma}_{i-1}})})$ as $|{\mathcal{A}}^{\sigma,\mathrm{in}}_{i}|\leq|{\mathcal{A}}|$. By taking maximum over $i$, the overall time complexity is thus $O(|{\mathcal{A}}|n^{C_{\sigma}+1})$.
Finally, following Lemma 3, we have
$G_{{\mathcal{A}}^{\sigma}_{i}}=\mathsf{elim}_{\sigma[i]}(G_{{\mathcal{A}}^{\sigma}_{i-1}}),$
and by finding the permutation $\sigma^{\ast}$ minimizing $C_{\sigma}$ and recalling Definition 7, we can conclude that there exists an algorithm such that the time complexity of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(|{\mathcal{A}}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$, which completes the proof. ∎
Appendix B Proof of Lemma 1
Here we prove an extended version of Lemma 1, as shown in Lemma 5 below. Lemma 1 then becomes a simple corollary of Lemma 5. We begin by introducing the following definitions.
Definition 14 (Refinement of partitions).
Let $\pi$ and $\rho$ be two partitions of a non-empty set $S$. $\pi$ is said to be finer than $\rho$, if for any $A\in\pi$, there exists $B\in\rho$ such that $A\subseteq B$, denoted by $\pi\preceq\rho$. Additionally, $\pi$ is strictly finer than $\rho$ if $\pi\neq\pi$, denoted by $\rho\prec\pi$.
Remark 8.
Refinement defines a partial order on partitions of a set and in fact forms a partition lattice, a classical result in lattice theory (birkhoff1940lattice).
Similar to restricted $V$-statistics defined in Definition 12, we define U-set and restricted $U$-statistics as follows.
Definition 15 (U-set).
Let $m,n$ be positive integers satisfying $m\leq n$ and $\pi$ be a partition of the set $[m]$. A U-set of size $n$ w.r.t. the partition $\pi$ is defined as
${\mathcal{U}}\left(n,\pi\right)\coloneqq\left\{(s_{1},s_{2},\ldots,s_{m})\in[n]^{m}\mid\text{where for any $i,j\in[m]$},i,j\in Q\text{ for some $Q\in\pi$}\iff s_{i}=s_{j}\right\}.$
If $\pi$ happens to be the finest partition of $[m]$, i.e. $\pi=\{\{1\},\{2\},\cdots,\{m\}\}$, the corresponding U-set is denoted by ${\mathcal{U}}(n,m)$ and in fact, ${\mathcal{U}}(n,m)\equiv\mathsf{Perm}(n,m)$. The cardinality of the given partition $\pi$ is referred to as the order of ${\mathcal{U}}\left(n,\pi\right)$.
Definition 16 (Restricted $U$-statistics).
Let $\mathbb{X}$} be a non-empty set, $m,n$ be positive integers satisfying $m\leq n$, $\pi$ be a partition of set $[m]$ and $h:{\mathbb{X}}^{m}\rightarrow{\mathbb{R}}$ be a function defined on ${\mathbb{X}}^{m}$. For any tuple $\bm{X}\in{\mathbb{X}}^{n}:n\geq m$, the $U$-statistic with kernel $h$ restricted by partition $\pi$ takes the following form
$\mathbb{U}[\pi](h,\bm{X})\coloneqq\sum_{\alpha\in{\mathcal{U}}(n,\pi)}h(\bm{X}[\alpha]).$
Armed with Definitions 1416, we are ready to present the following extended form of Lemma 1.
Lemma 5 (Extended Form of Lemma 1).
Let
  • $\mathbb{X}$} be a non-empty set,
  • $m,n$ be integers satisfying $m\leq n$,
  • $h$ be a function defined on ${\mathbb{X}}^{m}$,
  • $\bm{X}\in{\mathbb{X}}^{n}$.
Then for any $\pi\in\Pi_{m}$, it holds that
$\mathbb{U}[\pi](h,\bm{X})=\sum_{\begin{subarray}{c}\rho\in\Pi_{m}\\ \pi\preceq\rho\end{subarray}}\mu(\pi,\rho)\mathbb{V}[\rho](h,\bm{X}),$
where the coefficients $\mu(\pi,\rho)$ are given by
$\mu(\pi,\rho)=(-1)^{|\pi|-|\rho|}\prod_{C\in\rho}(\kappa_{\pi,C}-1)!,\text{ and }\kappa_{\pi,C}\coloneqq|\{Q\in\pi\mid Q\subseteq C\}|.$
In particular, if $\pi$ happens to be $\{\{k\}\mid k\in[m]\}$, the above result reduces to Lemma 1.
The proof of Lemma 5 can be found in Appendix B.2.
B.1 Preparatory Materials
Before presenting the proof in the next section, we introduce some useful preparatory materials, including: (1) the decomposition of a $V$-statistic into $U$-statistics and (2) a dual form of Möbius inversion formula (stanley2011enumerative). We first state the following technical result of decomposing a $V$-statistic into a linear combination of $U$-statistics.
Lemma 6 (Decomposition of a $V$-statistic into $U$-statistics).
Let
  • $\mathbb{X}$} be a non-empty set,
  • $m,n$ be positive integers satisfying $m\leq n$,
  • $h$ be a function defined on ${\mathbb{X}}^{m}$,
  • $\bm{X}\in{\mathbb{X}}^{n}$.
Then for any $\pi\in\Pi_{m}$, it holds that
$\mathbb{V}[\pi](h,\bm{X})=\sum_{\begin{subarray}{c}\rho\in\Pi_{m}\\ \pi\preceq\rho\end{subarray}}\mathbb{U}[\rho](h,\bm{X}),$
where
$\mathbb{U}[\rho](h,\bm{X})=\sum_{\alpha\in{\mathcal{U}}\left(n,\rho\right)}h(\bm{X}[\alpha]).$
Proof.
It is equivalent to prove that $\{{\mathcal{U}}\left(n,\rho\right)\mid\rho\in\Pi_{m}:\pi\preceq\rho\}$ forms a partition (i.e. disjoint union) of ${\mathcal{V}}(n,\pi)$, i.e.
${\mathcal{V}}(n,\pi)=\bigsqcup_{\begin{subarray}{c}\rho\in\Pi_{m}\\ \pi\preceq\rho\end{subarray}}{\mathcal{U}}\left(n,\rho\right).$
Part 1: Disjointness
We first prove that for any $\pi_{1},\pi_{2}\in\{\rho\in\Pi_{m}\mid\pi\preceq\rho\}:\pi_{1}\neq\pi_{2}$, it holds that
${\mathcal{U}}\left(n,\pi_{1}\right)\cap{\mathcal{U}}\left(n,\pi_{2}\right)=\emptyset.$
Since $\pi_{1}\neq\pi_{2}$, there must exist $i_{1},i_{2}\in[m]$ such that either
  • $\exists\,Q_{1}\in\pi_{1}$, such that $i_{1},i_{2}\in Q_{1}$, and
  • $\forall\,Q_{2}\in\pi_{2}$, $i_{1},i_{2}\notin Q_{2}$,
or
  • $\forall\,Q_{1}\in\pi_{1}$, $i_{1},i_{2}\notin Q_{1}$,
  • $\exists\,Q_{2}\in\pi_{2}$, such that $i_{1},i_{2}\in Q_{2}$.
Without loss of generality, we assume that the former holds. Following Definition 15, we have for any $\alpha_{1}\in{\mathcal{U}}\left(n,\pi_{1}\right)$, $\alpha_{1}[i_{1}]=\alpha_{1}[i_{2}]$, but for any $\alpha_{2}\in{\mathcal{U}}\left(n,\pi_{2}\right)$, $\alpha_{1}[i_{1}]\neq\alpha_{1}[i_{2}]$, which implies that ${\mathcal{U}}\left(n,\pi_{1}\right)\cap{\mathcal{U}}\left(n,\pi_{2}\right)=\emptyset$.
Part 2: Coverage
We are left to prove that for any $\alpha\in{\mathcal{V}}(n,\pi)$, there exists $\rho\in\Pi_{m}:\pi\preceq\rho$, such that $\alpha\in{\mathcal{U}}\left(n,\rho\right)$. We first construct a candidate $\rho$ as follows. For each $j\in\mathsf{set}(\alpha)$, let
$Q_{j}\coloneqq\{i\in[m]\mid\alpha[i]=j\},$
and then let $\rho\coloneqq\{Q_{j}\mid j\in\mathsf{set}(\alpha)\}$. It is straightforward to verify that $\rho$ is a partition of $[m]$.
Then we prove that $\rho$ is the partition satisfying the desiderata. Following Definition 15, we have $\alpha\in{\mathcal{U}}\left(n,\rho\right)$. Following Definition 11, we have for any $Q\in\pi$ and any $i_{1},i_{2}\in Q$, $\alpha[i_{1}]=\alpha[i_{2}]$, which implies that for any $Q\in\pi$, $Q\subseteq B_{j_{Q}}\in\rho$, for some $j_{Q}$ such that
$\alpha[j_{Q}]=\alpha[i],\quad\forall\,i\in Q,$
We then obtain that $\pi\preceq\rho$ following Definition 14, which complete the proof of coverage.
Combining Part 1 and Part 2, we complete the proof of this lemma. ∎
Next, we introduce the notion of Möbius inversion formula, which is used to reverse the decomposition from a $V$-statistic into $U$-statistics in Lemma 6 to obtain the decomposition from a $U$-statistic into $V$-statistics.
Definition 17 (Möbius Function, (stanley2011enumerative)).
Let $(P,\preceq)$ be a finite poset and, a function $\mu:P^{2}\to{\mathbb{R}}$ is said to be a Möbius function on $(P,\preceq)$, if
$\mu(s,s)=$ $\ 1,\quad\forall\ s\in P,\text{ and}$
$\mu(s,u)=$ $-\sum_{\begin{subarray}{c}t\in P\\ s\preceq t\prec u\end{subarray}}\mu(s,t),\quad\forall\ s,u\in P:s\prec u.$
Lemma 7 (Möbius Function on Partition Lattices, (stanley2011enumerative)).
Let $S$ be a non-empty set. Then the Möbius function on $(\Pi_{S},\preceq)$ takes the following form
$\mu(\pi,\rho)=(-1)^{|\pi|-|\rho|}\prod_{C\in\rho}(\kappa_{\pi,C}-1)!,\quad\forall\,\pi,\rho\in\Pi_{S}:\pi\preceq\rho\text{ and }\kappa_{\pi,C}=|\{A\in\pi\mid A\subseteq C\}|.$
Lemma 8 (Möbius Inversion Formula, Dual Form, (stanley2011enumerative)).
Let $(P,\preceq)$ be a finite poset and $f,g$ be functions defined on $P$. Then
$g(s)=\sum_{\begin{subarray}{c}t\in P\\ s\preceq t\end{subarray}}f(t),\quad\forall\,s\in P,$
if and only if
$f(s)=\sum_{\begin{subarray}{c}t\in P\\ s\preceq t\end{subarray}}\mu(s,t)g(t),\quad\forall\,s\in P,$
where $\mu$ is the Möbius function on $(P,\preceq)$.
B.2 Proof of Lemma 5
In this section, we prove Lemma 5. The argument is straightforward given the preparations in the previous section.
Proof of Lemma 5.
If we fix $h$ and $\bm{X}$, $f(\pi)\coloneqq\mathbb{U}[\pi](h,\bm{X})$ and $g(\rho)\coloneqq\mathbb{V}[\rho](h,\bm{X})$ are functions defined on $\Pi_{m}$.
Since $\Pi_{m}$ is finite and the refinement defined in Definition 14 defines a partial order on $\Pi_{m}$, we can apply Lemma 8 to functions $f$ and $g$.
Combining the Möbius function on $\Pi_{m}$ as stated in Lemma 7, the dual form of Möbius inversion formula given in Lemma 8 and the relationship between $f$ and $g$ as stated in Lemma 6, we immediately complete the proof. ∎
Appendix C Proof of Observation 1 and Results in Section 4.1
C.1 Proof of Observation 1
We begin by recalling some well-known bounds on treewidth. We then restrict our attention to simple graphs without isolated vertices and derive recurrence relations in several basic cases. The final proof proceeds by carefully verifying the conditions of Lemma 10 to be stated below.
Definition 18 (Degeneracy).
The degeneracy of a graph $G$, denoted by $D(G)$, is defined as the maximum value of the minimum degree $\delta(H)$ over all subgraphs $H$ of $G$.
Lemma 9 (Bounds on Treewidth).
Let $G=(V,E)$ be a graph. Then
  1. 1.
    $D(G)\leq\mathsf{tw}(G),$
  2. 2.
    if $H$ is a subgraph of $G$, then $\mathsf{tw}(H)\leq\mathsf{tw}(G)$,
  3. 3.
    for any vertex $v\in V$, it holds that $\mathsf{tw}(G)\leq\max\{\deg(v,G),\mathsf{tw}(\mathsf{elim}_{v}(G))\}$.
Proof.
The first inequality $D(G)\leq\mathsf{tw}(G)$ is established in Lemma 11 of bodlaender2011treewidth.
The second inequality is stated as Lemma 8 in the same reference.
The final claim follows directly from Definition 7: treewidth is defined as the minimum over all possible vertex elimination orderings, whereas the right-hand side corresponds to the class of orderings in which the vertex $v$ is eliminated first. ∎
Lemma 10 (Maximum Treewidth under Vertex and Edge Constraints).
Let $\mathcal{G}_{n,e}$ be the set of all simple graphs with $n$ vertices, $e$ edges, and no isolated vertices, and let $\mathcal{G}_{n,e}=\emptyset$ if no such graphs exist. Let $\mathrm{tt}(n,e)$ denote the maximum treewidth of the graphs in $\mathcal{G}_{n,e}$:
$\mathrm{tt}(n,e)\coloneqq\max\left\{\mathsf{tw}(G)\;\middle|\;G\in\mathcal{G}_{n,e}\right\}.$
The following properties hold:
  1. 1.
    Support set: For any fixed $e\in\mathbb{N}^{+}$, $\mathcal{G}_{n,e}$ is nonempty iff $n\in\mathrm{n}(e)\coloneqq\left\{n\in\mathbb{N}^{+}\,\middle|\,\binom{n}{2}\geq e,n\leq 2e\right\}$.
  2. 2.
    Representation of $\mathrm{t}(e):$ $\mathrm{t}(e)=\max_{n\in\mathrm{n}(e)}\mathrm{tt}(n,e).$
  3. 3.
    Monotonicity in edge counts: For any fixed $n:n\in\mathrm{n}(e+1)$, the function $\mathrm{tt}(n,e)$ is non-decreasing w.r.t. $e$.
  4. 4.
    Complete-graph treewidth: $\mathrm{tt}(n,\binom{n}{2})=n-1$, $\mathrm{tt}(n,\binom{n}{2}-1)=n-2$.
  5. 5.
    Recurrence relations in simple cases: recall that $\frac{2e}{n}$ is in fact the average degree of the graph and when $e>2$ and $n\in\mathrm{n}(e)$:
    • (1)
      If $1\leq\frac{2e}{n}<2$, then $\mathrm{tt}(n,e)\leq\max\{1,\mathrm{t}(e-1)\}$;
    • (2)
      If $1\leq\frac{2e}{n}<3$, then $\mathrm{tt}(n,e)\leq\max\{2,\mathrm{t}(e-1)\}$;
    • (3)
      If $1\leq\frac{2e}{n}<4$, then $\mathrm{tt}(n,e)\leq\max\{3,\mathrm{tt}(n-1,e),\mathrm{t}(e-1)\}$.
    • (4)
      $\mathrm{tt}(7,14)\leq\max\{4,\mathrm{t}(13)\}$.
Proof.
We prove the lemma in the following steps.
  1. 1.
    Since a simple graph with $n$ vertices and no isolated vertices can have at most $\binom{n}{2}$ edges, we must have $e\leq\binom{n}{2}$. Moreover, as isolated vertices are not allowed, every vertex must have degree at least one. In particular, the average degree satisfies $\frac{2e}{n}\geq 1$, which implies $n\leq 2e$.
  2. 2.
    The definition of $\mathrm{t}(e)$ does not exclude graphs with isolated vertices. However, for any graph $G$ with $e$ edges and isolated vertices, eliminating those vertices (with degree 0) does not affect the number of edges or the treewidth. Therefore, it suffices to take the maximum over all graphs in $\mathcal{G}_{n,e}$ for $n\in\mathrm{n}(e)$, i.e. taking the maximum over $\mathrm{tt}(n,e)$ for $n\in\mathrm{n}(e)$.
  3. 3.
    We observe that any graph $G\in\mathcal{G}_{n,e}$ can be extended to a graph $G^{\prime}\in\mathcal{G}_{n,e+1}$ by adding an edge, since $n\in\mathrm{n}(e+1)$. By the subgraph monotonicity of treewidth stated in Lemma 9, it follows that $\mathsf{tw}(G)\leq\mathsf{tw}(G^{\prime})$. Therefore, we have $\mathrm{t}(e)\leq\mathrm{t}(e+1)$.
  4. 4.
    First, consider the case where $e=\binom{n}{2}$. The only graph (up to isomorphism) with $n$ vertices and $\binom{n}{2}$ edges is the complete graph $K_{n}$. Since the degeneracy of $K_{n}$ is n1n-1, we have $\mathsf{tw}(K_{n})\geq n-1$. To show the upper bound, note that eliminating any vertex in $K_{n}$ removes a vertex of degree n1n-1 and yields $K_{n-1}$. By the third bound in Lemma 9, we get $\mathsf{tw}(K_{n})\leq\max\{n-1,\mathsf{tw}(K_{n-1})\}$. Repeating this process down to $K_{2}$, we have $\mathsf{tw}(K_{n})\leq\max\{n-1,\dots,\mathsf{tw}(K_{2})\}=n-1$, since $\mathsf{tw}(K_{2})=1$ is easy to verify. Therefore, $\mathrm{tt}(n,\binom{n}{2})=n-1$.
    Next, consider $e=\binom{n}{2}-1$. There is still only one graph in this case (up to isomorphism), obtained by deleting one edge from $K_{n}$. In this graph, two vertices have degree n2n-2 and the rest have degree n1n-1. Since it contains $K_{n-1}$ as a subgraph, we have a lower bound $\mathsf{tw}\geq n-2$ by the degeneracy. For the upper bound, eliminating one of the vertices with degree n2n-2 yields $K_{n-2}$. Thus, $\mathsf{tw}\leq\max\{n-2,\mathsf{tw}(K_{n-2})\}=n-2$. Therefore, $\mathrm{tt}(n,\binom{n}{2}-1)=n-2$.
  5. 5.
    Let us introduce another notation: let $\mathcal{G}_{e}\coloneqq\bigcup_{n\in\mathrm{n}(e)}\mathcal{G}_{n,e}$.
    Case (1): $1\leq\frac{2e}{n}<2$, the minimum degree in any graph in $\mathcal{G}_{n,e}$ must be exactly $1$. Eliminating such a vertex removes exactly one edge and yields a graph in $\mathcal{G}_{e-1}$. Therefore, by the third bound in Lemma 9, we have $\mathrm{tt}(n,e)\leq\max\{1,\mathrm{t}(e-1)\}$.
    Case (2): $1\leq\frac{2e}{n}<3$, the minimum degree in any graph from $\mathcal{G}_{n,e}$ must be either $1$ or $2$. If the minimum degree is $2$, eliminating such a vertex removes two edges (if its neighbors are adjacent) or one edge (if they are not, in which case we need to add one edge to connect them), and in both cases, one vertex is removed. The resulting graph lies in either $\mathcal{G}_{n-1,e-1}$ or $\mathcal{G}_{n-1,e-2}$. On the other hand, if the minimum degree is $1$, then as in the previous case, eliminating that vertex removes one edge and results in a graph in $\mathcal{G}_{e-1}$. Therefore, by the vertex elimination bound (Lemma 9), we have $\mathrm{tt}(n,e)\leq\max\{2,\mathrm{t}(e-1),\mathrm{tt}(n-1,e-1),\mathrm{tt}(n-1,e-2)\}$. Since $\mathrm{tt}(n-1,e-2)\leq\mathrm{tt}(n-1,e-1)\leq\mathrm{t}(e-1)$, we conclude $\mathrm{tt}(n,e)\leq\max\{2,\mathrm{t}(e-1)\}$.
    Case (3): $1\leq\frac{2e}{n}<4$. The minimum degree in any graph from $\mathcal{G}_{n,e}$ must be either $1$, 22, or $3$. If the minimum degree is $3$, then eliminating such a vertex removes between 0 and $3$ edges, depending on the connectivity of its neighbors, but in all cases it only removes exactly one vertex. Therefore, the resulting graph lies in $\mathcal{G}_{n-1,e}$, $\mathcal{G}_{n-1,e-1}$, $\mathcal{G}_{n-1,e-2}$, or $\mathcal{G}_{n-1,e-3}$. Combining this with the previous cases for minimum degree 11 and $2$, we obtain the recurrence relation: $\mathrm{tt}(n,e)\leq\max\{3,\mathrm{tt}(n-1,e),\mathrm{tt}(n-1,e-1),\mathrm{tt}(n-1,e-2),\mathrm{tt}(n-1,e-3),\mathrm{t}(e-1)\}\leq\max\{3,\mathrm{tt}(n-1,e),\mathrm{t}(e-1)\}$.
    Case (4): $n=7$, $e=14$. First we observe that $\frac{2e}{n}=4$, so the minimum degree in any graph $G=(V,E)$ from $\mathcal{G}_{n,e}$ must be $1$, 22, 33, or $4$.
    We focus on the case where the minimum degree is $4$, and argue that in this case, there must exist a vertex $v_{1}$ of degree exactly $4$ such that eliminating $v_{1}$ does not increase the number of edges.
    In general, eliminating a vertex of degree $4$ may remove 4 edges and add up to $\binom{4}{2}=6$ fill-in edges among its neighbors, so the net change in edge counts, denoted by $\Delta_{e}$, ranges from 4-4 to +2+2. We now show that in our setting, increasing the number of edges (i.e., $\Delta_{e}>0$) is impossible.
    Let $v_{1}$ be a vertex of degree $4$, and denote its neighborhood as $\{v_{2},v_{3},v_{4},v_{5}\}=N_{G}(v_{1})$. Partition the vertex set as $V=\{v_{1}\}\cup\{v_{2},v_{3},v_{4},v_{5}\}\cup\{v_{6},v_{7}\}$.
    Suppose that eliminating $v_{1}$ leads to $\Delta_{e}=2$. This implies that $N_{G}(v_{1})$ contains no edges, so 6 fill-in edges are added, making the neighborhood a clique. Now we count the maximum possible number of edges under this configuration: the 4 edges incident to $v_{1}$, no edges inside $\{v_{2},v_{3},v_{4},v_{5}\}$, up to 1 edge inside $\{v_{6},v_{7}\}$, and at most $2\cdot 4=8$ edges between $\{v_{6},v_{7}\}$ and $\{v_{2},v_{3},v_{4},v_{5}\}$. Hence, the total number of edges is at most $4+0+8+1=13<14=e$, contradicting the assumption. Therefore, $\Delta_{e}=2$ is impossible.
    Now consider the case where eliminating $v_{1}$ leads to $\Delta_{e}=1$. Then $N_{G}(v_{1})$ must contain exactly one edge, say connecting $v_{2}$ and $v_{3}$, and the edges between $\{v_{6},v_{7}\}$ and $\{v_{2},v_{3},v_{4},v_{5}\}$ and inside $\{v_{6},v_{7}\}$ all appear. This allows a maximum of $4+1+8+1=14$ edges. However, in this setting, $v_{4}$ and $v_{5}$ are not adjacent to each other or to $v_{2}$ or $v_{3}$, and their neighbors are only in $\{v_{1},v_{6},v_{7}\}$. Thus, each of them has degree at most $3$, contradicting the assumption that the minimum degree is $4$.
    Combining the above arguments, we conclude that eliminating $v_{1}$ does not increase the number of edges. Hence, if the minimum degree is $4$, then there exists a vertex of degree $4$ such that eliminating this vertex yields a graph in $\mathcal{G}_{n-1,e^{\prime}}$ with $e^{\prime}\leq e$.
    Now combining this with previous arguments for when the minimum degree is $1$, 22, or $3$, we obtain the recurrence relation: $\mathrm{tt}(n,e)\leq\max\{4,\mathrm{tt}(n-1,e),\mathrm{tt}(n-1,e-1),\mathrm{tt}(n-1,e-2),\mathrm{tt}(n-1,e-3),\mathrm{tt}(n-1,e-4),\mathrm{t}(e-1)\}\leq\max\{4,\mathrm{tt}(n-1,e),\mathrm{t}(e-1)\}$.
    Since $\mathrm{tt}(n-1,e)=\mathrm{tt}(6,14)=4$ by Property 4 shown earlier, we conclude: $\mathrm{tt}(7,14)\leq\max\{4,\mathrm{t}(13)\}$.
Armed with the above results on treewidth, we prove Observation 1.
Proof of Observation 1.
We verify the values of $\mathrm{t}(e)$ for each $1\leq e\leq 15$ by analyzing the sets $\mathcal{G}_{n,e}$ and showing that the lower and upper bounds coincide. Let $\mathrm{u}(e)$ (resp. $\mathrm{l}(e)$) denote the target upper (resp. lower) bound of $\mathrm{t}(e)$ for each $e$. Our goal is to prove that $\mathrm{t}(e)=\mathrm{l}(e)=\mathrm{u}(e)$ for all $e\in\{1,2,\dots,15\}$.
Lower bound. For each $e$, a specific graph $G_{e}$ is provided in Figure 3. Each $G_{e}$ is a simple graph with $e$ edges and contains a clique $K_{\mathrm{l}(e)+1}$ as a subgraph. By the definition of degeneracy and Lemma 9, we have:
$\mathrm{l}(e)\leq D(G_{e})\leq\mathsf{tw}(G_{e})\leq\mathrm{t}(e).$
Upper bound. We now show that $\mathrm{t}(e)\leq\mathrm{u}(e)$ by bounding $\mathrm{tt}(n,e)$ for all $n\in\mathrm{n}(e)$, using the properties in Lemma 10.
From Property 4, we obtain:
$\mathrm{tt}(2,1)$ $=1,\mathrm{tt}(3,3)=2,\mathrm{tt}(4,6)=3,$
$\mathrm{tt}(5,10)$ $=4,\mathrm{tt}(6,15)=5,\mathrm{tt}(5,9)=3,\mathrm{tt}(6,14)=4.$
Case $e=1,2$:
We have $\mathrm{n}(1)=\{2\}$ and $\mathrm{n}(2)=\{3,4\}$, so:
$\mathrm{t}(1)=\mathrm{tt}(2,1)=1.$
For $e=2$, both $n=3$ and $n=4$ satisfy $1\leq\frac{2e}{n}<2$. By Property 5:
$\mathrm{tt}(3,2)$ $\leq\max\{1,\mathrm{t}(1)\}=1,$
$\mathrm{tt}(4,2)$ $\leq\max\{1,\mathrm{t}(1)\}=1,$
and hence,
$\mathrm{t}(2)=1.$
Case $e=3,4,5$:
The support sets are:
$\mathrm{n}(3)$ $=\{3,4,5,6\},$
$\mathrm{n}(4)$ $=\{4,5,6,7,8\},$
$\mathrm{n}(5)$ $=\{4,5,\dots,10\}.$
Observe that
$\mathrm{n}(3)\setminus\{3\}\subseteq\mathrm{n}(4)\subseteq\mathrm{n}(5).$
By monotonicity (Part 3 in Lemma 10), we have:
$\max_{n\in\mathrm{n}(3)\setminus\{3\}}\mathrm{tt}(n,3)\leq\mathrm{t}(4)\leq\mathrm{t}(5).$
Since $\mathrm{tt}(3,3)=2$ and the lower bound is $2$, it follows that:
$2\leq\mathrm{t}(3)\leq\mathrm{t}(4)\leq\mathrm{t}(5).$
Now, for each $e\in\{3,4,5\}$ and all $n\in\mathrm{n}(e)$, we have $1\leq\frac{2e}{n}<3$. By Property 5:
$\mathrm{t}(5)\leq\max\{2,\mathrm{t}(4)\}\leq\max\{2,\mathrm{t}(3)\}\leq\max\{2,\mathrm{t}(2)\}=2,$
and thus:
$\mathrm{t}(3)=\mathrm{t}(4)=\mathrm{t}(5)=2.$
Case $e=6,7,8,9$:
The support sets are:
$\mathrm{n}(6)$ $=\{4,5,\dots,12\},$
$\mathrm{n}(7)$ $=\{5,6,\dots,14\},$
$\mathrm{n}(8)$ $=\{5,6,\dots,16\},$
$\mathrm{n}(9)$ $=\{5,6,\dots,18\}.$
We see that $\mathrm{tt}(4,6)=3$, and that
$\mathrm{n}(6)\setminus\{4\}\subseteq\mathrm{n}(7)\subseteq\mathrm{n}(8)\subseteq\mathrm{n}(9),$
the lower bound is $3$. Therefore, by monotonicity:
$3\leq\mathrm{t}(6)\leq\mathrm{t}(7)\leq\mathrm{t}(8)\leq\mathrm{t}(9).$
More specifically, note that
$\mathrm{tt}(5,6)\leq\mathrm{tt}(5,7)\leq\mathrm{tt}(5,8)\leq\mathrm{tt}(5,9)=3.$
For all $e\in\{6,7,8,9\}$ and $n\in\mathrm{n}(e)\setminus\{4,5\}$, we have $1\leq\frac{2e}{n}<4$, so by Property 5:
$\mathrm{tt}(n,e)\leq\max\{3,\mathrm{t}(e-1),\mathrm{tt}(5,e)\}\leq\max\{3,\mathrm{t}(e-1)\}.$
Hence,
$\mathrm{t}(9)\leq\max\{3,\mathrm{t}(8)\}\leq\max\{3,\mathrm{t}(7)\}\leq\max\{3,\mathrm{t}(6)\}\leq\max\{3,\mathrm{t}(5)\}=3,$
and therefore,
$\mathrm{t}(6)=\mathrm{t}(7)=\mathrm{t}(8)=\mathrm{t}(9)=3.$
Case $e=10,11,12,13,14$:
The support sets are:
$\mathrm{n}(10)$ $=\{5,6,\dots,20\},$
$\mathrm{n}(11)$ $=\{6,7,\dots,22\},$
$\mathrm{n}(12)$ $=\{6,7,\dots,24\},$
$\mathrm{n}(13)$ $=\{6,7,\dots,26\},$
$\mathrm{n}(14)$ $=\{6,7,\dots,28\}.$
We know that $\mathrm{tt}(5,10)=4$, the lower bound is $4$, and
$\mathrm{n}(10)\setminus\{5\}\subseteq\mathrm{n}(11)\subseteq\mathrm{n}(12)\subseteq\mathrm{n}(13)\subseteq\mathrm{n}(14),$
Thus, by monotonicity:
$4\leq\mathrm{t}(10)\leq\mathrm{t}(11)\leq\mathrm{t}(12)\leq\mathrm{t}(13)\leq\mathrm{t}(14).$
From Property 5 of Lemma 10 (Case (4)), we obtain:
$\mathrm{tt}(7,14)\leq\max\{4,\mathrm{t}(13)\}.$
And by monotonicity:
$\mathrm{tt}(6,10)\leq\mathrm{tt}(6,11)\leq\mathrm{tt}(6,12)\leq\mathrm{tt}(6,13)\leq\mathrm{tt}(6,14)=4,$
For $e=14$ and $n\in\mathrm{n}(14)\setminus\{5,6,7\}$, we have $1\leq\frac{2e}{n}<4$, so:
$\mathrm{tt}(n,14)\leq\max\{3,\mathrm{tt}(7,14),\mathrm{t}(13)\}\leq\max\{4,\mathrm{t}(13)\},$
and hence $\mathrm{t}(14)\leq\mathrm{t}(13)$.
For $e\in\{10,11,12,13\}$ and $n\in\mathrm{n}(e)\setminus\{5,6\}$, Property 5 gives:
$\mathrm{tt}(n,e)\leq\max\{3,\mathrm{t}(e-1),\mathrm{tt}(6,e)\}\leq\max\{4,\mathrm{t}(e-1)\}.$
Thus, we obtain:
$\mathrm{t}(14)\leq\mathrm{t}(13)\leq\max\{4,\mathrm{t}(12)\}\leq\max\{4,\mathrm{t}(11)\}\leq\max\{4,\mathrm{t}(10)\}\leq\max\{4,\mathrm{t}(9)\}=4,$
and therefore,
$\mathrm{t}(10)=\mathrm{t}(11)=\mathrm{t}(12)=\mathrm{t}(13)=\mathrm{t}(14)=4.$
Case $e=15$:
The support set is $\mathrm{n}(15)=\{6,7,\dots,30\}$.
We have $\mathrm{tt}(6,15)=5$, and $\mathrm{tt}(7,15)\leq\mathrm{tt}(7,20)=5$. For all other $n\in\mathrm{n}(15)\setminus\{6,7\}$, since $1\leq\frac{2e}{n}<4$, they are bounded by $\max\{3,\mathrm{tt}(7,15),\mathrm{t}(14)\}=5$, Thus we conclude $\mathrm{t}(15)=5$. ∎
(a) $G_{1}(K_{2})$
$e=1$, $\mathrm{t}(e)=1$
(b) $G_{2}$
$e=2$, $\mathrm{t}(e)=1$
(c) $G_{3}(K_{3})$
$e=3$, $\mathrm{t}(e)=2$
(d) $G_{4}$
$e=4$, $\mathrm{t}(e)=2$
(e) $G_{5}$
$e=5$, $\mathrm{t}(e)=1$
(f) $G_{6}(K_{4})$
$e=6$, $\mathrm{t}(e)=3$
(g) $G_{7}$
$e=7$, $\mathrm{t}(e)=3$
(h) $G_{8}$
$e=8$, $\mathrm{t}(e)=4$
(i) $G_{9}$
$e=9$, $\mathrm{t}(e)=3$
(j) $G_{10}(K_{5})$
$e=10$, $\mathrm{t}(e)=4$
(k) $G_{11}$
$e=11$, $\mathrm{t}(e)=4$
(l) $G_{12}$
$e=12$, $\mathrm{t}(e)=4$
(m) $G_{13}$
$e=13$, $\mathrm{t}(e)=4$
(n) $G_{14}$
$e=14$, $\mathrm{t}(e)=4$
(o) $G_{15}$ ($K_{6}$)
$e=15$, $\mathrm{t}(e)=5$
Figure 3: Graphs $G_{1}$ to $G_{15}$ with increasing edges and maximum treewidth.
C.2 Proof of Results in the Rightmost Column of Table 1
We can derive the sharp treewidth bounds presented in Table 1 using similar techniques to the above. For each $m\in\{2,3,\cdots,12\}$, the corresponding decomposition signature takes the following form: ${\mathcal{A}}_{m}=((1,2),\cdots,(m-1,m))$. Hence by Definition 10, the corresponding decomposition graph $G_{{\mathcal{A}}_{m}}$ is the graph with $V(G_{{\mathcal{A}}_{m}})=[m]$, $E(G_{{\mathcal{A}}_{m}})=\{\{i,i+1\}\}_{i=1}^{m-1}$, whose structure is illustrated in Figure 4(a).
The result is formally stated in the following observation.
Observation 2.
For $m=2,3,\cdots,12$, let ${\mathcal{A}}_{m}=((1,2),\cdots,(m-1,m))$, and $G_{{\mathcal{A}}_{m}}$ be the corresponding decomposition graph. Similar to Corollary 2, $M(m)$ is defined as
$M(m)\coloneqq\max\{\mathsf{tw}(G_{{\mathcal{A}}_{m}}/\pi)\mid\pi\in\Pi_{m}^{{\mathcal{A}}_{m}}\}.$
Then the following hold:
[…]
Proof.
Let $\mathcal{G}^{\mathsf{HOIF}}_{m}\coloneqq\{G_{{\mathcal{A}}_{m}}/\pi\mid\pi\in\Pi_{m}^{{\mathcal{A}}_{m}}\}$ then $M(m)\equiv\max\{\mathsf{tw}(G)\mid G\in\mathcal{G}^{\mathsf{HOIF}}_{m}\}$ and it holds that
$|E(G)|\leq m-1,\;\forall\,G\in\mathcal{G}^{\mathsf{HOIF}}_{m}.$
A direct application of Observation 1 gives us the following upper bounds: $M(m)\leq\mathrm{t}(m-1)$, summarized in Table 7. The desired result then follows from a characterization of the lower bound of $M(m)$.
Table 7: Upper bounds of $M(m)$ at different $m$’s.
$m$ $e$ $\mathrm{t}(e)$
2, 3 1 1
4, 5, 6 3, 4, 5 2
7, 8, 9, 10 6, 7, 8, 9 3
11, 12 10, 11 4
We observe that except for $m=7$, the upper bounds match the claimed results in (6). The lower bounds of $M(m)$ can be obtained by enlisting some concrete examples, which are given in Table 8.
Table 8: Lowers bound of $M(m)$: the exemplified graph whose treewidth reaches the lower bound of $M(m)$ at different $m$’s.
$m$ $\pi$ Graph $G_{{\mathcal{A}}_{m}}/\pi$ $\mathsf{tw}(G_{{\mathcal{A}}_{m}}/\pi)$
2 {{1}, {2}} Figure 3(a): $G_{1}$ 1
3 {{1}, {2}, {3}} Figure 3(b): $G_{2}$ 1
4 {{1, 4}, {2}, {3}} Figure 3(c): $G_{3}$ 2
5 {{1, 4}, {2}, {3}, {5}} Figure 3(d): $G_{4}$ 2
6 {{1, 4}, {2}, {3}, {5}, {6}} Figure 3(e): $G_{5}$ 2
7 {{1, 4}, {2}, {3}, {5}, {6}, {7}} Figure 4(b): $G_{6}^{\prime}$ 2
8 {{1, 4}, {2, 6}, {3, 7}, {5, 8}} Figure 3(g): $G_{7}$ 3
9 {{1, 4}, {2, 6}, {3, 7}, {5, 8}, {9}} Figure 3(h): $G_{8}$ 3
10 {{1, 4}, {2, 6}, {3, 7}, {5, 8}, {9}, {10}} Figure 3(i): $G_{9}$ 3
11 {{1, 6, 11}, {2, 9}, {3, 7}, {4, 10}, {5, 8}} Figure 3(j): $G_{10}$ 4
12 {{1, 6, 11}, {2, 9}, {3, 7}, {4, 10}, {5, 8}, {12}} Figure 3(k): $G_{11}$ 4
(a) $G_{{\mathcal{A}}_{m}}$ with $m=4$.
(b) $G^{\prime}_{6}:$ the example in Table 8 when $m=7$.
Figure 4: Two decomposition graphs used in Appendix C.2.
Combining Table 7 and Table 8, we can conclude that, except for $m=7$, the upper bound (given in Table 7) and the lower bound (given in Table 8) of $M(m)$ agree. For $m=7$, we have $M(7)\in\{2,3\}$. In the sequel, we will show that $M(7)<3$ so the claim (6) in Observation 2 holds.
To this end, we will first establish a lemma (Lemma 11) stating that a simple graph with at most $6$ edges has treewidth $3$ if and only if it is $K_{4}$ (i.e. the complete graph with $4$ vertices). Then, another lemma (Lemma 12) will show that the set of all quotient graphs of $G_{{\mathcal{A}}_{7}}$ does not contain $K_{4}$. The proof is now complete. ∎
We first state and prove Lemma 11.
Lemma 11.
Recall the definitions of $\mathrm{tt}(n,e)$ and $\mathcal{G}_{n,e}$ in Lemma 10. When $1\leq e\leq 6$, $\mathrm{tt}(n,e)=3$ if and only if $n=4,e=6$, and otherwise $\mathrm{tt}(n,e)<3$. Moreover, $\mathcal{G}_{4,6}=\{K_{4}\}$ up to isomorphism.
Proof.
We adopt the same notation as in Lemma 10.
First, by Observation 1, we know $\mathrm{t}(e)<3$ when $1\leq e\leq 5$, and $\mathrm{t}(6)=3$. Thus we only need to consider the case $e=6$. Note that the support set of $\mathcal{G}_{n,6}$ is $\mathrm{n}(6)=\{4,5,\dots,12\}$.
For $n=4$, we have $\binom{4}{2}=6$, and by Property 4 in Lemma 10, $\mathrm{tt}(4,6)=3$ directly follows.
For any $n\in\mathrm{n}(6)\setminus\{4\}$, we have $1\leq\frac{2\cdot 6}{n}<3$. Therefore, $\mathrm{tt}(n,6)\leq\max\{2,\mathrm{t}(5)\}<3$. This proves the first claim.
For the second claim, note that any simple graph with $4$ vertices can have at most $\binom{4}{2}=6$ edges, and the only such graph (up to isomorphism) is $K_{4}$. ∎
Lastly, we state and prove Lemma 12.
Lemma 12.
Let $G_{{\mathcal{A}}_{7}}$ be the decomposition graph of decomposition signature ${\mathcal{A}}_{7}=((i,i+1))_{i=1}^{6}$. Then no quotient graph of $G_{{\mathcal{A}}_{7}}$ is $K_{4}$ or contains $K_{4}$ as a subgraph.
Proof.
Consider a partition $\pi=\{Q_{1},Q_{2},\dots,Q_{K}\}\in\Pi_{7}$ of the vertex set of $G_{{\mathcal{A}}_{7}}$, as defined in Definition 8. The quotient graph $G_{{\mathcal{A}}_{7}}/\pi$ has $|\pi|=K\leq 7$ vertices, with each vertex corresponding to a subset $Q_{k}$ for $k\in[K]$ of the partition $\pi$. An edge exists between $Q_{i}$ and $Q_{j}$ ($i\neq j$) in the quotient graph if and only if there is at least one edge in $G_{{\mathcal{A}}_{7}}$ connecting a vertex in $Q_{i}$ to a vertex in $Q_{j}$.
We observe that $|E(G_{{\mathcal{A}}_{7}}/\pi)|\leq|E(G_{{\mathcal{A}}_{7}})|=6$ for all $\pi\in\Pi_{7}$, and the inequality is strict whenever any subset $Q_{k}$ contains adjacent vertices in $G_{{\mathcal{A}}_{7}}$.
Since $|E(G_{{\mathcal{A}}_{7}})|=6$ and $|E(K_{4})|=\binom{4}{2}=6$, no quotient graph of $G_{{\mathcal{A}}_{7}}$ can contain $K_{4}$ as a proper subgraph (i.e. a subgraph that is not the original graph). Therefore, it suffices to consider the case where $K_{4}$ is isomorphic to a quotient graph of $G_{{\mathcal{A}}_{7}}$. For simplicity, we write “==” to denote graph isomorphism.
Suppose that there exists a partition $\pi^{*}=\{Q_{1},Q_{2},Q_{3},Q_{4}\}\in\Pi_{7}$ such that $G_{{\mathcal{A}}_{7}}/\pi^{*}=K_{4}$. Then, $\pi^{*}$ must satisfy the following conditions:
  1. 1.
    No subset $Q_{i}$ contains adjacent numbers, i.e., $(k,k+1)$ for $k\in\{1,2,\dots,6\}$ and $i\in[4]$;
  2. 2.
    For any $i\neq j$, there exist $k\in Q_{i}$ and $k^{\prime}\in Q_{j}$ such that $|k-k^{\prime}|=1$.
The first condition ensures that $|E(G_{{\mathcal{A}}_{7}}/\pi)|=6$. The second condition ensures that every pair of subsets is connected by at least one edge in the quotient graph. Since $G_{{\mathcal{A}}_{7}}$ has only $6$ edges, each edge must correspond uniquely to a distinct pair of subsets, covering all $6$ edges of $K_{4}$.
Define a map $c:[7]\to[4]$ by
$c(k)=i\quad\text{for all }k\in Q_{i}.$
This map can be interpreted as a coloring of the vertices: each vertex $k$ is colored by the index of the subset $B_{i}$ which it belongs to, so that all vertices in $Q_{i}$ are dyed by color $i$. By above analysis, each pair of adjacent vertices is assigned different colors, thus for each $i\in[6]$, $\{Q_{c(i)},Q_{c(i+1)}\}$ is an edge in $G_{{\mathcal{A}}_{7}}/\pi^{*}$.
Then, the sequence of edges
$\{Q_{c(1)},Q_{c(2)}\},\ \{Q_{c(2)},Q_{c(3)}\},\ \{Q_{c(3)},Q_{c(4)}\},\ \{Q_{c(4)},Q_{c(5)}\},\ \{Q_{c(5)},Q_{c(6)}\},\ \{Q_{c(6)},Q_{c(7)}\}$
corresponds to an Euler trail (i.e., a sequence of edges in which consecutive edges share a common endpoint, visiting every edge of the graph exactly once without repetition.) in the quotient graph $G_{{\mathcal{A}}_{7}}/\pi^{*}=K_{4}$.
However, by Corollary 4.1 in bondy1976graph, a connected undirected graph admits an Euler trail if and only if it has exactly 0 or $2$ vertices of odd degree. Since $K_{4}$ has $4$ vertices, each of degree $3$, all vertices have odd degrees, and thus no Euler trail exists. This contradiction shows that no partition $\pi^{*}$ can produce $K_{4}$ as a quotient graph of $G_{{\mathcal{A}}_{7}}$. ∎

### v2

@@
 Despite their widespread appearance, a comprehensive study of their computational complexity is surprisingly lacking.
-This paper aims to fill that gap by presenting several results related to the computational aspect of $U$-statistics.
-First, we derive a useful decomposition from an $m$-th order $U$-statistic to a linear combination of $V$-statistics with orders not exceeding $m$, which are generally more feasible to compute.
-Second, we explore the connection between exactly computing $V$-statistics and Einstein summation, a tool often used in computational mathematics, quantum computing, and quantum information sciences for accelerating tensor computations.
+This paper aims to fill this gap by presenting several results related to the computational aspect of $U$-statistics.
+First, we derive a useful decomposition from a $m$-th order $U$-statistic to a linear combination of $V$-statistics with orders not exceeding $m$, which are generally more feasible to compute.
+Second, we explore the connection between exactly computing $V$-statistics and Einstein summation, a tool often used in computational mathematics and quantum computing to accelerate tensor computations.
 Third, we provide an optimistic estimate of the time complexity for exactly computing $U$-statistics, based on the treewidth of a particular graph associated with the $U$-statistic kernel.
-The above ingredients lead to a new, much more runtime-efficient algorithm of exactly computing general higher-order $U$-statistics.
-We also wrap our new algorithm into an open-source Python package called u-stats.
-We demonstrate via three statistical applications that u-stats achieves impressive runtime performance compared to existing benchmarks.
-This paper aspires to achieve two goals: (1) to capture the interest of researchers in both statistics and other related areas further to advance the algorithmic development of $U$-statistics, and (2) to offer the package u-stats as a valuable tool for practitioners, making the implementation of methods based on higher-order $U$-statistics a more delightful experience.
-Keywords: $U$-statistics, Statistical/Scientific Computing, Tensors, Einstein summation, Treewidth
+The above ingredients lead to (1) a new, much more runtime-efficient algorithm to exactly compute general higher-order $U$-statistics, and (2) a more streamlined characterization of runtime complexity of computing $U$-statistics.
+We develop an accompanying open-source package called u-stats in both Python and R.
+We demonstrate through three examples in statistics that u-stats achieves impressive runtime performance compared to existing benchmarks.
+This paper also aspires to achieve two goals: (1) to capture the interest of researchers in both statistics and other related areas to further advance the algorithmic development of $U$-statistics and (2) to lift the burden of implementing higher-order $U$-statistics from practitioners.
+Keywords: $U$-statistics, Statistical Computing, Tensors, Einstein Summation, Graphs
 1 Introduction
-Higher-order $U$-statistics are prevalent in various disciplines, including statistics (robins2008higher; bonvini2022fast; bonvini2024doubly; neykov2021minimax; yao2018testing; breunig2024adaptive; shao2025u), machine learning (clemenccon2008ranking; clemenccon2013maximal; shen2025engression), random graphs (chatterjee2024higher), Bayesian networks and statistical physics (pearl1988probabilistic; koller2009probabilistic; shpitser2011efficient; li2020random), just to name a few.
-It is well-known in the statistical community that $U$-statistics are highly time-consuming to compute at large.
-In the worst-case, an $m$-th order $U$-statistic over a sample of size $n$ takes roughly $O(n^{m})$ steps to compute, because by definition (see Definition 1) one needs to sum over all $m$-out-of-$n$ permutations (resp.
-over $n^{m}$ terms).
+Higher-order $U$-statistics are prevalent in various disciplines, including statistics (robins2008higher; bonvini2022fast; bonvini2024doubly; yao2018testing; breunig2024adaptive; shao2025u), machine learning (clemenccon2008ranking; clemenccon2013maximal), random graphs (chatterjee2024higher), and Bayesian networks (shpitser2011efficient), just to name a few.
+It is well-known that $U$-statistics are highly time-consuming to compute at large.
+In the worst-case, an $m$-th order $U$-statistic over a sample of size $n$ takes roughly $O(n^{m})$ steps to compute, because by definition (see Definition 1) one needs to sum over all $m$-out-of-$n$ permutations.
 If $n=10^{4}$ and $m=7$, $n^{m}\approx 10^{28}$!
-More specifically, in causal inference, robins2008higher; robins2016technical; robins2017minimax developed a so-called “higher-order influence function (HOIF)” framework for constructing nearly rate-optimal estimators of average treatment effects under ignorability (hernan2023causal) under various structural assumptions, using higher-order $U$-statistics with order roughly of $\sqrt{\log n}$.
-Despite their appealing statistical properties, implementing such complex estimators remains challenging in practice and computational complexity is one important bottleneck.
-Here we quote a sentence from a recent graduate-level textbook on causal inference by wager2024causal (page 44) to highlight such a quandary:
-One challenge with the HOIF estimator of robins2017minimax, however, is that to date it has been challenging to implement in practical applications …
-This paper was initially motivated by (at least partially) addressing the computational aspect of this challenge, but we will present general-purpose theory and algorithm for computing general higher-order $U$-statistics beyond those for HOIF estimators.
+More specifically, in causal inference, robins2008higher developed a so-called “higher-order influence function (HOIF)” framework for constructing nearly rate-optimal estimators of average treatment effects under ignorability under various structural assumptions, using higher-order $U$-statistics with order roughly of $\sqrt{\log n}$.
+Despite their appealing statistical properties, the implementation remains a challenge in practice, especially computationally.
+Here we quote from wager2024causal to highlight this dilemma:
+One challenge with the HOIF estimator … is that to date it has been challenging to implement in practical applications …
+This point is echoed in a recent review article on the future of causal inference (cinelli2025challenges) (also see zheng2025perturbed for a related discussion):
+… where estimators are available that can improve on doubly robust style methods … the (higher-order) estimators are often computationally intensive and require careful tuning.
+There are many opportunities to make these methods more practical, automatic, and user-friendly.
+Our paper was initially motivated by (at least partially) addressing the computational aspect of this challenge, but we will present a general-purpose algorithm (and some theory) for computing generic higher-order $U$-statistics beyond those for HOIF estimators.
 We will revisit HOIF estimators in Section 4.1.
-To tackle the difficulty of computing higher-order $U$-statistics, statisticians and computer scientists have long considered the so-called “incomplete $U$-statistics”, where one sums over only a random subset of all $m$-out-of-$n$ permutations (blom1976some; janson1984asymptotic; nowicki1988subgraph; shao2025u).
+To tackle the difficulty of computing higher-order $U$-statistics, statisticians and computer scientists have long resorted to the so-called “incomplete $U$-statistics”, where one sums over only a (random) subset of all $m$-out-of-$n$ permutations (blom1976some; nowicki1988subgraph; shao2025u).
 Another natural approach is to choose a subsample of size $n^{\prime}<n$ from the whole sample and then compute a complete $U$-statistic from that subsample.
-Both strategies are akin to randomized or sketching algorithms in the theoretical computer science literature (mahoney2011randomized).
-The randomization approach certainly injects extra randomness in the final statistical results, which may result in inconsistent statistical conclusions across different randomly selected permutations/subsamples (guo2025rank).
-For a restricted class of higher-order $U$-statistics with decomposable and symmetric (or more precisely permutation-invariant) kernels, kong2018estimating cleverly recognize that exact computation can be accomplished by repeated matrix multiplications.
+The randomization approach injects extra randomness into the final statistical results, which can produce inconsistent statistical conclusions between different randomly selected permutations/subsamples (guo2025rank).
+For a restricted class of higher-order $U$-statistics with decomposable and permutation-invariant kernels, kong2018estimating astutely recognized that exact computation can be accomplished by repeated matrix multiplications.
+For asymmetric kernels, their strategy essentially computes an incomplete $U$-statistic instead.
 A somewhat similar strategy has also been considered in zhang2025adaptive, in which an incomplete $U$-statistic with a special kernel structure is computed more efficiently via dynamic programming.
-For asymmetric kernels, their strategy essentially computes an incomplete $U$-statistic instead.
-All the above approaches may suffer from efficiency loss by not using all the available information.
-The literature on (the complexity of) exactly computing higher-order $U$-statistics is rather scarce.
-This is hardly surprising because the exact time complexity of computing higher-order $U$-statistics seems to simply equal the total number of summations.
-However, as we will demonstrate in the paper, exactly computing $U$-statistics can be improved compared to a brute-force summation over all summands (using for-loops) when the kernel enjoys a particular decomposable structure specified in Definition 9 later.
-Interestingly, as we will exhibit in Section 4, such a structure holds in many statistical applications.
-It will also be clear later that it is helpful to bring in tools such as the Einstein summation from the computational mathematics community to obtain improved algorithms of computing $U$-statistics by leveraging such a structure.
-It is worth noting that Section 3 of liu2024hypothesis considered to use di-graphs to represent the structure of a class of $U$-statistics, a strategy conceptually similar to ours.
-Yet our new algorithm is designed to cover general higher-order $U$-statistics.
+All of the above approaches may suffer from efficiency loss by not using all the available information.
+[…]
 Main contributions
-The main contributions of our paper can be summarized as below.
-  • We design a new algorithmic framework for exactly computing higher-order $U$-statistics and another closely related objects called $V$-statistics by leveraging the Einstein summation operation and related concepts from combinatorics and graph theory.
-Thanks to the connection with graph theory, we can provide an optimistic estimate of the time complexity of exactly computing $U$-statistics (see Section 3.3 for details).
-Although the idea is conceptually simple, to the best of our knowledge, it has not been thoroughly explored in the existing statistical literature.
-  • Despite the ubiquity of $U$-statistics (see later this section), a dedicated, general-purpose software package for computing them seems to be lacking.
-To fill in this void, we developed a Python package called u-stats, which implemented the aforementioned algorithm.
-We demonstrate that the runtime of u-stats achieves the state-of-the-art performance in various examples, including computing HOIF estimators, motif counts of networks, and the distance covariance in the independence testing problem.
+The main contributions of our paper can be summarized as follows.
+  • We design a new method for exactly computing higher-order $U$-statistics and another related object called $V$-statistics by leveraging the Einstein summation operation and related concepts from graph theory.
+Using graph theory, we provide an optimistic estimate of the time complexity of exactly computing $U$-statistics (see Section 3.3).
+Although the idea is conceptually simple, to the best of our knowledge, it has not been thoroughly explored in the statistical literature.
+  • Despite the ubiquity of $U$-statistics, a general-purpose software package to compute them is lacking.
+To fill in this void, we developed a Python package u-stats (also available in R), which implemented the aforementioned algorithm.
+We demonstrate that the runtime of u-stats achieves the state-of-the-art performance in various examples, including computing HOIF estimators, motif counts of networks, and distance covariances in independence testing problems.
 In fact, as will be clear in Section 4, u-stats can dramatically improve the runtime compared to popular benchmarks.
-u-stats currently supports both CPU and GPU infrastructure.
+u-stats currently supports both CPU and GPU infrastructures.
 Having access to GPU can further improve the runtime because our algorithm heavily utilizes tensor operations such as the Einstein summation.
-We envision that the package u-stats can be a valuable tool for practitioners.
-Remark 1.
-We focus on scalar-valued $U$-statistics in this paper, but we plan to upgrade our package u-stats to handle vector- or matrix-valued cases in its next version.
-In this paper, one floating point operation is considered as the basic computation unit.
 Related Literature
-In this part, we provide an incomplete survey on the related literature of $U$-statistics, and the relevant techniques/tools for developing our new algorithm.
+In this part, we provide an incomplete survey of the related literature on $U$-statistics and the relevant techniques/tools to develop our new algorithm.
 Statistics and machine learning
@@
 For example, in certain machine learning tasks (clemenccon2008ranking; clemenccon2013maximal; shen2025engression), the empirical training risk can be written as a $U$-statistic instead of a sample mean.
-Various (conditional) dependence measures can also be represented as $U$-statistics (yao2018testing; neykov2021minimax; chatterjee2021new; azadkia2021simple; shao2025u; zhang2025adaptive); see Section 4.3 for an example.
+Various (conditional) dependence measures can also be represented as $U$-statistics (yao2018testing; he2021asymptotically; shao2025u; zhang2025adaptive); see Section 4.3 for an example.
 As mentioned before, in causal inference, HOIF has been shown to be a theoretically appealing framework for constructing nearly rate-optimal estimators for various causal effect parameters.
 Estimators based on HOIF are higher-order $U$-statistics; see Section 4.1 for concrete forms.
-As mentioned, due to the computational difficulty, HOIF estimators have not been comprehensively applied in real-life applications.
+Our work is also related to the literature on probabilistic graphical models (pearl1988probabilistic; shpitser2011efficient), in which a common problem is to compute certain marginal distributions from the joint distribution, factorized according to some Bayesian network.
+It can be reduced to computing $U$-statistics with kernels satisfying a particular structure to be defined later.
 Motif counts in random graphs
 Higher-order $U$-statistics also appear very often in statistical inference over networks, often modeled as random graphs (chatterjee2024higher; jin2025counting).
-For instance, chatterjee2024higher estimate motif counts (such as triangles, 4-cliques, and etc.) by higher-order $U$-statistics and use these estimated motif counts to test certain properties of the underlying network.
+For instance, chatterjee2024higher estimate motif counts (such as triangles, 4-cliques, etc.) by higher-order $U$-statistics and use these estimated motif counts to test certain properties of the underlying network.
 Motif counts also have widespread applications in bioinformatics (wernicke2006fanmod).
@@
 In particular, our algorithm utilizes the Einstein summation operation, an important tool for tensor computations.
-For example, in quantum computing/information, tensor networks are often used to describe many-body quantum systems and a very useful operation called tensor contraction is implemented by the Einstein summation in practice (markov2008simulating; stoudenmire2018learning).
+For example, in quantum computing/information, tensor networks are often used to describe many-body quantum systems and a very useful operation called tensor contraction is implemented by the Einstein summation in practice (markov2008simulating).
 As will be seen later, our proposed algorithm is similar to tensor contraction in spirit.
-Probabilistic graphical models and statistical physics
-Our work is also related to the literature on probabilistic graphical models (pearl1988probabilistic; shpitser2011efficient), in which a common problem is to compute certain marginal distributions from the joint distribution, (nested) factorized according to Bayesian network.
-It can be reduced to computing $U$-statistics with kernels satisfying the decomposable structure to be defined later.
-Similarly, in both classical (edwards1975theory; li2020random) and quantum many-body systems (bernaschi2024quantum), the corresponding log-Gibbs measure can often be written as a multi-linear sum over all particles.
-The summation is taken over all possible configurations of each particle.
-Thus the log-Gibbs measure can be viewed as (a linear combination of) higher-order $U$-statistics.
 Notation and organization
@@
 We adopt the following convention: we use the demarcation $\{\}$ for a set, representing an unordered collection of distinct elements; whereas we use the demarcation $()$ for a tuple, meaning an ordered collection of elements which may contain duplicates.
-Given a positive integer $m$, we let $m]\coloneqq\{1,2,\ldots,m\}$.
+Given a positive integer $m$, let $m]\coloneqq\{1,2,\ldots,m\}$.
 To reduce clutter, we write $\bar{i}_{m}\equiv(i_{1},\cdots,i_{m})$ for short.
-$\mathbb{N}$ denotes the set of natural numbers including zero, while $\mathbb{N}^{+}$ denotes the set of all positive integers.
-Given a tuple $\bm{s}$, we let $\mathsf{set}(\bm{s})\coloneqq\{s_{k}\mid k\in[m]\}$ denote the set containing all elements appear in $\bm{s}$.
-Given a non-empty and finite set $S$ and a positive integer $m$, all $m$-permutations of elements in $S$ is denoted by $\mathsf{Perm}(S,m)$.
-All partitions of $S$ is denoted by $\Pi_{S}$.
-If $S$ happens to be integer set $[m]$, $\Pi_{S}$ is simplified to $\Pi_{m}$.
+$\mathbb{N}$ denotes the set of natural numbers that includes zero, while $\mathbb{N}^{+}$ denotes the set of all positive integers.
+Given a tuple $\bm{s}$, we let $\mathsf{set}(\bm{s})\coloneqq\{s_{k}\mid k\in[m]\}$ denote the set containing all the elements that appear in $\bm{s}$.
+Given a non-empty and finite set $S$ and a positive integer $m$, all $m$-permutations of elements in $S$ are denoted by $\mathsf{Perm}(S,m)$.
+All partitions of $S$ are denoted by $\Pi_{S}$.
+If $S$ happens to be an integer set $[m]$, $\Pi_{S}$ is simplified to $\Pi_{m}$.
 Given sets $S_{1},\cdots,S_{m}$, $\times_{i=1}^{m}S_{i}$ denotes their Cartesian product.
@@
 Given a tuple $\bm{s}\in S^{m}$ and a tuple $\alpha\in[m]^{*}$, we let $\bm{s}[\alpha]\coloneqq(s_{\alpha_{k}})_{k=1}^{|\alpha|}$.
+To describe computational complexity, we write $\Theta(\cdot)$ to indicate equality up to a constant factor and $O(\cdot)$ to indicate an upper bound up to a constant factor.
 Terminologies from graph theory will be frequently encountered.
 Throughout this paper, all graphs are assumed to be finite and undirected without self-loops, but are allowed to have isolated vertices, unless stated otherwise.
-A graph $G$ is denoted by a pair $(V,E)$, where $V\equiv V(G)$ is the finite vertex set, and $E\equiv E(G)$ is the edge set of $G$.
+A graph $G$ is associated with a pair $(V,E)$, where $V\equiv V(G)$ denotes the set of vertices and $E\equiv E(G)$ denotes the set of edges in $G$.
 Two vertices $u$ and $v$ are said to be adjacent if $\{u,v\}\in E$.
+A complete graph is a graph with every pair of vertices connected by an edge, and a complete graph with $m$ vertices is denoted by $K_{m}$.
 For any vertex $v\in V$, its neighbor set is denoted as $N_{G}(v)\coloneqq\{u\in V\mid\{u,v\}\in E\}$, and the degree of $v$ is defined as $\deg(v,G)\coloneqq|N_{G}(v)|$.
 A graph $G^{\prime}$ is a subgraph of graph $G$, if $V(G^{\prime})\subseteq V(G)$ and $E(G^{\prime})\subseteq E(G)$.
-For any subset $V^{\prime}\subseteq V$, the induced subgraph $G_{V^{\prime}}$ has vertex set $V^{\prime}$ and edge set $E_{V^{\prime}}\coloneqq E(G_{V^{\prime}})\equiv\{\{u,v\}\in E\mid\forall\,u,v\in V^{\prime}\}$.
-A subset of vertices $C\subseteq V$ is said to be a clique if $G_{C}$ is complete, where a complete graph is simply a graph with every pair of vertices connected by an edge.
+For any subset $V^{\prime}\subseteq V$, the induced subgraph $G_{V^{\prime}}$ has a vertex set $V^{\prime}$ and an edge set $E_{V^{\prime}}\coloneqq E(G_{V^{\prime}})\equiv\{\{u,v\}\in E\mid\forall\,u,v\in V^{\prime}\}$.
+A subset of vertices $C\subseteq V$ is said to be a clique if $G_{C}$ is complete.
 The remainder of our paper is structured as follows.
 Section 2 introduces and reviews several concepts central to the paper, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation (Section 2.2), and notions (including vertex elimination, treewidth, and quotient graphs) from graph theory (Section 2.3).
-Readers who are familiar with related materials are recommended to skip the relevant subsections.
-In Section 3, the new algorithm for exactly computing $U$-statistics is presented.
-We also illustrate how the algorithm proceeds using concrete examples to facilitate understanding.
-We then demonstrate the utility of our algorithm using three examples in Section 4, including computing HOIF estimators, motif counts in networks (Section 4.2), and computing distance covariances (Section 4.3).
-In our own opinion, Section 4 is the most important part of this paper.
+Readers who are familiar with the related materials are recommended to skip the relevant subsections.
+In Section 3, the new algorithm is presented.
+We then demonstrate the practical merit of our algorithm using three numerical examples in Section 4, including the computations of HOIF estimators, motif counts in networks (Section 4.2), and distance covariances (Section 4.3).
 Section 5 concludes the paper with a discussion of future directions.
@@
 In this section, we will review concepts important for later development, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation operation (Section 2.2), and useful notions from graph theory (Section 2.3).
-The way that we define $U$- or $V$-statistics may look strange to statisticians, but we find such definitions more convenient to work with when combined with the Einstein summation formulation.
-2.1 UU/$V$-Statistics and Tensors
+The way in which we define $U$- or $V$-statistics may look strange to statisticians, but we find such definitions more convenient to work with when combined with the Einstein summation formulation.
+2.1 UU/$V$-Statistic and Tensor
 We define $U$-statistics and $V$-statistics as follows. As we just focus on computation, the sample space is regarded as a non-empty set.
-Definition 1 ($U$-statistics).
+Definition 1 ($U$-Statistic).
 Let $\mathbb{X}$ be a non-empty set, $m,n$ be positive integers satisfying $m\leq n$, and $h:\mathbb{X}^{m}\rightarrow{\mathbb{R}}$ be a function.
@@
 $\mathbb{U}(h,\bm{X})\coloneqq\sum_{\alpha\in\mathsf{Perm}([n],m)}h(\bm{X}[\alpha]).$
-Definition 2 ($V$-statistics).
+Definition 2 ($V$-Statistic).
 Let $\mathbb{X}$ be a non-empty set, $m,n$ be positive integers and $h:\mathbb{X}^{m}\rightarrow{\mathbb{R}}$ be a function.
@@
 $\mathbb{V}(h,\bm{X})\coloneqq\sum_{\alpha\in[n]^{m}}h(\bm{X}[\alpha]).$
-Definition 3 (Tensors).
+Definition 3 (Tensor).
 Given positive integers $m,n>0$, an $m$-th order tensor (i.e.
@@
 As will be clear later in this paper, $V$-statistics can be represented as summations over higher-order tensors.
-This unique property renders computing $V$-statistics, but not $U$-statistics, amenable for tensor operations such as the Einstein summation (einstein1916grundlage).
+This unique property renders computing $V$-statistics, but not $U$-statistics, amenable for tensor operations such as the Einstein summation.
 In particular, if the kernel function $h$ satisfies certain decomposable structure (see Definition 9 later), the computational complexity of $V$-statistics can be notably reduced.
 Our algorithmic framework is thus to first turn a $U$-statistic into a linear combination of $V$-statistics and then compute each individual $V$-statistic via the Einstein summation.
-2.2 The Einsum Notation and Einsum Operation
+2.2 Einsum Notation and Einsum Operation
 As mentioned, our algorithm to be revealed will rely on the Einstein summation operation (abbreviated as the Einsum operation) over tensors.
 In this section, we give a brief but (hopefully) pedagogical overview for readers unfamiliar with these related concepts.
-We also steer readers to klaus2023compiling or blacher2024einsum for more details.
-The Einsum operation to be defined serves as a fundamental computational primitive in our proposed algorithm for efficiently and exactly computing UU/$V$-statistics.
-The Einsum operation generalizes the classical Einstein summation convention introduced in (einstein1916grundlage).
+We also steer readers to blacher2024einsum for more details.
+The Einsum operation to be defined serves as a fundamental computational primitive in our proposed algorithm to efficiently and exactly compute the UU/$V$-statistics.
+The Einsum operation generalizes the classical Einstein summation convention.
 The modern Einsum operation, now implemented in popular Python packages such as numpy and pytorch, can be used for efficient tensor computation.
 To explain the Einsum operation, we first introduce a concept called the Einsum notation.
-Definition 4 (The Einsum Notation).
+Definition 4 (Einsum Notation).
 Let $m$ and $K$ be two positive integers. An Einsum notation with respect to [w.r.t.] $[m]$ is a pair
@@
 where we recall that $m]^{\ast}$ denotes all tuples possibly generated by elements in $[m]$, and $B$ needs to satisfy the following constraint:
-  • $B[i]\neq B[j]$, for any $i,j\in[|B|]suchthati\neq j$,
+  • $B[i]\neq B[j]$, for any $i,j\in[|B|]:i\neq j$,
   • $\mathsf{set}(B)\subseteq\bigcup_{k\in[K]}\mathsf{set}(A_{k})$.
 $\mathcal{A}\coloneqq(A_{k})_{k=1}^{K}$ is referred to as the input tuple list of $\mathcal{N}$ and $B$ is referred as the output tuple.
-The index set of $\mathcal{N}$} is the set of all indices appearing in the input tuple list, and is denoted by $\mathsf{Indices}({\mathcal{N}})$.
-Without loss of generality, we always take $m$ to be the smallest $m$ such that $m]=\mathsf{Indices}({\mathcal{N}})$.
+The index set of $\mathcal{N}$ is the set of all indices appearing in the input tuple list, and is denoted by $\textsc{Indices}(\mathcal{N})$.
+Without loss of generality, we always take $m$ to be the smallest $m$ such that $m]=\textsc{Indices}(\mathcal{N})$.
 If $B=\emptyset$, we then simply write $\mathcal{A}$ as $\mathcal{N}=(\mathcal{A};\emptyset)$ and call $\mathcal{A}$ as the Einsum notation.
-It is noteworthy that in the above definition, the index set $[m]$ is simply chosen for notational convenience and in principle the indices do not have to be integers or even numbers.
+It should be noted that, in the above definition, the index set $[m]$ is simply chosen for notational convenience and, in principle, the indices need not be integers or numbers.
 We now define the Einsum operation based on the Einsum notation.
-Definition 5 (The Einsum Operation).
+Definition 5 (Einsum Operation).
 Let $\mathcal{N}=(\mathcal{A};B)$ be an Einsum notation w.r.t.
@@
 $m_{k}=|A_{k}|$ for any $k\in[K]$.
-An Einsum operation, denoted as $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$, takes $\mathcal{T}$} and $\mathcal{N}$} as inputs and outputs another tensor $T^{\rm out}:[n]^{|B|}\rightarrow{\mathbb{R}}$ as follows: for any $\alpha\in[n]^{|B|}$,
+An Einsum operation, denoted as $\textsc{Einsum}(\mathcal{T},\mathcal{N})$, takes $\mathcal{T}$ and $\mathcal{N}$ as inputs and outputs another tensor $T^{\rm out}:[n]^{|B|}\rightarrow{\mathbb{R}}$ as follows: for any $\alpha\in[n]^{|B|}$,
 $T^{\rm out}(\alpha)=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{m}\\ \alpha^{\prime}[B]=\alpha\end{subarray}}\prod_{k=1}^{K}T_{k}(\alpha^{\prime}[A_{k}]).$
@@
 $T^{\rm out}=\sum_{\alpha^{\prime}\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha^{\prime}[A_{k}]).$
-We now use several concrete examples to illustrate the above definitions.
-Example 1.
-We take matrix multiplication as an example to illustrate the Einsum operation just defined.
-Let $T_{1},T_{2}$ be two input matrices, which are also 2nd-order tensors.
-Therefore ${\mathcal{T}}=(T_{1},T_{2})$.
-Suppose we are interested in multiplying $T_{1}$ and $T_{2}$, denoted as $T^{\rm out}$: for any $i,j\in[n]$,
-$T^{\rm out}(i,j)=\sum_{k=1}^{n}T_{1}(i,k)T_{2}(k,j).$
-The matrix multiplication corresponds to the Einsum operation with Einsum notation ${\mathcal{N}}=({\mathcal{A}};B)$, with input tuple list ${\mathcal{A}}=((1,2),(2,3))$ and output tuple $B=(1,3)$ and $|B|=2$.
-The index set is simply $I=[3]$.
-This means that the operation sums over the dimensions of $T_{1}$ and $T_{2}$ corresponding to index $2$ and reserves indices 11 and $3$.
-$T^{\rm out}(\alpha)$ for any $\alpha=(i,j)\in[n]^{2}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ because by definition:
-$\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ $\equiv\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}\prod_{k=1}^{2}T_{k}(\alpha^{\prime}[A_{k}])$
-$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])$
-$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[1,2])T_{2}(\alpha^{\prime}[2,3])$
-$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[3])$
-$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha[3])$
-$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(i,\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],j),$
-so the claim holds by identifying $\alpha^{\prime}[2]=k$.
-Example 2.
-Here we consider an example involving higher-order tensors.
-Given three tensors ${\mathcal{T}}=(T_{1},T_{2},T_{3})$, each of size $n$, where $T_{1}$ and $T_{2}$ are 3rd-order tensors while $T_{3}$ is a matrix, suppose that we are interested in computing the following 3rd-order tensor: for any $(i,j,k)\in[n]^{3}$
-$T^{\rm out}(i,j,k)=\sum_{a=1}^{n}\sum_{b=1}^{n}\sum_{c=1}^{n}T_{1}(i,a,b)T_{2}(a,c,j)T_{3}(b,k).$
-$T^{\rm out}(i,j,k)$ is in fact a partial trace, a common operation taken in quantum mechanics.
-The corresponding Einsum notation is ${\mathcal{N}}=({\mathcal{A}};B)$ with ${\mathcal{A}}=(A_{1}=(1,2,3),A_{2}=(2,4,5),A_{3}=(3,6))$ with $B=(1,5,6)$ and the index set $I=[6]$.
-$T^{\rm out}(\alpha)$ for any $\alpha=(i,j,k)\in[n]^{3}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ because by definition:
-$\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ $\equiv\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])T_{3}(\alpha^{\prime}[A_{3}])$
-$=\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
-$=\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
-$=\sum_{\alpha^{\prime}[2]=1}^{n}\sum_{\alpha^{\prime}[3]=1}^{n}\sum_{\alpha^{\prime}[4]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2],\alpha^{\prime}[3])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[4],\alpha[5])T_{3}(\alpha^{\prime}[3],\alpha[6]),$
-so the claim holds by identifying $\alpha^{\prime}[2]=a$, $\alpha^{\prime}[3]=b$, and $\alpha^{\prime}[4]=c$.
+[…]
 Example 1.
@@
 The corresponding Einsum notation is $\mathcal{A}=(A_{1}=(1,2),A_{2}=(2,3),A_{3}=(3,4))$ with $B=\emptyset$ and the index set $I=[4]$.
-$T^{\rm out}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ because by definition:
-$\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ $\equiv\sum_{\alpha\in[n]^{4}}\prod_{k=1}^{3}T_{k}(\alpha[A_{k}])=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[A_{1}])T_{2}(\alpha[A_{2}])T_{3}(\alpha[A_{3}])$
+$T^{\rm out}$ can be written in the Einsum operation form $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ because by definition:
+$\textsc{Einsum}(\mathcal{T},\mathcal{A})$ $\equiv\sum_{\alpha\in[n]^{4}}\prod_{k=1}^{3}T_{k}(\alpha[A_{k}])=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[A_{1}])T_{2}(\alpha[A_{2}])T_{3}(\alpha[A_{3}])$
 $=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[1,2])T_{2}(\alpha[2,3])T_{3}(\alpha[3,4])$
@@
 The Einsum operation has been implemented in highly optimized Python libraries such as numpy.einsum and pytorch.einsum.
-These efficient algorithmic subroutines, enabling efficient Einsum operation in practice, serve as the building blocks for our algorithm.
+These efficient algorithmic subroutines, enabling efficient Einsum operation in practice, serve as the building blocks of our algorithm.
 Notably, developed for training large overparameterized neural networks in tensor forms, pytorch.einsum is also designed to leverage GPU computing to further enhance the scalability and speed of the Einsum operation, as we will demonstrate using an example from statistics in Section 4.1.
@@
 The Einsum operation is frequently employed in scientific computing problems.
-For example, the Einsum operation serves as a basic algorithmic subroutine for computation over tensor (hyper)networks, a useful mathematical language adopted by the quantum computing and quantum information community (markov2008simulating; stoudenmire2018learning).
+For example, the Einsum operation serves as a basic algorithmic subroutine for computation over tensor (hyper)networks, a useful mathematical language adopted by the quantum computing and quantum information community (markov2008simulating).
 Although we do not directly adopt the tensor network formulation, our strategy is quite similar in that we also establish a correspondence between the Einsum operation and operations on graphs – see Section 3.2 for more details.
 2.3 Useful Concepts from Graph Theory
-In this section, we recall three key concepts from graph theory that are central to our complexity-theoretic characterization: vertex elimination and treewidth, which captures the dominant computational complexity; and quotient graph, which is related to the decomposition of a $U$-statistic into a linear combination of $V$-statistics.
-Here we adopt the terminologies from bodlaender2010treewidth.
+In this section, we recall three key concepts from graph theory that are central to our complexity-theoretic characterization.
+We follow the terminology of bodlaender2010treewidth for vertex elimination and treewidth, which captures the dominant computational complexity of our problem.
+We also recall the notion of quotient graph, adopted from sanders2012high, which underlies the decomposition of a $U$-statistic into a linear combination of $V$-statistics.
 Definition 6 (Vertex Elimination).
@@
 $E_{\pi}=\left\{\{B_{i},B_{j}\}\subseteq\pi\mid\exists\,v_{1}\in B_{i},v_{2}\in B_{j},\text{ such that }\{v_{1},v_{2}\}\in E\right\}.$
-3 A New Algorithm of Exactly Computing $U$-Statistics
+3 A New Algorithm of Computing $U$-Statistics
 We are ready to present our new algorithm for exactly computing $U$-statistics, along with an analysis of the time complexity.
-Our algorithm has been implemented in an open source Python package u-stats.
-We begin by introducing the multiplicative-decomposable (MD) structure of the UU/$V$-statistic kernel in Section 3.1.
-We then describe our algorithm and the complexity analysis for $V$-statistics in Section 3.2, followed by those for $U$-statistics in Section 3.3.
+Our algorithm has been implemented in an open source Python package u-stats (and its R interface).
+We begin by introducing the multiplicative-decomposable (MD) structure of the UU/$V$-statistic kernel in Section 3.1, followed by the new algorithm and complexity analysis for $V$- and $U$-statistics in Section 3.2 and Section 3.3.
 3.1 The Multiplicative-Decomposable Kernels
 […]
-Definition 9 (Multiplicative Decomposition of a Function).
+Definition 9 (Multiplicative Decomposition).
 Let $\mathbb{X}$ be a non-empty set, $m$ a positive integer, and $h:\mathbb{X}^{m}\to{\mathbb{R}}$.
 Suppose $\mathcal{A}\coloneqq(A_{k})_{k=1}^{K}\in[m]^{**}$ satisfies $\bigcup_{k=1}^{K}\mathsf{set}(A_{k})=[m]$, and let $\mathcal{H}\coloneqq(h_{k})_{k=1}^{K}$ be a tuple of functions where each $h_{k}$ is defined on $\mathbb{X}^{|A_{k}|}$.
-${\mathcal{H}}\times{\mathcal{A}}$, comprised of ordered pairs of the form $(h_{k},A_{k})$ for $k\in[K]$, is called a multiplicative decomposition of the function $h$, if for any $\bm{x}\in{\mathbb{X}}^{m}$,
+$\mathcal{H}\times\mathcal{A}$, consisting of ordered pairs of the form $(h_{k},A_{k})$ for $k\in[K]$, is called a multiplicative decomposition of the function $h$, if for any $\bm{x}\in\mathbb{X}^{m}$,
 $h(\bm{x})=\prod_{k=1}^{K}h_{k}\big(\bm{x}[A_{k}]\big).$
-In the sequel, we refer to the function tuple $\mathcal{H}$} as the kernel components and the index tuple $\mathcal{A}$} as the decomposition signature.
-For convenience, we say that a kernel $h$ permits a (nontrivial) MD structure ${\mathcal{H}}\times{\mathcal{A}}$ when $K>1$.
-$U$-statistic kernels with MD structures are pervasive in statistical applications. We will provide three examples in Section 4.
-To utilize Einsum operation to compute UU/$V$-statistics, we need to store pieces of the underlying statistics in tensor format.
+[…]
+To utilize the Einsum operation to compute UU/$V$-statistics, we need to store pieces of the underlying statistics in tensor format.
 We term this step as tensorization.
@@
 $T_{k}(\alpha_{k})=h_{k}(\bm{X}[\alpha_{k}]),\quad\forall\,\alpha_{k}\in[n]^{|A_{k}|},\,\forall\,k\in[K].$
-Then following Definition 2 and Definition 9, we have
+Then following Definition 2 and Definition 9,
 $\mathbb{V}(h,\bm{X})=\sum_{\alpha\in[n]^{m}}h(\bm{X}[\alpha])\equiv\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}h_{k}(\bm{X}[\alpha[A_{k}]])\equiv\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha[A_{k}]),$ (1)
@@
 3:A sample $\bm{X}=(X_{1},\dots,X_{n})\in\mathbb{X}^{n}$ from a non-empty set $\mathbb{X}$
-4:A tuple of tensors ${\mathcal{T}}\coloneqq(T_{k}:[n]^{m_{k}}\to{\mathbb{R}})_{k=1}^{K}$, where $T_{k}$ is the tensorized version of $h_{k}$
+4:A tuple of tensors $\mathcal{T}\coloneqq(T_{k}:[n]^{m_{k}}\to{\mathbb{R}})_{k=1}^{K}$, where $T_{k}$ is the tensorized $h_{k}$
 5:function Tensorization($\mathcal{H},\bm{X}$)
@@
 The $V$-statistic and $U$-statistic with kernel $h$ are respectively:
-$\mathbb{V}(h,\bm{X})=\mathbb{V}({\mathcal{T}},{\mathcal{A}})=\textsf{Einsum}{}({\mathcal{T}},{\mathcal{A}}),\quad\mathbb{U}(h,\bm{X})=\mathbb{U}({\mathcal{T}},{\mathcal{A}}).$
-3.2 The Algorithm and Time Complexity of Computing $V$-Statistics
+$\mathbb{V}(h,\bm{X})=\mathbb{V}(\mathcal{T},\mathcal{A})=\textsc{Einsum}{}(\mathcal{T},\mathcal{A}),\quad\mathbb{U}(h,\bm{X})=\mathbb{U}(\mathcal{T},\mathcal{A}).$
+3.2 The Algorithm of Computing $V$-Statistics
 In this section, we show that $V$-statistics can be computed via the Einsum operation.
@@
 By definition, $\mathcal{A}$ constitutes an Einsum notation and $\mathcal{T}$ can be represented by $\mathcal{A}$ (see Definition 5), then we have
-$\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})=\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha[A_{k}]).$
+$\textsc{Einsum}(\mathcal{T},\mathcal{A})=\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha[A_{k}]).$
 Comparing the above equation with (1), we immediately have
-$\mathbb{V}(h,\bm{X})=\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}),$
+$\mathbb{V}(h,\bm{X})=\textsc{Einsum}(\mathcal{T},\mathcal{A}),$
 i.e. computing a $V$-statistic is the same as performing an Einsum operation. Algorithm 2 describes how $V$-statistics can be computed by the Einsum operation.
@@
 6:  $\mathcal{T}\leftarrow$ Tensorization($\mathcal{H},\bm{X}$) $\triangleright$ see Algorithm 1
-7:  $v\leftarrow\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ [^1] In our implementation of u-stats, we directly call numpy.einsum or pytorch.einsum, with the summation ordering optimized by opt-einsum.
+7:  $v\leftarrow\textsc{Einsum}(\mathcal{T},\mathcal{A})$ [^1] In our implementation of u-stats, we directly call numpy.einsum or pytorch.einsum, with the summation ordering optimized by opt-einsum.
 8:  return $v$
@@
 The time complexity can in turn be characterized by the treewidth (Definition 7) of that graph.
-Similar idea has also been exploited in markov2008simulating, in which the time complexity of as the so-called tensor contraction operation (a special case of the Einsum operation) is also characterized via the treewidth of the underlying tensor network.
+Similar idea has also been exploited in markov2008simulating, in which the time complexity of tensor contraction (a special case of the Einsum operation) is characterized by the treewidth of the underlying tensor network.
 To facilitate presentation, we first introduce the notion of “decomposition graph” associated with a given Einsum notation, which captures all the information needed to establish the time complexity of computing a particular $V$-statistic.
-Definition 10 (The Decomposition Graph Associated with an Einsum Notation).
+Definition 10 (Decomposition Graph).
 Let $\mathcal{A}$ be an Einsum notation with no output.
 The decomposition graph associated with $\mathcal{A}$ (or equivalently the decomposition signature $\mathcal{A}$) is a graph $G_{\mathcal{A}}=(V_{\mathcal{A}},E_{\mathcal{A}})$ where
-  • $V_{{\mathcal{A}}}=\mathsf{Indices}({\mathcal{A}})$,
+  • $V_{\mathcal{A}}=\textsc{Indices}(\mathcal{A})$,
   • $E_{\mathcal{A}}=\big\{\{i,j\}\subseteq V_{\mathcal{A}}\mid\exists\,A\in\mathcal{A};\text{ such that }i,j\in\mathsf{set}(A)\big\}$.
 We use several concrete examples to illustrate the above definition.
-Example 4 (Correspondence between the Einsum notation and the decomposition graph).
+Example 2 (Einsum notations and its decomposition graphs).
 Figure 1 below illustrates several examples of decomposition graphs associated with different decomposition signatures.
@@
 $(1,2,4),(2,3,4))$
-Figure 1: Examples of decomposition graphs associated with different decomposition signatures.
-We are ready to present the following “optimistic” estimate of the time complexity of the Einsum operation. The proof of the proposition below is deferred to Appendix A.
+Figure 1: Examples of decomposition graphs of different decomposition signatures.
+[…]
 Proposition 1.
 Let $\mathcal{A}$ be an Einsum notation with no output, and $G_{\mathcal{A}}$ be the corresponding decomposition graph of $\mathcal{A}$.
-Then there exists an algorithm with a particular Einsum ordering such that for any set of tensors $\mathcal{T}$} that can be represented by the Einsum notation $\mathcal{A}$}, the time complexity of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(|{\mathcal{A}}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$.
-Remark 4 (The Einsum ordering in practice).
+Then there exists an algorithm with a particular Einsum ordering such that for any set of tensors $\mathcal{T}$ that can be represented by the Einsum notation $\mathcal{A}$, the time complexity of $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ is $O(|\mathcal{A}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$.
+Remark 3 (Einsum summation ordering in practice).
 We say that the time complexity given in the above proposition is only an optimistic estimate, because although it can be achieved with the optimal Einsum ordering corresponding to the treewidth (Definition 7), this particular ordering is usually difficult to find in practice.
 arnborg1987complexity have shown that finding the ordering that gives rise to the treewidth is an NP-complete problem.
-While it is ideal to evaluate $\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ in Algorithm 2 using this optimal ordering, we decide to directly plug in the Python library opt-einsum (daniel2018opt) to find the ordering in practice, because such a preexisting library has been optimized and stabilized over the years.
+While it is ideal to evaluate $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ in Algorithm 2 using this optimal ordering, we decide to directly plug in the Python library opt-einsum (daniel2018opt) to find the ordering in practice, because such a preexisting library has been optimized and stabilized over the years.
 opt-einsum offers both heuristic strategies such as greedy algorithms or exhaustive search such as dynamic programming, but the latter can be time-consuming when dealing with large-scale problems.
@@
 Since the ordering only depends on the MD structure of the kernel $h$, for commonly encountered kernels, one could in principle find the optimal ordering via exhaustive search without accessing the data sample and re-use this ordering whenever the same type of kernel appears.
-Finally, we briefly remark that opt-einsum uses a different scheme to describe the summation ordering from the one used in this paper.
-We will not describe such a difference in any detail as doing so will necessarily introduce extra notations and definitions that are only tangential to the main theme of the paper.
 With Proposition 1, we immediately obtain the following optimistic estimate of the time complexity of exactly computing a $V$-statistic.
@@
 Then there exists an algorithm such that its time complexity of computing $\mathbb{V}(h,\bm{X})$ exactly is $O(|\mathcal{A}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$.
-3.3 The Algorithm and Time Complexity of Computing $U$-Statistics
+3.3 The Algorithm of Computing $U$-Statistics
 As discussed in Section 2.1, unlike $V$-statistics, $U$-statistics cannot be directly computed via the Einsum operation.
@@
 The cardinality of the given partition $\pi$ is referred to as the order of $\mathcal{V}(n,\pi)$.
-Definition 12 (Restricted $V$-statistics).
+Definition 12 (Restricted $V$-Statistic).
 Let $\mathbb{X}$ be a non-empty set, $m,n$ be positive integers, $\pi$ be a partition of set $[m]$ and $h:\mathbb{X}^{m}\rightarrow{\mathbb{R}}$ be a function defined on $\mathbb{X}^{m}$.
@@
 Having introduced the above concepts, we state the following lemma that describes how to decompose a $U$-statistic into a linear combination of $V$-statistics.
-Lemma 1 (Decomposition of $U$-statistics into $V$-statistics).
+Lemma 1 (Decomposition of $U$-Statistic into $V$-Statistics).
 Let $\mathbb{X}$ be a non-empty set and $h:\mathbb{X}^{m}\rightarrow{\mathbb{R}}$ be a kernel function. Then for any tuple $\bm{X}\in\mathbb{X}^{n}$ with $n\geq m$,
 $\mathbb{U}(h,\bm{X})=\sum_{\pi\in\Pi_{m}}\mu_{\pi}\mathbb{V}[\pi](h,\bm{X}),$
-where
-$\mu_{\pi}=(-1)^{(m-|\pi|)}\prod_{C\in\pi}(|C|-1)!.$
-The proof of the above lemma can be found in Appendix B and leverages Möbius inversion from combinatorics.
-We now discuss how to extend the tensorization algorithm from $V$-statistics (Algorithm 1) to $U$-statistics.
-Let $\mathbb{U}(h,\bm{X})$ be an $m$-th order $U$-statistic with kernel function $h$ and sample $\bm{X}\in{\mathbb{X}}^{n}$, where $h$ permits a MD structure ${\mathcal{H}}\times{\mathcal{A}}$.
-On the surface, it seems that we have to apply Algorithm 2 to each resulting $V$-statistics in the decomposition given in Lemma 1.
-But in fact, we only need to apply Algorithm 2 once since all the resulting $V$-statistics can be tensorized by the same tensor tuple ${\mathcal{T}}=\textsc{Tensorizaion}({\mathcal{H}},\bm{X})$.
+where $\mu_{\pi}=(-1)^{(m-|\pi|)}\prod_{C\in\pi}(|C|-1)!$.
+[…]
 Given any $\pi\in\Pi_{m}$, we can also compute the restricted $V$-statistic $\mathbb{V}[\pi](h,\bm{X})$ by turning it into a new Einsum notation $\mathcal{A}_{\pi}$ constructed from $\mathcal{A}$ based on $\pi$.
@@
 We then have the following identity that relates the restricted $V$-statistic by the partition $\pi$ and the Einsum operation under the induced decomposition signature $\mathcal{A}_{\pi}$:
-$\mathbb{V}[\pi](h,\bm{X})=\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}_{\pi}).$ (2)
+$\mathbb{V}[\pi](h,\bm{X})=\textsc{Einsum}(\mathcal{T},\mathcal{A}_{\pi}).$ (2)
 One can immediately realize that (2) holds because, by Definitions 11 and 12, $\mathbb{V}[\pi](h,\bm{X})$ is a summation over all tuples of $n]^{m}$, under the restriction that indices belonging to the same subset of the partition $\pi$ need to be identified.
@@
 1:
-2:An Einsum notation ${\mathcal{A}}=(A_{k})_{k=1}^{K}$ with index set $[m]$
-3:A partition $\pi\in\Pi_{m}$ of size $J$
-4:The Einsum notation ${\mathcal{A}}_{\pi}$ induced by $\pi$ on $\mathcal{A}$}
-5:function InducedNotation(${\mathcal{A}};\pi$)
-6:  Let $(Q_{j})_{j=1}^{J}$ be a permutation of $\pi$ $\triangleright$ Label subsets by $j\in[J]$
-7:  Define $\psi\colon[m]\to[J]$ as the subset assignment map:
-8:  function $\psi$(ii)
-9:   return $j$ if $i\in Q_{j}$ $\triangleright$ Map all indices in $Q_{j}$ to $j$
-10:  end function
-11:  Initialize ${\mathcal{A}}_{\pi}\leftarrow{\mathcal{A}}$
-12:  for $k\in[K]$ do
-13:   for $i\in[|A_{k}|]$ do
-14:     ${\mathcal{A}}_{\pi}[k][i]\leftarrow\psi(A_{k}[i])$ $\triangleright$ Replace index with its subset label
+2:An Einsum notation $\mathcal{A}=(A_{k})_{k=1}^{K}$ with index set $[m]$ and a partition $\pi\in\Pi_{m}$ of size $J$
+3:The Einsum notation $\mathcal{A}_{\pi}$ induced by $\pi$ on $\mathcal{A}$
+4:function InducedNotation($\mathcal{A};\pi$)
+5:  Let $(Q_{j})_{j=1}^{J}$ be a permutation of $\pi$ $\triangleright$ Label subsets by $j\in[J]$
+6:  Define $\psi\colon[m]\to[J]$ as the subset assignment map:
+7:  function $\psi$(ii)
+8:   return $j$ if $i\in Q_{j}$ $\triangleright$ Map all indices in $Q_{j}$ to $j$
+9:  end function
+10:  Initialize $\mathcal{A}_{\pi}\leftarrow\mathcal{A}$
+11:  for $k\in[K]$ do
+12:   for $i\in[|A_{k}|]$ do
+13:     $\mathcal{A}_{\pi}[k][i]\leftarrow\psi(A_{k}[i])$ $\triangleright$ Replace index with its subset label
+14:   end for
 15:  end for
-16:  end for
-17:  return ${\mathcal{A}}_{\pi}$
-18:end function
+16:  return $\mathcal{A}_{\pi}$
+17:end function
 According to Lemma 1, when computing an $m$-th order $U$-statistic, the number of $V$-statistics that we need to compute is the same as the total number of partitions of $[m]$, or equivalently the Bell number $\mathsf{B}_{m}$.
@@
 $\begin{split}\widetilde{T}_{k}(\alpha)=\begin{cases}h_{k}(\bm{X}[\alpha])&\text{if }\forall\,i,j\in[|A_{k}|]:i\neq j,\,\alpha[i]\neq\alpha[j]\\ 0&\text{otherwise},\end{cases}\,\forall\,\alpha\in[n]^{|A_{k}|},\,\forall\,k\in[K].\end{split}$ (3)
-Then it holds that
-$\mathbb{U}({\mathcal{T}},{\mathcal{A}})=\mathbb{U}(\widetilde{{\mathcal{T}}},{\mathcal{A}}).$
+Then $\mathbb{U}(\mathcal{T},\mathcal{A})=\mathbb{U}(\widetilde{\mathcal{T}},\mathcal{A})$.
 The proof of Lemma 2 is straightforward as none of the indices $\alpha$ such that $\widetilde{T}_{k}(\alpha)\neq T_{k}(\alpha)$ (when the second case of (3) holds) contributes to the value of the underlying $U$-statistic, because by definition $U$-statistic sums over distinct indices.
@@
 In Table 4 of Section 4.1, we use higher-order influence functions (HOIFs) mentioned in the beginning of this paper as a realistic example to also exhibit the effectiveness of this sparsification trick in reducing the number of $V$-statistics needed to compute.
-With the sparsification trick, the final algorithm of computing $U$-statistics is displayed in Algorithm 4 below.
+Finally, we present the new algorithm for computing $U$-statistics in Algorithm 4 below.
 Algorithm 4 Computation of $U$-Statistics via Tensorization
@@
 5:function Ustats($\mathcal{T},\mathcal{A};\bm{X}$)
-6:  Let $\widetilde{{\mathcal{T}}}$ be constructed as in Lemma 2
-7:  Initialize $u\leftarrow 0$
-8:  for $\pi\in\Pi_{m}$ do
-9:   if $\forall Q\in\pi,\forall A\in{\mathcal{A}};|Q\cap\mathsf{set}(A)|<2$ then
-10:     $\mu_{\pi}\leftarrow(-1)^{m-|\pi|}\prod_{C\in\pi}(|C|-1)!$ $\triangleright$ See Lemma 1
-11:     ${\mathcal{A}}_{\pi}\leftarrow$ InducedNotation(${\mathcal{A}};\pi$) $\triangleright$ See Algorithm 3
-12:     $v_{\pi}\leftarrow\mathsf{Einsum}(\widetilde{{\mathcal{T}}},{\mathcal{A}}_{\pi})$ (see footnote 1)
-13:     $u\leftarrow u+\mu_{\pi}v_{\pi}$
-14:   end if
-15:  end for
-16:  return $u$
-17:end function
+6:  Let $\widetilde{\mathcal{T}}$ be constructed as in Lemma 2 and initialize $u\leftarrow 0$
+7:  for $\pi\in\Pi_{m}$ do
+8:   if $\forall Q\in\pi,\forall A\in\mathcal{A};|Q\cap\mathsf{set}(A)|<2$ then
+9:     $\mu_{\pi}\leftarrow(-1)^{m-|\pi|}\prod_{C\in\pi}(|C|-1)!$ $\triangleright$ See Lemma 1
+10:     $\mathcal{A}_{\pi}\leftarrow$ InducedNotation($\mathcal{A};\pi$) $\triangleright$ See Algorithm 3
+11:     $v_{\pi}\leftarrow\textsc{Einsum}(\widetilde{\mathcal{T}},\mathcal{A}_{\pi})$ $\triangleright$ See Footnote 1
+12:     $u\leftarrow u+\mu_{\pi}v_{\pi}$
+13:   end if
+14:  end for
+15:  return $u$
+16:end function
 Example 1(b).
@@
 $\mathbb{U}(\widetilde{\mathcal{T}},\mathcal{A})$ $=\mathbb{V}[\{1\}\{2\}\{3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})-\cancel{\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})}-\mathbb{V}[\{1,3\}\{2\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})$
-$-\mathbb{V}[\{1,4\}\{2\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})-\cancel{\mathbb{V}[\{2,3\}\{1\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}-\mathbb{V}[\{2,4\}\{1\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})-\cancel{\mathbb{V}[\{3,4\}\{1\}\{2\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
-$+\cancel{\mathbb{V}[\{1,2\}\{3,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\mathbb{V}[\{1,3\}\{2,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})+\cancel{\mathbb{V}[\{1,4\}\{2,3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
-$+\cancel{2\mathbb{V}[\{1,2,3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\cancel{2\mathbb{V}[\{1,2,4\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\cancel{2\mathbb{V}[\{1,3,4\}\{2\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
-$+\cancel{2\mathbb{V}[\{2,3,4\}\{1\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}-\cancel{\mathbb{V}[\{1,2,3,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}.$
-As shown in the above display, there were originally 15 $V$-statistics in total, whereas only 5 $V$-statistics are left after applying the sparsification trick!
+$-\mathbb{V}[\{1,4\}\{2\}\{3\}](\widetilde{\mathcal{T}},\mathcal{A})-\cancel{\mathbb{V}[\{2,3\}\{1\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})}-\mathbb{V}[\{2,4\}\{1\}\{3\}](\widetilde{\mathcal{T}},\mathcal{A})$
+$-\cancel{\mathbb{V}[\{3,4\}\{1\}\{2\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{\mathbb{V}[\{1,2\}\{3,4\}](\widetilde{\mathcal{T}},\mathcal{A})}+\mathbb{V}[\{1,3\}\{2,4\}](\widetilde{\mathcal{T}},\mathcal{A})$
+$+\cancel{\mathbb{V}[\{1,4\}\{2,3\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{2\mathbb{V}[\{1,2,3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{2\mathbb{V}[\{1,2,4\}\{3\}](\widetilde{\mathcal{T}},\mathcal{A})}$
+$+\cancel{2\mathbb{V}[\{1,3,4\}\{2\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{2\mathbb{V}[\{2,3,4\}\{1\}](\widetilde{\mathcal{T}},\mathcal{A})}-\cancel{\mathbb{V}[\{1,2,3,4\}](\widetilde{\mathcal{T}},\mathcal{A})}.$
+As shown in the above display, there were originally 15 non-redundant $V$-statistics in total, whereas only 5 $V$-statistics remain after applying the sparsification trick!
 Taking the second term $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})$ on the RHS of the above display as an example: $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})\equiv 0$ because the intersection between $Q=\{1,2\}$ and $A_{1}=(1,2)$ is 2, violating the criterion described in (4).
-Armed with the above preparations, we can bound the time complexity of exactly computing $U$-statistics using Proposition 1.
 Given a partition $\pi$ and the decomposition signature $\mathcal{A}$, the decomposition graph corresponding to induced Einsum notation $\mathcal{A}_{\pi}$ can be characterized by the quotient graph $G_{\mathcal{A}}/\pi$ (see Definition 8) of the decomposition graph $G_{\mathcal{A}}$ induced by $\pi$.
+Armed with the above observation, we can bound the time complexity of exactly computing $U$-statistics by combining Corollary 1, Lemma 1, and Lemma 2, leading to the following result.
 Corollary 2.
@@
 Under the same conditions of Corollary 1, there exists an algorithm such that the time complexity of computing the $U$-statistic $\mathbb{U}(h,\bm{X})$ is $O(|\mathcal{A}|\sum_{l=1}^{M}N_{l}n^{l+1})$, where
-$M$ $\coloneqq\max\{\mathsf{tw}(G_{\mathcal{A}}/\pi)\mid\pi\in\Pi_{m}^{\mathcal{A}}\},$
-$N_{l}$ $\coloneqq|\{\pi\in\Pi_{m}^{{\mathcal{A}}}\mid\mathsf{tw}(G_{{\mathcal{A}}}/\pi)=l\}|,\;\forall\,l\in[M].$
+$M\coloneqq\max\{\mathsf{tw}(G_{\mathcal{A}}/\pi)\mid\pi\in\Pi_{m}^{\mathcal{A}}\},\quad N_{l}\coloneqq|\{\pi\in\Pi_{m}^{\mathcal{A}}\mid\mathsf{tw}(G_{\mathcal{A}}/\pi)=l\}|,\;\forall\,l\in[M].$
 and $\Pi_{m}^{\mathcal{A}}$ is defined in (4).
 Remark 5.
-In the above result, quantities $N_{l}$ and $M$ are introduced.
-$N_{l}$ is the number of partitions in $\Pi_{m}^{{\mathcal{A}}}$ such that the treewidth of the decomposition graph $G_{{\mathcal{A}}}$ quotient by $\pi$ is $l$, is defined.
-Although abstract, $N_{l}$ and $M$ can in principle be computed for concrete examples.
-For instance, in Section 4.1, we report the values of $M$ for HOIF estimators between orders $2$ to $12$; see the last column of Table 1.
-At this moment, we are not yet aware of a good strategy to count $N_{l}$, except a loose upper bound by the total number of $V$-statistics in the decomposition; see the third column of Table 1.
+[…]
 Remark 6.
-In this paper, we have used the treewidth of the decomposition graph (or its quotient as in Corollary 2) to characterize the optimistic time-complexity of exactly computing UU/$V$-statistics, and the Einsum operation.
-Although treewidth has been a widely used concept for establishing the time-complexity of algorithms on graphs (dong2021nearly), it is in general nontrivial to compute or derive a tight bound.
-To our knowledge, Theorem 4.7 in (kneis2009bound) proved a bound for treewdith linear in the edge size.
-We do not know if this bound can be further improved.
-We will discuss this as an open problem at the end of this paper (Section 5).
-Before moving on to applications of our algorithm in the next section, we summarize the steps of exactly computing a $U$-statistic using our proposed algorithm.
-Given a $U$-statistic kernel $h$ and a sample $\bm{X}$:
+[…]
+Before moving on to applications of our algorithm in the next section, we summarize our proposed algorithm below. Given a $U$-statistic kernel $h$ and a sample $\bm{X}$:
   1. (1)
@@
     add all evaluated $V$-statistics together sequentially as in Algorithm 4.
+Remark 7.
+[…]
 4 Examples and Numerical Results
-In this section, we evaluate the performance of our new algorithm for computing UU/$V$-statistics exactly in three concrete statistical applications, including computing higher-order influence functions (Section 4.1), computing motif counts in random graphs (Section 4.2) and computing distance covariance measures (Section 4.3).
-As mentioned, we have made our algorithm available for public use as a Python package u-stats, which is available to download from either its GitHub repository (clicking u-stats) or the Python package index PyPi by typing the following on the command line:
-pip install u-stats
-We also refer interested readers to the GitHub link for how to implement u-stats in practice.
+In this section, we evaluate the performance of our new algorithm for computing UU/$V$-statistics in three concrete statistical applications [^2] The replication code can be found in this link..
+We have made our algorithm available for public use, which is available to download from the GitHub repository (by clicking u-stats or its R interface) or the Python package index PyPi.
+For the first application (Section 4.1) on Higher-Order Influence Functions (HOIFs) (robins2008higher), we also provide a standalone implementation based on u-stats, given the growing interest in using HOIFs in practice in the causal inference community (wager2024causal; zheng2025perturbed; cinelli2025challenges).
 4.1 Application 1: Higher-Order Influence Functions
-As mentioned in the Introduction, Higher-Order Influence Functions (HOIFs) (robins2008higher; robins2016technical) can help construct rate-optimal estimators for important parameters such as the treatment-specific mean under ignorability under certain complexity-reducing assumptions on the data generating mechanism.
+As mentioned in the Introduction, HOIFs (robins2008higher) can help construct rate-optimal estimators for important parameters such as the treatment-specific mean under ignorability under certain complexity-reducing assumptions on the data generating mechanism.
 However, estimators constructed via HOIFs are high-order $U$-statistics, which are time-consuming to compute in practice.
@@
 where for every $j=2,\cdots,m$, given $\bar{i}_{j}\in[n]^{j}$,
-$h^{(j)}(\bm{X}[\bar{i}_{j}])=A_{i_{1}}\phi(Z_{i_{1}})^{\top}\left\{\prod_{s=3}^{j}\left(A_{i_{s}}\phi(Z_{i_{s}})A_{i_{s}}\phi(Z_{i_{s}})^{\top}-\mathbb{I}\right)\right\}\phi(Z_{i_{2}})Y_{i_{2}}$
+$h^{(j)}(\bm{X}[\bar{i}_{j}])=A_{i_{1}}\phi(Z_{i_{1}})^{\top}\left\{\prod_{s=3}^{j}\left(A_{i_{s}}\phi(Z_{i_{s}})\phi(Z_{i_{s}})^{\top}-\mathbb{I}\right)\right\}\phi(Z_{i_{2}})Y_{i_{2}}$
 and we assume for simplicity that the population mean of $A\phi(Z)\phi(Z)^{\top}$ is $\mathbb{I}$.
@@
 $h_{j}^{(m)}(\bm{X}[\bar{i}_{j}])\coloneqq A_{i_{1}}\phi(Z_{i_{1}})^{\top}\left\{\prod_{s=3}^{j}A_{i_{s}}\phi(Z_{i_{s}})A_{i_{s}}\phi(Z_{i_{s}})^{\top}\right\}\phi(Z_{i_{2}})Y_{i_{2}}.$ (5)
-Then obviously $h_{j}^{(m)}$ permits a MD structure by letting $f_{l}(X_{1},X_{2})\coloneqq A_{1}\phi(Z_{1})^{\top}A_{2}\phi(Z_{2})$ for $l<j$ and $f_{j}(X_{1},X_{2})\coloneqq A_{1}\phi(Z_{1})^{\top}\phi(Z_{2})Y_{2}$.
-The corresponding decomposition signature of $h_{j}^{(m)}$ is $((i,i+1))_{i=1}^{j-1}$, which is denoted by ${\mathcal{A}}_{\mathsf{HOIF},j}$, so the decomposition graph of ${\mathcal{A}}_{\mathsf{HOIF},j}$ is a chain graph with e=j1e=j-1 edges.
-We also demonstrate the effectiveness of the sparsification trick using the HOIF example (see Lemma 2).
-The following table enlists the total number of $V$-statistics and the number of floating-point operations (FLOPs) that need to be computed before and after applying the sparsification trick when the sample size is $10^{4}$.
-The rightmost column of Table 1 also records the value of $M$ defined in Corollary 2.
-The derivation of this column can be found in Appendix C.2 (and unfortunately we do not know a more systematic way of deriving these).
-Table 1: Exact number of floating-point operations (FLOPs), $V$-statistic decompositions, and intermediate tensor sizes across HOIF orders from $2$ to $12$, with and without the sparsification trick, where the Einsum ordering is found by applying opt-einsum with optimization scheme greedy.
-Sample size is fixed at $n=10000$.
-Order ($m$)
-V-stat terms
-(Bell number)
-V-stat terms
-FLOPs
-FLOPs
-$M$
-2 2 1 $2.00020\times 10^{8}$ $2.00000\times 10^{8}$ 11
-3 5 2 $060\times 10^{12} 2.00020$$020\times 10^{12}$ 11
-4 15 5 $8.00310\times 10^{12}$ $4.00150\times 10^{12}$ 22
-5 52 15 $5.20133\times 10^{13}$ $60045\times 10^{13} 22$
-6 203 52 $2.94069\times 10^{14}$ $1.20020\times 10^{14}$ 22
-7 877 203 $1.78236\times 10^{15}$ $6.32091\times 10^{14}$ 22
-8 4140 877 $31054\times 10^{17} 2.2$$23430\times 10^{17}$ 33
-9 21147 4140 $5.29126\times 10^{18}$ $47959\times 10^{18} 33$
-10 115975 21147 $2.07990\times 10^{21}$ $2.04048\times 10^{21}$ 33
-11 678570 115975 $5.50026\times 10^{22}$ $3.72144\times 10^{22}$ 44
-12 4213597 678570 $1.13401\times 10^{24}$ $6.62181\times 10^{23}$ 44
-In Tables 2 and 3, we report the average runtime over 3 runs when computing the $U$-statistic $\mathbb{U}(h_{j}^{(m)},\bm{X})$ where kernel function $h_{j}^{(m)}$ is as in (5) with $m=7$ and $j\in\{2,3,\cdots,7\}$ and sample size $n\in\{1000,2000,4000,8000,10000\}$, respectively using CPU parallel computation and using a single GPU.
+[…]
+In Table 1, we report the average runtime over 10 runs when computing the $U$-statistic $\mathbb{U}(h_{j}^{(m)},\bm{X})$ where kernel function $h_{j}^{(m)}$ is as in (5) with $m=7$ and $j\in\{2,3,\cdots,7\}$ and sample size $n\in\{1000,2000,4000,8000,10000\}$, respectively using CPU parallel computation and using a single GPU.
 We would like to highlight that even when $m=7$, $n=10000$, it only takes about 5 minutes using our new package u-stats, with CPU parallel computation.
-If one instead has access to a single GPU, the runtime even reduces to about 12 seconds, a substantial speedup!
-We remark in passing that all the runtime reported in Table 2 and Table 3 excludes the time of kernel evaluation and tensorization.
-Table 2: Average runtime (in seconds) using CPU parallel computation.
-Experiments were conducted on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB, evaluated across varying sample sizes and orders of HOIF-type $U$-statistics.
+If one instead has access to a single GPU, the runtime even reduces to about 8 seconds, a substantial speedup!
+We remark in passing that all the runtime reported in Table 1 excludes the time of kernel evaluation and tensorization.
+Table 1: Average runtime (in seconds) using CPU (Intel Xeon ICX Platinum 8358 CPUs [2.6GHz, 64 total cores] with memory of 512 GB) with parallel computation (upper cell) and using a single GPU (NVIDIA RTX 5090, 32GB) with parallel computation (lower cell), evaluated across varying sample sizes and orders of HOIFs.
 Order 1000 2000 4000 8000 10000
-2 0.64 0.00141 0.01298 0.03174 0.04746
-3 0.00396 0.01518 0.07392 0.26849 0.45976
-4 0.02321 0.07313 0.32765 2.09545 2.36766
-5 0.09853 0.36239 1.71419 9.11349 14.21350
-6 0.38878 1.47719 7.44444 40.22143 58.85063
-7 1.91677 6.75947 34.08805 195.31414 290.54295
-Table 3: Average runtime (in seconds) using a single GPU (NVIDIA RTX 4090, 24GB) with parallel computation, across varying sample sizes and orders of HOIF-type $U$-statistics.
-Order 1000 2000 4000 8000 10000
-2 0.00184 0.00151 0.00155 0.00238 0.00339
-3 0.00172 0.00147 0.00193 0.00577 0.00745
-4 0.00305 0.00353 0.00707 0.03612 0.06245
-5 0.00835 0.01089 0.03456 0.19807 0.36160
-6 0.02608 0.03906 0.16981 1.12072 2.13328
-7 0.12031 0.19132 0.91884 6.45225 12.21350
-4.2 Application 2: Motif Counts in Random Graph Models and Networks
+2 0.00217 0.00070 0.00515 0.00142 0.01965 0.00492 0.03466 0.02779 0.04811 0.04398
+3 0.00388 0.00124 0.01548 0.00274 0.06968 0.01337 0.27130 0.06413 0.43110 0.08977
+4 0.01963 0.00242 0.07295 0.00475 0.32076 0.02375 1.81642 0.10690 2.33967 0.17264
+5 0.09356 0.00567 0.36376 0.00922 1.91751 0.04619 9.80295 0.23752 16.92521 0.41339
+6 0.45275 0.01863 1.52683 0.02545 7.76946 0.13763 41.78568 0.89001 62.54028 1.66450
+7 1.73287 0.07340 6.80758 0.10322 36.41280 0.64650 197.05856 4.49332 288.89672 8.58869
+4.2 Application 2: Motif Counts in Networks
 Motif (or subgraph) counts refer to the number of occurrences of a particular subgraph type $\mathsf{r}$ within a larger graph $G$, denoted by $\mathsf{C}(\mathsf{r},G)$.
@@
 Leveraging this observation, we show that various motif counts can be computed using a single Einsum operation.
-While this approach may not capture fine-grained structural information such as sparsity of the adjacency matrix, it is highly efficient in the context of dense graphs, significantly outperforming a popular software called igraph designed for computing motif counts.
+While this approach may not capture fine-grained structural information such as sparsity of the adjacency matrix, it is highly efficient in the context of dense graphs.
 Motif counts in a simple graph $G$ with $n$ vertices and the adjacency matrix $A\in\{0,1\}^{n\times n}$ can be represented as $U$-statistics of the following form:
 $\mathsf{C}(\mathsf{r},G)=\frac{1}{|\mathrm{Aut}(\mathsf{r})|}\sum_{\bar{i}_{m}\in\mathsf{Perm}(n,m)}h_{\mathsf{r}}(\bar{i}_{m};A),$
-where the kernel function $h_{\mathsf{r}}$ encodes the motif structure, and $|\mathrm{Aut}(\mathsf{r})|$ denotes the number of automorphisms of motif $\mathsf{r}$, which counts the number of symmetric repetitions.
-To be concrete, letting $B\equiv 1-A$, the corresponding $U$-statistics for motif counts for specific $3$-vertex motifs and $4$-vertex motifs (see Figure 2 for which motifs $\mathsf{r}_{1}$ to $\mathsf{r}_{8}$ encode) read as follows:
-$\mathsf{C}(\mathsf{r}_{1},G)$ $=\frac{1}{2}\sum_{\bar{i}_{3}\in\mathsf{Perm}(n,3)}A_{i_{1},i_{2}}A_{i_{2},i_{3}}B_{i_{3},i_{1}},$
-$\mathsf{C}(\mathsf{r}_{2},G)$ $=\frac{1}{6}\sum_{\bar{i}_{3}\in\mathsf{Perm}(n,3)}A_{i_{1},i_{2}}A_{i_{2},i_{3}}A_{i_{3},i_{1}},$
-$\mathsf{C}(\mathsf{r}_{3},G)$ $=\frac{1}{6}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}B_{i_{2},i_{3}}B_{i_{2},i_{4}}B_{i_{3},i_{4}},$
-$\mathsf{C}(\mathsf{r}_{4},G)$ $=\frac{1}{2}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}B_{i_{1},i_{3}}B_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
-$\mathsf{C}(\mathsf{r}_{5},G)$ $=\frac{1}{2}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}B_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
-$\mathsf{C}(\mathsf{r}_{6},G)$ $=\frac{1}{8}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}B_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
-$\mathsf{C}(\mathsf{r}_{7},G)$ $=\frac{1}{4}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}A_{i_{2},i_{4}}B_{i_{3},i_{4}},$
-$\mathsf{C}(\mathsf{r}_{8},G)$ $=\frac{1}{24}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}A_{i_{2},i_{4}}A_{i_{3},i_{4}}.$
-We compare our package u-stats with igraph (csardi2006igraph) on exactly computing all 3-vertex and all 4-vertex motif counts.
-igraph computes motif counts by using a highly optimized algorithm based on wernicke2006fanmod.
-We generate random graphs using the classical Erdös-Rényi model $G(n,p)$, where each of the $\binom{n}{2}$ possible edges between $n$ vertices is present independently with probability $p$.
-We report the runtime comparisons in Table 4 and Table 5.
-All experiments were conducted on a server equipped with Intel Xeon Platinum 8358 CPUs (2.6GHz, 64 cores) and memory of 512 GB.
-Our implementation makes full use of all 64 CPU cores via multiprocessing, while igraph runs in a single-threaded setting, due to its current implementation limitation.
-To quantify the performance, we consider two metrics:
-  • Speedup: the ratio of igraph runtime to the runtime of u-stats with parallel computation;
-  • Single-core speedup: a very rough estimate of per-core speedup, computed as the speedup divided by the number of cores (i.e. $64$).
-As shown in Table 4 and Table 5, igraph has better runtime performance than u-stats when $p$ is small, corresponding to a sparse graph.
-Because u-stats is designed as a general purpose package for computing $U$-statistics, we do not further take into account the potential sparsity of the adjacency matrix.
-When $p$ increases (i.e., as the graph becomes denser), the runtime of igraph grows considerably even when $p$ is as low as $0.1$.
-Yet u-stats still maintains roughly the same runtime throughout, which is much faster than that of igraph when $p$ grows.
-(a) $\mathsf{r}_{1}$: V-shape
-(b) $\mathsf{r}_{2}$: Triangle
-(c) $\mathsf{r}_{3}$ : 3-star
-(d) $\mathsf{r}_{4}$ : Fork
-(e) $\mathsf{r}_{5}$ : Tailed triangle
-(f) $\mathsf{r}_{6}$: Square
-(g) $\mathsf{r}_{7}$: House
-(h) $\mathsf{r}_{8}$ : 4-clique
-Figure 2: All non-degenerate isomorphism classes of 3-vertex and 4-vertex simple undirected graphs.
-Table 4: Runtime comparison of exact all 33-vertex motif counts using u-stats and igraph on Erdös-Rényi graphs $G(n,p)$ with $n=5623$.
-Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB.
-“Speedup ratio” compares igraph to our parallel method; “single core speedup” assumes single-core execution.
-Edge
-Prob. $p$
-u-stats
-Time (s)
-igraph
-Time (s)
-Speedup
-Ratio
-Single-Core
-Speedup
-0.001 0.617 0.025 0.040 0.001
-0.005 0.649 0.353 0.544 0.008
-0.010 0.692 1.795 2.595 0.041
-0.020 0.783 10.594 13.529 0.211
-0.050 1.038 136.272 131.288 2.051
-0.080 1.298 514.702 396.415 6.194
-0.100 1.453 960.299 660.690 10.323
-0.150 1.864 3003.490 1611.389 25.178
-0.200 2.275 6725.515 2955.656 46.182
-Table 5: Runtime comparison of exact all 44-vertex motif counts using u-stats and igraph on Erdös-Rényi graphs $G(n,p)$ with $n=2000$.
-Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB.
-“Speedup ratio” compares igraph to our parallel method; “single core speedup” assumes single-core execution.
-Edge
-Prob. $p$
-u-stats
-Time (s)
-igraph
-Time (s)
-Speedup
-Ratio
-Single-Core
-Speedup
-0.001 70.959 0.004 0. 0.
-0.005 71.947 0.168 0.002 0.
-0.010 68.463 1.453 0.021 0.
-0.020 65.634 15.573 0.237 0.004
-0.050 65.786 394.532 5.997 0.094
-0.080 66.866 2371.373 35.465 0.554
-0.100 68.219 5396.057 79.099 1.236
-0.150 65.969 23003.041 348.692 5.448
-0.200 65.259 61796.760 946.940 14.796
-4.3 Application 3: Distance Covariance Measures for Independence Testing
+[…]
+We compare our package u-stats with several state-of-the-art systems for motif counting, including the widely used graph analysis library igraph (csardi2006igraph), the parallel motif counting system Peregrine (jamshidi2020peregrine), and NVIDIA’s GPU-based library cuGraph (fender2022rapids).
+These systems represent different computational paradigms.
+igraph relies on highly optimized enumeration algorithms derived from wernicke2006fanmod and runs in a single-threaded setting.
+Peregrine implements parallel, motif-specific optimizations for fast CPU-based counting.
+cuGraph provides specialized GPU kernels, but currently supports only a limited set of motifs (e.g., triangles).
+In contrast, u-stats adopts a unified formulation based on $U$-statistics and Einsum, without requiring implementations specific to motif counting.
+[…]
+The results reveal distinct performance regimes across methods.
+On sparse graphs (small $p$), Peregrine and igraph outperform u-stats due to their ability to exploit sparsity through enumeration-based strategies.
+However, as the graph becomes denser, their runtime increases significantly, with igraph failing to complete within the time limit even at moderate densities.
+In contrast, u-stats maintains nearly constant runtime across all values of $p$, resulting in substantial performance advantages in dense regimes.
+The results for 4-vertex motifs in Table 7 exhibit similar trends.
+On GPUs, cuGraph achieves faster runtimes on sparse graphs due to specialized kernels, but its performance degrades with increasing density and eventually runs out of memory, whereas u-stats remains stable without requiring motif-specific implementations.
+Table 2: Runtime comparison of exact all $3$-vertex motif counts using u-stats, Peregrine and igraph on Erdős–Rényi graphs $G(n,p)$ with $n=20000$.
+Experiments were run on Intel Xeon Scalable Cascade Lake 6248 CPUs (2.5GHz, 40 total cores) with memory of 192 GB.
+“OOT” denotes instances that exceeded the time limit of 1800 seconds.
+Edge Prob. $p$ u-stats Time (s) Peregrine Time (s) igraph Time (s)
+0.0005 10.3478 0.0703 0.1915
+0.0010 10.4452 0.0740 0.7259
+0.0050 10.6196 0.0780 40.7373
+0.0100 10.3616 0.1190 289.5082
+0.0500 9.8357 1.1782 OOT
+0.2000 10.6010 17.1063 OOT
+0.4000 10.8445 67.6727 OOT
+0.8000 8.8623 158.7377 OOT
+4.3 Application 3: Distance Covariance Measures
 In the last example, we revisit the problem considered in Section 5.1 of shao2025u, which concerns the problem of computing the squared distance covariance $\mathsf{dCov}^{2}$, a recently proposed dependence measure between two random vectors (szekely2007measuring; yao2018testing).
 In particular, let $O_{i}\coloneqq(X_{i},Y_{i}),i\in[n]$ be the observed sample.
-$\mathsf{dCov}^{2}(X,Y)$ can then be written as a $4$-th order $U$-statistic
+$\mathsf{dCov}^{2}(X,Y)$ can be written as a $4$-th order $U$-statistic: Denoting $\|\cdot\|_{2}$ as the vector $\ell_{2}$-norm, then
 $\mathsf{dCov}^{2}(X,Y)=\frac{(n-4)!}{n!}\sum_{\bar{i}_{4}\in\mathsf{Perm}([n],4)}$ $\Bigg(h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{3}},O_{i_{4}})+h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{1}},O_{i_{2}})$
 $-h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{1}},O_{i_{3}})-h_{1}(O_{i_{1}},O_{i_{2}})h_{2}(O_{i_{2}},O_{i_{4}})\Bigg),$
-where $h_{1}(O_{i},O_{j})\coloneqq\|X_{i}-X_{j}\|_{2}$ and $h_{2}(O_{i},O_{j})\coloneqq\|Y_{i}-Y_{j}\|_{2}$ for $i,j\in[n]$. Here $\|\cdot\|_{2}$ denotes the vector $\ell_{2}$-norm.
-Quoting from Table 5 of shao2025u, it takes “8099.73 seconds” to compute $\mathsf{dCov}^{2}$ with $n=138$ among 11 random vectors.
+where $h_{1}(O_{i},O_{j})\coloneqq\|X_{i}-X_{j}\|_{2}$ and $h_{2}(O_{i},O_{j})\coloneqq\|Y_{i}-Y_{j}\|_{2}$ for $i,j\in[n]$.
+Quoting from Table 5 of shao2025u, it takes “8099.73 seconds” to compute $\mathsf{dCov}^{2}$ among 11 random vectors, together with its confidence interval, when $n=138$.
 Each vector corresponds to a different stock sector—such as Health Care or Industrials—and its components represent the monthly log-returns of individual stocks within that sector, using historical S&P 500 data.
-To bypass such a high computational burden, shao2025u resort to computing a randomized incomplete $U$-statistic instead.
+shao2025u resort to computing a randomized incomplete $U$-statistic instead.
 In the same Table 5, shao2025u demonstrated that there is indeed a significant speedup when using randomized incomplete $U$-statistics.
-It is noteworthy that another motivation of shao2025u to use randomized incomplete $U$-statistic is due to its asymptotic normality (chen2019randomized) even when the original complete $U$-statistic has a Gaussian chaos limiting distribution.
-But we will not further discuss this issue as it is not directly relevant to the main theme of our paper.
 In Table 3, we display the average runtime over 10 runs of exactly computing $\mathsf{dCov}^{2}$ using our package u-stats, or computing $\mathsf{dCov}^{2}$ exactly or approximately with the randomized incomplete $U$-statistic using the MATLAB code attached in the supplementary materials of shao2025u.
-We rerun the code of shao2025u together with u-stats on the same computing platform to ensure a fair comparison.
+We rerun the code of shao2025u together with u-stats on the same computing platform to ensure a fair comparison and we only focus on point estimation.
 In particular, the algorithm proposed in shao2025u will randomly choose $O(n^{\alpha})$ permutations, with $\alpha>0$ a tuning parameter determined by the available computation budget.
@@
 We vary the tuning parameter $\alpha$ within the set $\{1.5,2,2.5\}$.
-Our package u-stats again exhibits quite impressive runtime performance: u-stats without parallel computation takes slightly shorter time than the randomized incomplete $U$-statistic when $\alpha=2$, whereas u-stats with parallel computation takes less than half of the time as the case when $\alpha=1.5$!
+Our package u-stats again exhibits quite impressive runtime performance: u-stats without parallel computation takes slightly shorter time than the randomized incomplete $U$-statistic when $\alpha=2$, whereas u-stats with parallel computation takes less than half of the time of the case when $\alpha=1.5$!
 Table 3: Runtime (in seconds) comparison of various methods for computing $\mathsf{dCov}^{2}$. Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB.
 u-stats shao2025u’s MATLAB code
-No Parallel
-Parallel
-Randomized
-$\alpha=1.5$
-Randomized
-$\alpha=2.0$
-Randomized
-$\alpha=2.5$
-Complete
+No Parallel Parallel Randomized $\alpha=1.5$ Randomized $\alpha=2.0$ Randomized $\alpha=2.5$ Complete
 4.0928 0.1847 0.4211 4.5744 53.0265 3395.4001
 5 Concluding Remarks
-In this paper, we have conducted a delicate and systematic study on computing and the complexity of computing higher-order $U$-statistics (and also $V$-statistics), exactly.
-Explicit connections between the complexity of computing higher-order UU/$V$-statistics and concepts from tensor computation and graph theory are established.
-An accompanying Python package u-stats is provided to help practitioners compute UU/$V$-statistics more efficiently in applications.
-Although Python becomes more and more popular among statisticians (dai2023rehline), an R interface of u-stats will also be available in due course.
-It will also be interesting to test u-stats on other problems involving $U$-statistics (qiu2024neural; shen2025engression).
+In this paper, we have conducted a delicate and systematic study on exactly computing and the complexity of exactly computing higher-order $U$-statistics (and also $V$-statistics).
+Explicit connections are established between the complexity of computing higher-order UU/$V$-statistics and concepts from tensor computation and graph theory.
+An accompanying Python package u-stats and its R interface are provided to help practitioners compute UU/$V$-statistics more efficiently in applications.
+It will also be interesting to test u-stats on other problems involving $U$-statistics (shen2025engression).
 One limitation of our work is that we prioritize time complexity over space complexity.
-As a result, when we need to store a very high-order tensor, the sample size $n$ cannot be too large – this is why we have only evaluated the performance of u-stats on computing HOIF estimators with order $m\leq 7$ when $n=10000$.
-An important future direction is thus to explore how to balance the time-complexity and space-complexity when the sample size $n$ is large.
-Finally, we conclude our paper with an open problem that we ourselves find intriguing.
-As we indicated in Remark 7 at the end of Section 3, the treewidth of the decomposition graph associated with the Einsum notation of the underlying $V$-statistic relates to its time complexity.
-But the treewidth of a graph is not easy to compute either analytically (harvey2018treewidth) or numerically (arnborg1987complexity).
-It will be interesting to get a sharp bound on treewidth using the number of vertices or edges of the graph.
-As mentioned in Remark 7, kneis2009bound has proved that the treewidth is bounded at most linearly by the edge size, or more precisely bounded by $e/5.769+O(\log n)$ where $n\coloneqq|V|$ denotes the number of vertices in $G$.
-Yet we do not know if this bound can be further improved.
-Below we provide the maximum tree-width possible for edge size no greater than $15$, to hopefully serve as a starting point of fully addressing this problem.
-Observation 1 (Treewidth of Graphs with a Given Number of Edges).
-Let
-$\mathrm{t}(e)\coloneqq\max\left\{\mathsf{tw}(G)\;\middle|\;G\text{ is a graph where }|E(G)|=e\right\}.$
-We have computed the exact values of $\mathrm{t}(e)$ for all $1\leq e\leq 15$ as follows:
+As a result, when we need to store a very high-order tensor, the sample size $n$ cannot be too large; this is why we have only evaluated the performance of u-stats on computing HOIF estimators of order $m\leq 7$ when $n=10000$.
+Therefore, an important future direction is to explore how to balance time and space complexities when both $m$ and $n$ are even larger.
+SUPPLEMENTARY MATERIAL
 […]
-The proof can be found in Appendix C.1. However, extending the above result to $e>15$ is currently beyond our reach.
-Appendix A Proof of Proposition 1
+In this section, we provide two additional examples to illustrate the definitions of the Einsum notation and the Einsum operation, supplementing Example 1.
+Example 2.
+We take matrix multiplication as an example to illustrate the Einsum operation just defined.
+Let $T_{1},T_{2}$ be two input matrices, which are also 2nd-order tensors.
+Therefore $\mathcal{T}=(T_{1},T_{2})$.
+Suppose we are interested in multiplying $T_{1}$ and $T_{2}$, denoted as $T^{\rm out}$: for any $i,j\in[n]$, $T^{\rm out}(i,j)=\sum\limits_{k=1}^{n}T_{1}(i,k)T_{2}(k,j)$.
+The matrix multiplication corresponds to the Einsum operation with Einsum notation $\mathcal{N}=(\mathcal{A};B)$, with input tuple list $\mathcal{A}=((1,2),(2,3))$ and output tuple $B=(1,3)$ and $|B|=2$.
+The index set is simply $I=[3]$.
+This means that the operation sums over the dimensions of $T_{1}$ and $T_{2}$ corresponding to index $2$ and reserves indices 11 and $3$.
+$T^{\rm out}(\alpha)$ for any $\alpha=(i,j)\in[n]^{2}$ can be written in the Einsum operation form $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ because by definition:
+$\textsc{Einsum}(\mathcal{T},\mathcal{N})$ $\equiv\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}\prod_{k=1}^{2}T_{k}(\alpha^{\prime}[A_{k}])=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])$
+$=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[1,2])T_{2}(\alpha^{\prime}[2,3])=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[3])$
+$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha[3])=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(i,\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],j),$
+so the claim holds by identifying $\alpha^{\prime}[2]=k$.
+Example 3.
+Here we consider an example involving higher-order tensors.
+Given three tensors $\mathcal{T}=(T_{1},T_{2},T_{3})$, each of size $n$, where $T_{1}$ and $T_{2}$ are 3rd-order tensors while $T_{3}$ is a matrix, suppose that we are interested in computing the following 3rd-order tensor: for any $(i,j,k)\in[n]^{3}$
+$T^{\rm out}(i,j,k)=\sum_{a=1}^{n}\sum_{b=1}^{n}\sum_{c=1}^{n}T_{1}(i,a,b)T_{2}(a,c,j)T_{3}(b,k).$
+$T^{\rm out}(i,j,k)$ is in fact a partial trace, a common operation taken in quantum mechanics.
+The corresponding Einsum notation is $\mathcal{N}=(\mathcal{A};B)$ with $\mathcal{A}=(A_{1}=(1,2,3),A_{2}=(2,4,5),A_{3}=(3,6))$ with $B=(1,5,6)$ and the index set $I=[6]$.
+$T^{\rm out}(\alpha)$ for any $\alpha=(i,j,k)\in[n]^{3}$ can be written in the Einsum operation form $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ because by definition:
+$\textsc{Einsum}(\mathcal{T},\mathcal{N})$ $\equiv\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{4}\\ \alpha^{\prime}[1,5,6]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])T_{3}(\alpha^{\prime}[A_{3}])$
+$=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{4}\\ \alpha^{\prime}[1,5,6]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
+$=\sum_{\alpha^{\prime}[2]=1}^{n}\sum_{\alpha^{\prime}[3]=1}^{n}\sum_{\alpha^{\prime}[4]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2],\alpha^{\prime}[3])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[4],\alpha[5])T_{3}(\alpha^{\prime}[3],\alpha[6]).$
+Thus the claim holds by identifying $\alpha^{\prime}[2]=a$, $\alpha^{\prime}[3]=b$, and $\alpha^{\prime}[4]=c$.
+[…]
+In he2021asymptotically; lai2023block, higher-order $U$-statistics are employed to test diagonal and block-diagonal structures of high-dimensional covariance matrices.
+A detailed computational analysis, including an iterative algorithm and complexity bounds, is provided in he2021asymptotically.
+Consider the sample $\bm{X}=(X_{i})_{i=1}^{n}$ with $X_{i}\in\mathbb{R}^{p}$. The $U$-statistic used to construct the test (omitting the normalizing constant) takes the form
+$\mathcal{U}(a)=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{\bar{i}_{2a}\in\mathsf{Perm}([n],2a)}\prod_{k=1}^{a}(X_{i_{2k-1},j_{1}}X_{i_{2k-1},j_{2}}-X_{i_{2k-1},j_{1}}X_{i_{2k},j_{2}}),$ (6)
+where $\mathbf{I}\subseteq[p]^{2}$ depends on the specific testing goal. For diagonal testing, he2021asymptotically uses $\mathbf{I}=[p]^{2}\setminus\bigcup_{j=1}^{p}\{(j,j)\}$.
+When the mean of each $X_{i}$ is known to be zero, the statistic simplifies to
+$\widetilde{\mathcal{U}}(a)=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{\bar{i}_{a}\in\mathsf{Perm}([n],a)}\prod_{k=1}^{a}(X_{i_{k},j_{1}}X_{i_{k},j_{2}}).$
+For this simplified form, he2021asymptotically presents an iterative algorithm (which is a special case of our algorithm) and establishes a computational complexity of $O(p^{2}n)$ independent of the order $a$.
+In terms of the original test statistic $\mathcal{U}(a)$ in (6), however, he2021asymptotically only provides an algorithm and complexity analysis for $a\leq 2$. They remark that:
+When $a\geq 3$, a similar iterative method can be applied, but a closed-form expression may be difficult to derive directly.
+The difficulty in obtaining a closed-form expression stems precisely from the decomposition challenge addressed in Lemma 1 of our paper, which expresses a general $U$-statistic as a linear combination of $V$-statistics.
+To circumvent this issue, he2021asymptotically instead considers a modified test statistic $\mathcal{U}_{c}(a)$ resembling $\widetilde{\mathcal{U}}(a)$ and shows that it exhibits asymptotic properties similar to $\mathcal{U}(a)$.
+Our algorithm u-stats can directly compute $\mathcal{U}(a)$, and our complexity analysis developed in Section 3.3 yields the same runtime complexity $O(p^{2}n)$.
+To see this, we first expand the product using the binomial theorem:
+$\mathcal{U}(a)$ $=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{\bar{i}_{2a}\in\mathsf{Perm}([n],2a)}\sum_{S\subseteq[a]}(-1)^{|S|}\prod_{k\notin S}(X_{i_{2k-1},j_{1}}X_{i_{2k-1},j_{2}})\prod_{k\in S}(X_{i_{2k-1},j_{1}}X_{i_{2k},j_{2}})$
+$=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{S\subseteq[a]}(-1)^{|S|}\,\mathbb{U}(h_{S,j_{1},j_{2}},\bm{X}),$
+where
+$h_{S,j_{1},j_{2}}(\bar{X}_{2a})=\prod_{k\notin S}(X_{i_{2k-1},j_{1}}X_{i_{2k-1},j_{2}})\prod_{k\in S}(X_{i_{2k-1},j_{1}})\prod_{k\in S}(X_{i_{2k},j_{2}}).$
+Thus, $h_{S,j_{1},j_{2}}$ admits a MD structure (Definition 9) with
+$\mathcal{H}_{S,j_{1},j_{2}}$ $=\bigl(\underbrace{h_{1}(X_{i})=X_{i,j_{1}}X_{i,j_{2}}}_{\text{repeated }a-|S|\text{ times}},\ \underbrace{h_{2}(X_{i})=X_{i,j_{1}}}_{\text{repeated }|S|\text{ times}},\ \underbrace{h_{3}(X_{i})=X_{i,j_{2}}}_{\text{repeated }|S|\text{ times}}\bigr),$
+$\mathcal{A}_{S}$ $=\bigl((2k-1)_{i\notin S},(2k-1)_{i\in S},(2k)_{i\in S}\bigr).$
+The corresponding decomposition graph $G_{\mathcal{A}_{S}}$ (Figure 2) consists solely of isolated vertices (i.e., it is edgeless).
+Consequently, all quotient graphs are also edgeless and have treewidth 0.
+By Corollary 2, each $\mathbb{U}(h_{S,j_{1},j_{2}},\bm{X})$ can be computed in $O(n)$ time for fixed $S$ and $(j_{1},j_{2})$.
+Since $|\mathbf{I}|=O(p^{2})$ and there are $2^{a}$ possible subsets $S\subseteq[a]$, the total complexity of computing $\mathcal{U}(a)$ is $O(2^{a}p^{2}n)$.
+In the regime where $a$ is small relative to $n$ and $p$ (the setting considered in he2021asymptotically), this reduces to $O(p^{2}n)$.
+[…]
+Figure 2: The decomposition graph $G_{\mathcal{A}_{S}}$, which consists of $a+|S|$ isolated vertices.
+[…]
 Our proposed algorithm for computing a tensor $\mathcal{T}$ is based on iterating the Einsum operation w.r.t.
@@
 This is why the treewidth of the decomposition graph can be used to give an optimistic estimate of the time complexity.
-Definition 13 (Einsum Notation Sequence under a Particular Summation Ordering).
-Let ${\mathcal{A}}=(A_{1},\cdots,A_{K})$ be an Einsum notation (recall from Definition 4) with no output of.
+Definition 13 (Einsum Notation Sequence).
+Let $\mathcal{A}=(A_{1},\cdots,A_{K})$ be an Einsum notation (recall from Definition 4) with no output.
 Let $I=[m]$ be the index set of $\mathcal{A}$ and $\sigma$ be a permutation over $I$.
-The Einsum notation sequence of $\mathcal{N}$} under $\sigma$ is the sequence ${\mathcal{A}}^{\sigma}_{0},{\mathcal{A}}^{\sigma}_{1},\ldots,{\mathcal{A}}^{\sigma}_{m}$ defined recursively as follows:
+The Einsum notation sequence of $\mathcal{A}$ under $\sigma$ is the sequence $\mathcal{A}^{\sigma}_{0},\mathcal{A}^{\sigma}_{1},\ldots,\mathcal{A}^{\sigma}_{m}$ defined recursively as follows:
   • $\mathcal{A}^{\sigma}_{0}=\mathcal{A}$.
@@
 The following lemma establishes a one-to-one correspondence between the Einsum notation sequence under a particular summation ordering $\sigma$ defined in Definition 13 and vertex elimination given in Definition 6.
-Lemma 3 (The Einsum Notation Sequence, Vertex Elimination Correspondence).
+Lemma 3.
 Let $\mathcal{A}$ be an Einsum notation with index set $I=[m]$ and $\sigma$ be a permutation of $I$.
@@
 Proof.
-We immediately have |B|=d1|B|=d-1, which implies that the result of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ has $n^{d-1}$ entries.
+We immediately have |B|=d1|B|=d-1, which implies that the result of $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ has $n^{d-1}$ entries.
 For each entry, as one index is summed over, it takes n1n-1 summation operations and $(|\mathcal{A}|-1)n$ multiplication operations.
-Therefore, the total time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ is $\big{(}(n-1)n^{d-1}+(|{\mathcal{A}}|-1)n^{d}\big{)}$, which is $\Theta(|{\mathcal{A}}|n^{d})$.
+Therefore, the total time complexity of computing $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ is $\big((n-1)n^{d-1}+(|\mathcal{A}|-1)n^{d}\big)$, which is $\Theta(|\mathcal{A}|n^{d})$.
 ∎
-[…]
+We need to introduce another notation before presenting the final proof: given two lists of tuples $\mathcal{A}$ and $\mathcal{B}$, we write $\mathcal{A}\sqsubseteq\mathcal{B}$ to mean that $\mathcal{A}$ is a sublist of $\mathcal{B}$.
 Proof of Proposition 1.
@@
 We define the $\mathcal{A}^{\sigma}_{0},\mathcal{A}^{\sigma}_{1},\ldots,\mathcal{A}^{\sigma}_{m}$ be the Einsum notation sequence under $\sigma$ as in Definition 13.
-We first show that there is an algorithm such that the time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(n^{C_{\sigma}+1})$ where
+We first show that there is an algorithm such that the time complexity of computing $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ is $O(n^{C_{\sigma}+1})$ where
 $C_{\sigma}\coloneqq\max_{i\in[m]}\deg(\sigma[i],G_{\mathcal{A}^{\sigma}_{i-1}}).$
@@
     and similarly let $\mathcal{T}^{\sigma,\mathrm{in}}_{i}\sqsubseteq\mathcal{T}^{\sigma}_{i-1}$ be the tensor tuple corresponding to $\mathcal{A}^{\sigma,\mathrm{in}}_{i}$;
-  • Let $B^{\sigma,\mathrm{out}}_{i}$ be a permutation of the set $\bigcup_{\begin{subarray}{c}A\in{\mathcal{A}}^{\sigma}_{i-1}\\ \sigma[i]\in A\end{subarray}}(\mathsf{set}(A)\setminus\{\sigma[i]\})$ as the output tuple and $T^{\sigma,\mathrm{out}}_{i}\coloneqq\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$;
+  • Let $B^{\sigma,\mathrm{out}}_{i}$ be a permutation of the set $\bigcup_{\begin{subarray}{c}A\in\mathcal{A}^{\sigma}_{i-1}\\ \sigma[i]\in A\end{subarray}}(\mathsf{set}(A)\setminus\{\sigma[i]\})$ as the output tuple and $T^{\sigma,\mathrm{out}}_{i}\coloneqq\textsc{Einsum}(\mathcal{T}^{\sigma,\mathrm{in}}_{i},(\mathcal{A}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$;
   • Update the tensor tuple $\mathcal{T}^{\sigma}_{i}$ from $\mathcal{T}^{\sigma}_{i-1}$ by
@@
 According to the distributive property of multiplication over summation, it holds that for any $i,j\in[m]$,
-$\mathsf{Einsum}({\mathcal{T}}^{\sigma}_{i},{\mathcal{A}}^{\sigma}_{i})\equiv\mathsf{Einsum}({\mathcal{T}}^{\sigma}_{j},{\mathcal{A}}^{\sigma}_{j})\equiv\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}).$
-Thus the time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ just comes from transferring from one step to the next step, which is $\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$.
-Following Lemma 4, we obtain that the time complexity of computing $\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$ is $\Theta(|{\mathcal{A}}^{\sigma,\mathrm{in}}_{i}|n^{\deg(x,G_{{\mathcal{A}}^{\sigma}_{i-1}})})$, which is also $\Theta(|{\mathcal{A}}|n^{\deg(\sigma[i],G_{{\mathcal{A}}^{\sigma}_{i-1}})})$ as $|{\mathcal{A}}^{\sigma,\mathrm{in}}_{i}|\leq|{\mathcal{A}}|$.
+$\textsc{Einsum}(\mathcal{T}^{\sigma}_{i},\mathcal{A}^{\sigma}_{i})\equiv\textsc{Einsum}(\mathcal{T}^{\sigma}_{j},\mathcal{A}^{\sigma}_{j})\equiv\textsc{Einsum}(\mathcal{T},\mathcal{A}).$
+Thus the time complexity of computing $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ just comes from transferring from one step to the next step, which is $\textsc{Einsum}(\mathcal{T}^{\sigma,\mathrm{in}}_{i},(\mathcal{A}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$.
+Following Lemma 4, we obtain that the time complexity of computing $\textsc{Einsum}(\mathcal{T}^{\sigma,\mathrm{in}}_{i},(\mathcal{A}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$ is $\Theta(|\mathcal{A}^{\sigma,\mathrm{in}}_{i}|n^{\deg(x,G_{\mathcal{A}^{\sigma}_{i-1}})})$, which is also $\Theta(|\mathcal{A}|n^{\deg(\sigma[i],G_{\mathcal{A}^{\sigma}_{i-1}})})$ as $|\mathcal{A}^{\sigma,\mathrm{in}}_{i}|\leq|\mathcal{A}|$.
 By taking maximum over $i$, the overall time complexity is thus $O(|\mathcal{A}|n^{C_{\sigma}+1})$.
@@
 $G_{\mathcal{A}^{\sigma}_{i}}=\mathsf{elim}_{\sigma[i]}(G_{\mathcal{A}^{\sigma}_{i-1}}),$
-and by finding the permutation $\sigma^{\ast}$ minimizing $C_{\sigma}$ and recalling Definition 7, we can conclude that there exists an algorithm such that the time complexity of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(|{\mathcal{A}}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$, which completes the proof.
+and by finding the permutation $\sigma^{\ast}$ minimizing $C_{\sigma}$ and recalling Definition 7, we can conclude that there exists an algorithm such that the time complexity of $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ is $O(|\mathcal{A}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$, which completes the proof.
 ∎
-Appendix B Proof of Lemma 1
+[…]
 Here we prove an extended version of Lemma 1, as shown in Lemma 5 below. Lemma 1 then becomes a simple corollary of Lemma 5. We begin by introducing the following definitions.
@@
 Let $m,n$ be positive integers satisfying $m\leq n$ and $\pi$ be a partition of the set $[m]$. A U-set of size $n$ w.r.t. the partition $\pi$ is defined as
-${\mathcal{U}}\left(n,\pi\right)\coloneqq\left\{(s_{1},s_{2},\ldots,s_{m})\in[n]^{m}\mid\text{where for any $i,j\in[m]$},i,j\in Q\text{ for some $Q\in\pi$}\iff s_{i}=s_{j}\right\}.$
+$\mathcal{U}\left(n,\pi\right)\coloneqq\left\{\overline{s}_{m}\in[n]^{m}\mid\text{where for any $i,j\in[m]$},i,j\in Q\text{ for some $Q\in\pi$}\iff s_{i}=s_{j}\right\}.$
 If $\pi$ happens to be the finest partition of $[m]$, i.e.
@@
 The cardinality of the given partition $\pi$ is referred to as the order of $\mathcal{U}\left(n,\pi\right)$.
-Definition 16 (Restricted $U$-statistics).
+Definition 16 (Restricted $U$-Statistic).
 Let $\mathbb{X}$ be a non-empty set, $m,n$ be positive integers satisfying $m\leq n$, $\pi$ be a partition of set $[m]$ and $h:\mathbb{X}^{m}\rightarrow{\mathbb{R}}$ be a function defined on $\mathbb{X}^{m}$.
@@
 In particular, if $\pi$ happens to be $\{\{k\}\mid k\in[m]\}$, the above result reduces to Lemma 1.
-The proof of Lemma 5 can be found in Appendix B.2.
-B.1 Preparatory Materials
+[…]
+[…]
 Before presenting the proof in the next section, we introduce some useful preparatory materials, including: (1) the decomposition of a $V$-statistic into $U$-statistics and (2) a dual form of Möbius inversion formula (stanley2011enumerative).
 We first state the following technical result of decomposing a $V$-statistic into a linear combination of $U$-statistics.
-Lemma 6 (Decomposition of a $V$-statistic into $U$-statistics).
+Lemma 6 (Decomposition of $V$-Statistic into $U$-Statistics).
 Let
@@
 Next, we introduce the notion of Möbius inversion formula, which is used to reverse the decomposition from a $V$-statistic into $U$-statistics in Lemma 6 to obtain the decomposition from a $U$-statistic into $V$-statistics.
-Definition 17 (Möbius Function, (stanley2011enumerative)).
+Definition 17 (Möbius Function, stanley2011enumerative).
 Let $(P,\preceq)$ be a finite poset and, a function $\mu:P^{2}\to{\mathbb{R}}$ is said to be a Möbius function on $(P,\preceq)$, if
@@
 $\mu(s,u)=$ $-\sum_{\begin{subarray}{c}t\in P\\ s\preceq t\prec u\end{subarray}}\mu(s,t),\quad\forall\ s,u\in P:s\prec u.$
-Lemma 7 (Möbius Function on Partition Lattices, (stanley2011enumerative)).
+Lemma 7 (Möbius Function on Partition Lattices, stanley2011enumerative).
 Let $S$ be a non-empty set. Then the Möbius function on $(\Pi_{S},\preceq)$ takes the following form
 $\mu(\pi,\rho)=(-1)^{|\pi|-|\rho|}\prod_{C\in\rho}(\kappa_{\pi,C}-1)!,\quad\forall\,\pi,\rho\in\Pi_{S}:\pi\preceq\rho\text{ and }\kappa_{\pi,C}=|\{A\in\pi\mid A\subseteq C\}|.$
-Lemma 8 (Möbius Inversion Formula, Dual Form, (stanley2011enumerative)).
+Lemma 8 (Möbius Inversion Formula, Dual Form, stanley2011enumerative).
 Let $(P,\preceq)$ be a finite poset and $f,g$ be functions defined on $P$. Then
@@
 where $\mu$ is the Möbius function on $(P,\preceq)$.
-B.2 Proof of Lemma 5
+[…]
 In this section, we prove Lemma 5. The argument is straightforward given the preparations in the previous section.
@@
 ∎
-Appendix C Proof of Observation 1 and Results in Section 4.1
-C.1 Proof of Observation 1
+[…]
+In this section, we present an open problem that we ourselves find intriguing.
+As we indicated in Remark 6 at the end of Section 3, the treewidth of the decomposition graph associated with the Einsum notation of the underlying $V$-statistic relates to its time complexity.
+But the treewidth of a graph is not easy to compute either analytically (harvey2018treewidth) or numerically (arnborg1987complexity).
+It will be interesting to get a sharp bound on treewidth using the number of vertices or edges of the graph.
+As mentioned in Remark 6, kneis2009bound has shown that the treewidth is bounded at most linearly by the edge size, or more precisely bounded by $e/5.769+O(\log n)$ where $n\coloneqq|V|$ denotes the number of vertices in $G$.
+However, we do not know whether this bound can be further improved.
+Below we provide the maximum tree-width that is possible for a graph with edge size no greater than $15$, to hopefully serve as a starting point of fully addressing this problem.
+Observation 1 (Treewidth and Edge Count).
+Let
+$\mathrm{t}(e)\coloneqq\max\left\{\mathsf{tw}(G)\;\middle|\;G\text{ is a graph where }|E(G)|=e\right\}.$
+We have computed the exact values of $\mathrm{t}(e)$ for all $1\leq e\leq 15$ as follows:
+[…]
+[…]
+[…]
 We begin by recalling some well-known bounds on treewidth.
@@
 The degeneracy of a graph $G$, denoted by $D(G)$, is defined as the maximum value of the minimum degree $\delta(H)$ over all subgraphs $H$ of $G$.
-Lemma 9 (Bounds on Treewidth).
+We first present some results on inequalities involving treewidth. The first two inequalities are from bodlaender2011treewidth, while the last one follows directly from Definition 7.
+Lemma 9.
 Let $G=(V,E)$ be a graph. Then
@@
     for any vertex $v\in V$, it holds that $\mathsf{tw}(G)\leq\max\{\deg(v,G),\mathsf{tw}(\mathsf{elim}_{v}(G))\}$.
-Proof.
-The first inequality $D(G)\leq\mathsf{tw}(G)$ is established in Lemma 11 of bodlaender2011treewidth.
-The second inequality is stated as Lemma 8 in the same reference.
-The final claim follows directly from Definition 7: treewidth is defined as the minimum over all possible vertex elimination orderings, whereas the right-hand side corresponds to the class of orderings in which the vertex $v$ is eliminated first.
-∎
-Lemma 10 (Maximum Treewidth under Vertex and Edge Constraints).
+We then present results concerning the treewidth of graphs subject to constraints on the number of edges or vertices.
+Lemma 10.
 Let $\mathcal{G}_{n,e}$ be the set of all simple graphs with $n$ vertices, $e$ edges, and no isolated vertices, and let $\mathcal{G}_{n,e}=\emptyset$ if no such graphs exist.
@@
   1. 1.
-    Support set: For any fixed $e\in\mathbb{N}^{+}$, $\mathcal{G}_{n,e}$ is nonempty iff $n\in\mathrm{n}(e)\coloneqq\left\{n\in\mathbb{N}^{+}\,\middle|\,\binom{n}{2}\geq e,n\leq 2e\right\}$.
+    Support set: For any fixed $e\in\mathbb{N}^{+}$, $\mathcal{G}_{n,e}$ is nonempty if and only if
+    $n\in\mathrm{n}(e)\coloneqq\left\{n\in\mathbb{N}^{+}\,\middle|\,\binom{n}{2}\geq e,n\leq 2e\right\}.$
   2. 2.
@@
   4. 4.
-    Complete-graph treewidth: $\mathrm{tt}(n,\binom{n}{2})=n-1$, $\mathrm{tt}(n,\binom{n}{2}-1)=n-2$.
+    Treewidth of complete graph: $\mathrm{tt}(n,\binom{n}{2})=n-1$, $\mathrm{tt}(n,\binom{n}{2}-1)=n-2$.
   5. 5.
@@
     • (1)
-      If $1\leq\frac{2e}{n}<2$, then $\mathrm{tt}(n,e)\leq\max\{1,\mathrm{t}(e-1)\}$;
+      If $1\leq\frac{2e}{n}<2$, then $\mathrm{tt}(n,e)\leq\max\{1,\mathrm{t}(e-1)\}$.
     • (2)
-      If $1\leq\frac{2e}{n}<3$, then $\mathrm{tt}(n,e)\leq\max\{2,\mathrm{t}(e-1)\}$;
+      If $1\leq\frac{2e}{n}<3$, then $\mathrm{tt}(n,e)\leq\max\{2,\mathrm{t}(e-1)\}$.
     • (3)
@@
 Figure 3: Graphs $G_{1}$ to $G_{15}$ with increasing edges and maximum treewidth.
-C.2 Proof of Results in the Rightmost Column of Table 1
-We can derive the sharp treewidth bounds presented in Table 1 using similar techniques to the above.
-For each $m\in\{2,3,\cdots,12\}$, the corresponding decomposition signature takes the following form: ${\mathcal{A}}_{m}=((1,2),\cdots,(m-1,m))$.
-Hence by Definition 10, the corresponding decomposition graph $G_{{\mathcal{A}}_{m}}$ is the graph with $V(G_{{\mathcal{A}}_{m}})=[m]$, $E(G_{{\mathcal{A}}_{m}})=\{\{i,i+1\}\}_{i=1}^{m-1}$, whose structure is illustrated in Figure 4(a).
+[…]
+The following table lists the total number of $V$-statistics and the number of floating-point operations (FLOPs) that need to be computed before and after applying the sparsification trick when the sample size is $10^{4}$.
+The rightmost column of Table 4 also records the value of $M$ defined in Corollary 2.
+Table 4: Exact number of floating-point operations (FLOPs), $V$-statistic decompositions, intermediate tensor sizes, and the maximum treewidth $M$ across all decomposition graphs for HOIF orders from $2$ to $12$, with and without the sparsification trick, where the Einsum ordering is found by applying opt-einsum with optimization scheme greedy.
+Sample size is fixed at $n=10000$.
+Order ($m$) V-stat terms (Bell number) V-stat terms (sparsification) FLOPs (initial) FLOPs (sparsification) $M$
+2 2 1 $2.00020\times 10^{8}$ $2.00000\times 10^{8}$ 11
+3 5 2 $060\times 10^{12} 2.00020$$020\times 10^{12}$ 11
+4 15 5 $8.00310\times 10^{12}$ $4.00150\times 10^{12}$ 22
+5 52 15 $5.20133\times 10^{13}$ $60045\times 10^{13} 22$
+6 203 52 $2.94069\times 10^{14}$ $1.20020\times 10^{14}$ 22
+7 877 203 $1.78236\times 10^{15}$ $6.32091\times 10^{14}$ 22
+8 4140 877 $31054\times 10^{17} 2.2$$23430\times 10^{17}$ 33
+9 21147 4140 $5.29126\times 10^{18}$ $47959\times 10^{18} 33$
+10 115975 21147 $2.07990\times 10^{21}$ $2.04048\times 10^{21}$ 33
+11 678570 115975 $5.50026\times 10^{22}$ $3.72144\times 10^{22}$ 44
+12 4213597 678570 $1.13401\times 10^{24}$ $6.62181\times 10^{23}$ 44
+[…]
 The result is formally stated in the following observation.
@@
 The desired result then follows from a characterization of the lower bound of $M(m)$.
-Table 7: Upper bounds of $M(m)$ at different $m$’s.
+Table 5: Upper bounds of $M(m)$ for HOIFs: via $M(m)\leq\mathrm{t}(m-1)$.
 $m$ $e$ $\mathrm{t}(e)$
@@
 We observe that except for $m=7$, the upper bounds match the claimed results in (7). The lower bounds of $M(m)$ can be obtained by enlisting some concrete examples, which are given in Table 6.
-Table 8: Lowers bound of $M(m)$: the exemplified graph whose treewidth reaches the lower bound of $M(m)$ at different $m$’s.
+Table 6: Lowers bound of $M(m)$ for HOIFs: via finding the exemplified graph.
 $m$ $\pi$ Graph $G_{\mathcal{A}_{m}}/\pi$ $\mathsf{tw}(G_{\mathcal{A}_{m}}/\pi)$
@@
 (b) $G^{\prime}_{6}:$ the example in Table 6 when $m=7$.
-Figure 4: Two decomposition graphs used in Appendix C.2.
+[…]
 Combining Table 5 and Table 6, we can conclude that, except for $m=7$, the upper bound (given in Table 5) and the lower bound (given in Table 6) of $M(m)$ agree.
@@
 Then, another lemma (Lemma 12) will show that the set of all quotient graphs of $G_{\mathcal{A}_{7}}$ does not contain $K_{4}$.
-The proof is now complete.
+The proof is then complete.
 ∎
@@
 $c(k)=i\quad\text{for all }k\in Q_{i}.$
-This map can be interpreted as a coloring of the vertices: each vertex $k$ is colored by the index of the subset $B_{i}$ which it belongs to, so that all vertices in $Q_{i}$ are dyed by color $i$.
-By above analysis, each pair of adjacent vertices is assigned different colors, thus for each $i\in[6]$, $\{Q_{c(i)},Q_{c(i+1)}\}$ is an edge in $G_{{\mathcal{A}}_{7}}/\pi^{*}$.
-Then, the sequence of edges
-$\{Q_{c(1)},Q_{c(2)}\},\ \{Q_{c(2)},Q_{c(3)}\},\ \{Q_{c(3)},Q_{c(4)}\},\ \{Q_{c(4)},Q_{c(5)}\},\ \{Q_{c(5)},Q_{c(6)}\},\ \{Q_{c(6)},Q_{c(7)}\}$
-corresponds to an Euler trail (i.e., a sequence of edges in which consecutive edges share a common endpoint, visiting every edge of the graph exactly once without repetition.) in the quotient graph $G_{{\mathcal{A}}_{7}}/\pi^{*}=K_{4}$.
+By above analysis, all 6 edges in $G_{\mathcal{A}_{7}}/\pi^{*}$ must be
+$\{Q_{c(1)},Q_{c(2)}\},\ \{Q_{c(2)},Q_{c(3)}\},\ \{Q_{c(3)},Q_{c(4)}\},\ \{Q_{c(4)},Q_{c(5)}\},\ \{Q_{c(5)},Q_{c(6)}\},\ \{Q_{c(6)},Q_{c(7)}\},$
+which forms an Euler trail (i.e., a sequence of edges in which consecutive edges share a common endpoint, visiting every edge of the graph exactly once without repetition.) in the quotient graph $G_{\mathcal{A}_{7}}/\pi^{*}=K_{4}$.
 However, by Corollary 4.1 in bondy1976graph, a connected undirected graph admits an Euler trail if and only if it has exactly 0 or $2$ vertices of odd degree.
@@
 ∎
+[…]
+In this section, we first present the $U$-statistic representations for all 3-vertex and 4-vertex motifs mentioned in Section 4.2.
+Letting $B\equiv 1-A$, the corresponding $U$-statistics for motif counts for specific $3$-vertex motifs and $4$-vertex motifs (see Figure 5 for which motifs $\mathsf{r}_{1}$ to $\mathsf{r}_{8}$ encode) read as follows:
+$\mathsf{C}(\mathsf{r}_{1},G)$ $=\frac{1}{2}\sum_{\bar{i}_{3}\in\mathsf{Perm}(n,3)}A_{i_{1},i_{2}}A_{i_{2},i_{3}}B_{i_{3},i_{1}},$
+$\mathsf{C}(\mathsf{r}_{2},G)$ $=\frac{1}{6}\sum_{\bar{i}_{3}\in\mathsf{Perm}(n,3)}A_{i_{1},i_{2}}A_{i_{2},i_{3}}A_{i_{3},i_{1}},$
+$\mathsf{C}(\mathsf{r}_{3},G)$ $=\frac{1}{6}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}B_{i_{2},i_{3}}B_{i_{2},i_{4}}B_{i_{3},i_{4}},$
+$\mathsf{C}(\mathsf{r}_{4},G)$ $=\frac{1}{2}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}B_{i_{1},i_{3}}B_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
+$\mathsf{C}(\mathsf{r}_{5},G)$ $=\frac{1}{2}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}B_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
+$\mathsf{C}(\mathsf{r}_{6},G)$ $=\frac{1}{8}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}B_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}B_{i_{2},i_{4}}A_{i_{3},i_{4}},$
+$\mathsf{C}(\mathsf{r}_{7},G)$ $=\frac{1}{4}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}A_{i_{2},i_{4}}B_{i_{3},i_{4}},$
+$\mathsf{C}(\mathsf{r}_{8},G)$ $=\frac{1}{24}\sum_{\bar{i}_{4}\in\mathsf{Perm}(n,4)}A_{i_{1},i_{2}}A_{i_{1},i_{3}}A_{i_{1},i_{4}}A_{i_{2},i_{3}}A_{i_{2},i_{4}}A_{i_{3},i_{4}}.$
+(a) $\mathsf{r}_{1}$: V-shape
+(b) $\mathsf{r}_{2}$: Triangle
+(c) $\mathsf{r}_{3}$ : 3-star
+(d) $\mathsf{r}_{4}$ : Fork
+(e) $\mathsf{r}_{5}$ : Tailed triangle
+(f) $\mathsf{r}_{6}$: Square
+(g) $\mathsf{r}_{7}$: House
+(h) $\mathsf{r}_{8}$ : 4-clique
+Figure 5: All non-degenerate isomorphism classes of 3-vertex and 4-vertex simple undirected graphs.
+Therefore, motif counts for $m$-vertex motifs naturally correspond to a complete decomposition signature $\mathcal{A}_{\mathsf{Motif},m}=((s,t))_{1\leq s<t\leq m}$, where the associated decomposition graph $G_{\mathcal{A}_{\mathsf{Motif},m}}$ is the complete graph $K_{m}$.
+This structure has two important implications.
+First, by leveraging our sparsification technique (Lemma 2 and Remark 4), the computation reduces to a single dominant term, i.e.,
+$\Pi_{m}^{\mathcal{A}_{\mathsf{Motif},m}}=\{\pi_{m}\},\quad\text{where }\pi_{m}\coloneqq\big\{\{1\},\{2\},\ldots,\{m\}\big\}.$
+Second, this also reveals a fundamental limitation. Even after sparsification, the resulting term still induces the same complete decomposition structure,
+$G_{\mathcal{A}_{\mathsf{Motif},m}}/\pi_{m}=K_{m},$
+and therefore the computational complexity of exact motif counting cannot be reduced below $O(n^{m})$ within our framework.
+We next present the experimental results for 4-vertex motif counting and GPU-based triangle counting, which are summarized in Table 7 and Table 8, respectively.
+The overall message is quite similar to that in Table 2 presented in Section 4.2.
+Table 7: Runtime comparison of exact all $4$-vertex motif counts using u-stats, Peregrine and igraph on Erdős–Rényi graphs $G(n,p)$ with $n=2000$.
+Experiments were run on Intel Xeon Scalable Cascade Lake 6248 CPUs (2.5GHz, 40 total cores) with memory of 192 GB.
+“OOT” denotes instances that exceeded the time limit of 3600 seconds.
+Edge Prob. $p$ u-stats Time (s) Peregrine Time (s) igraph Time (s)
+0.0005 3.8413 0.1024 0.0752
+0.0010 18.7463 0.1141 0.0873
+0.0050 41.0730 0.1132 0.2645
+0.0100 40.7701 0.1116 1.7099
+0.0500 41.5188 0.2079 610.0180
+0.1000 40.5648 0.8243 OOT
+0.2000 41.2671 5.8045 OOT
+0.4000 43.3735 48.7363 OOT
+0.6000 39.6022 159.4699 OOT
+0.8000 40.6578 292.4395 OOT
+Table 8: Runtime comparison of exact triangle counting using u-stats and cuGraph on Erdős–Rényi graphs $G(n,p)$ with $n=10000$.
+Experiments were run on a single GPU (NVIDIA RTX 4090, 24GB) “OOM” indicates out-of-memory.
+Edge Prob. $p$ u-stats Time (s) cuGraph Time (s)
+0.001 2.8014 0.1536
+0.005 2.9564 0.5664
+0.010 3.3363 1.0379
+0.020 3.7830 1.9797
+0.050 5.0856 4.9804
+0.080 6.3764 8.1346
+0.100 7.2829 10.3661
+0.150 9.6873 15.3828
+0.200 11.7283 OOM
+0.800 37.6880 OOM